### v1

# IsoSEL: Isometric Structural Entropy Learning for Deep Graph Clustering in Hyperbolic Space

Abstract
Graph clustering is a longstanding topic in machine learning. In recent years, deep learning methods have achieved encouraging results, but they still require predefined cluster numbers $K$, and typically struggle with imbalanced graphs, especially in identifying minority clusters. The limitations motivate us to study a challenging yet practical problem: deep graph clustering without $K$ considering the imbalance in reality. We approach this problem from a fresh perspective of information theory (i.e., structural information). In the literature, structural information has rarely been touched in deep clustering, and the classic definition falls short in its discrete formulation, neglecting node attributes and exhibiting prohibitive complexity. In this paper, we first establish a new Differentiable Structural Information, generalizing the discrete formalism to continuous realm, so that the optimal partitioning tree, revealing the cluster structure, can be created by the gradient backpropagation. Theoretically, we demonstrate its capability in clustering without requiring $K$ and identifying the minority clusters in imbalanced graphs, while reducing the time complexity to $O(N)$ w.r.t. the number of nodes. Subsequently, we present a novel IsoSEL framework for deep graph clustering, where we design a hyperbolic neural network to learn the partitioning tree in the Lorentz model of hyperbolic space, and further conduct Lorentz Tree Contrastive Learning with isometric augmentation. As a result, the partitioning tree incorporates node attributes via mutual information maximization, while the cluster assignment is refined by the proposed tree contrastive learning. Extensive experiments on five benchmark datasets show the IsoSEL outperforms $14$ recent baselines by an average of $+1.3\%$ in NMI.
Index Terms:
Graph Clustering, Deep Learning, Structural Entropy, Hyperbolic Space, Lorentz Group.
Codeshttps://github.com/RiemannGraph/DSE_clustering
I Introduction
Graph clustering shows fundamental importance for understanding graph data and has a wide range of applications, including biochemical analysis, recommender system and community detection [1, 2]. In recent years, deep learning has been introduced to graph clustering, where graph neural networks [3, 4] learn informative embeddings with a clustering objective performed in the representation space [5, 6]. While achieving remarkable success, deep graph clustering still faces two major challenges:
  • The requirement of predefined cluster numbers $K$. So far, deep graph clustering cannot work without a predefined cluster number $K$, and thus one is required to correctly provide $K$ before clustering, which is often impractical in real cases. On the one hand, it is challenging to estimate the number of node clusters. The methods, e.g., Bayesian information criterion [7] and recent RGC [8], repeatedly train and test the deep model to search the optimal $K$, and the searching process typically results in high computation overhead. On the other hand, existing solutions without graph structures (e.g., Bayesian non-parametric methods [9], DBSCAN [10], and hierarchical clustering [11]) cannot be directly applied owing to the inter-correlation among nodes. That is, deep graph clustering without $K$ is largely underexplored.
  • Clustering the imbalanced graphs. The imbalance is prevalent in real-world graphs, where a large proportion of vertices belong to the majority clusters, while minority clusters contain significantly fewer vertices [12, 13]. Take the popular Cora dataset [14] as an example: $26.8\%$ of the papers belong to Neural Network cluster, whereas only $4.8\%$ fall under Rule Learning cluster. Deep graph clustering models typically rely on the representation learning, and minority clusters tend to be underrepresented due to their limited sample sizes compared to the majority clusters [15]. This imbalance increases the risk of generating biased clustering boundaries that favor majority clusters, as evidenced by empirical observations in our Case Study. Also, the challenge becomes even greater when the number of clusters is unknown.
Figure 1: In IsoSEL, we learn the partitioning tree of structural information with the LSEnet in hyperbolic space, and then propose Lorentz tree contrastive learning to refine the partitioning tree for graph clustering without $K$.
In this paper, we revisit deep graph clustering with an unknown $K$, considering the imbalance in real-world graphs, and approach this problem from a fresh perspective based on information theory. Stemming from Shannon entropy, structural information [16] has been proposed to measure the uncertainty in graphs. According to structural information minimization, the optimal partitioning tree can be constructed to describe a graph’s self-organization without requiring $K$. While this sheds light on the targeted problem, it presents several significant challenges to deep clustering: (1) Connection to Clustering. Structural information has recently garnered increasing research attention [17, 18, 19], but its theoretical connection to clustering remains unclear. As a result, structural information has not yet been introduced to graph clustering. (2) Discrete Formulation. The discrete nature of structural information prevents gradient backpropagation, posing a major challenge in training deep graph models. (3) Attribute Ignorance. Structural information operates solely on the graph structure overlooking node attributes, which are often equally important to graph clustering. (4) Computational Complexity. Constructing a partitioning tree of height $3$ yields the complexity of $O(\lvert N\rvert^{3}\log^{2}\lvert N\rvert)$, where $N$ is the number of nodes. Note that, the complexity scales exponentially with tree height [16], making it infeasible for real-world graphs.
To tackle these challenges, we establish a new differentiable structural information (DSI), generalizing the discrete formalism to the continuous realm. In DSI, we reformulate the structural information with level-wise assignment matrices. We prove that DSI upper bounds the graph conductance, a well-defined clustering objective, and thus DSI emerges as a new differentiable graph clustering objective without requiring the number of clusters $K$. Thus we create a deep partitioning tree $\mathcal{T}_{\operatorname{net}}$ parametrized by a neural network. Different from the partitioning tree of [16] constructed by heuristic algorithm, the proposed deep tree can be learned by gradient backpropagation according to DSI minimization. Theoretically, we show that the structural entropy of the optimal $\mathcal{T}_{\operatorname{net}}$ well approximates that of [16] under slight constraint. Next, we reduce the time complexity to the linear $\mathcal{O}(N)$ with respect to the number of nodes by exploring the equivalence in level-wise assignments. Furthermore, we discover another advantage of the deep partitioning tree that it is able to identify the minority clusters from the majority ones, which is suitable for clustering real graphs that often exhibit imbalanced cluster distributions.
Grounded on DSI theory, we present a novel Isometric Structural Entropy Learning (referred to as IsoSEL) to learn and refine the deep partitioning tree in the hyperbolic space for graph clustering without $K$. The overall architecture is illustrated in Fig. 1, consisting of the Lorentz Structural Entropy net (LSEnet) from the conference version [^1] The preliminary results have been published in the proceedings of the 41st International Conference on Machine Learning (ICML) as an Oral Paper [20]., and Lorentz tree contrastive learning proposed in this full version. As in the part B of Fig. 1, LSEnet is designed to parameterize the deep partitioning tree in the Lorentz model of hyperbolic space. The novelty lies in exploring the alignment between hyperbolic space and hierarchical/tree-like structures. Specifically, we first embed the leaf nodes of the partitioning tree using Lorentz convolution over the graph. Then, we recursively learn the parent nodes from the bottom (leaf nodes) to the top (root node) using the proposed Lorentz assigner.
In IsoSEL, another key innovation is the proposed Lorentz Tree Contrastive Learning (as shown in part C of Fig. 1), which refines the deep partitioning tree in the hyperbolic space. To achieve this, we first formulate the isometric augmentation to generate an augmented graph, integrating attribute similarity into structural information from the representation perspective using the notion of Gauss curvature. Second, we design a projection head for tree contrastive learning. Unlike graph contrastive learning, the main challenge in contrasting trees lies in ensuring consistency between leaf and parent nodes in hyperbolic space, given the unified nature of the partitioning tree. To address this, we instantiate a neural Lorentz group in hyperbolic space. Specifically, we parameterize the transformation in the Lorentz group using a closed form matrix-vector multiplication, which supports gradient descent optimization. This parameterization ensures that the projection head preserves the hyperbolic manifold and maintains consistency between leaf and parent nodes, leveraging the linearity of the Lorentz group. Finally, we utilize the Minkowski inner product to contrast the deep partitioning trees, refining the hyperbolic representations and enhancing cluster assignment. As a result, IsoSEL seamlessly integrates the strengths of both structural entropy and hyperbolic space for graph clustering.
Contribution Highlights. The key contributions are five-fold:
  • Problem. We revisit deep graph clustering without a predefined cluster number, considering the imbalance of real graphs and, to the best of our knowledge, bridge graph clustering and structural entropy for the first time.
  • Theory. We establish the differentiable structural information (DSI) in continuous realm, so that a deep partitioning tree with linear computational complexity is created by gradient backpropagation for graph clustering without $K$.
  • Neural Architecture. We present a novel IsoSEL approach where the hyperbolic LSEnet is designed to learn the deep partitioning tree with Lorentz convolution and assigner, connecting DSI to hyperbolic geometry.
  • Learning Paradigm. We propose a Lorentz tree contrastive learning with the isometric augmentation to refine the deep partitioning tree in the hyperbolic space, while incorporating the attribute information.
  • Experiment. Extensive real world results show the superiority of IsoSEL for graph clustering, and we examine that biased boundary among imbalanced clusters is alleviated with case study and visualization.
II Preliminaries and Notations
Different from the typical graph clustering, we study a challenging yet practical problem of graph clustering without the number of clusters, considering the imbalance of real graphs. This section formally reviews the basic concepts of structural information, hyperbolic space and Lorentz group, and introduce the main notations.
II-1 Graph & Graph Clustering
An attribute graph $G=(\mathcal{V},\mathcal{E},\mathbf{X})$ is defined on the node set $\mathcal{V}$, $|\mathcal{V}|=N$. The edge set $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$ is associated with a weight function $w:\mathcal{E}\rightarrow\mathbb{R}^{+}$. Alternatively, $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the adjacency matrix collecting the edge weights. Each node is attached to a $d$-dimension attribute vector, summarized in $\mathbf{X}\in\mathbb{R}^{N\times D}$. Graph clustering aims to group nodes into several disjoint clusters, so that nodes share higher similarity with the nodes in the same clusters than that of different clusters. Deep learning has become the dominant solution to graph clustering in recent years. In deep graph clustering, a self-supervised loss is carefully designed to generate informative embeddings, and clustering is conducted in the representation space. On the one hand, the cluster number $K$ is a typical prerequisite of deep models, but offering $K$ before clustering is impractical in real cases. On the other hand, real graphs are often imbalanced, exhibiting skewed class distributions. The minority cluster of fewer nodes tends to be underrepresented in the self-supervised learning owing to limited samples. However, clustering on imbalanced graphs is still underexplored in the literature. To bridge the gaps, we are interested in deep graph clustering without $K$, considering the imbalance of real graphs.
II-2 Structural Information & Structural Entropy
Given a graph $G=(\mathcal{V},\mathcal{E})$ with weight function $w$, the weighted degree of node $v$ is written as $d_{v}=\sum\nolimits_{u\in\mathcal{N}(v)}w(v,u)$, the summation of the edge weights surrounding it. $\mathcal{N}(v)$ denotes the neighborhood of $v$. For an arbitrary node set $\mathcal{U}\subset\mathcal{V}$, its volume $\operatorname{Vol}(\mathcal{U})$ is defined by summing of weighted degrees of the nodes in $\mathcal{U}$, and $\operatorname{Vol}(G)$ denotes the volume of the graph. The one-dimensional structural entropy is defined as follows,
$\mathcal{H}^{1}(G)=-\sum_{v\in\mathcal{V}}\frac{d_{v}}{\operatorname{Vol}(G)}\log_{2}\frac{d_{v}}{\operatorname{Vol}(G)},\vspace{-0.03in}$ (1)
yielding as a constant for certain given graph. The $H$-dimensional structural information is accompanied with a partitioning tree $\mathcal{T}$ of height $H$. Concretely, in $\mathcal{T}$ with root node $\lambda$, each tree node $\alpha$ is associated with a subset of $\mathcal{V}$, referred to as module $T_{\alpha}\subset\mathcal{V}$ [^2] The vertex of partitioning tree is termed as “tree node” to distinguish from the node in graphs. and its immediate predecessor is written as $\alpha^{-}$. The leaf node of the tree corresponds to the node of graph $G$. Accordingly, the module of the leaf node is a singleton, while $T_{\lambda}$ is the node set $\mathcal{V}$. The structural information assigned to each non-root node $\alpha$ is defined as
$\mathcal{H}^{\mathcal{T}}(G;\alpha)=-\frac{g_{\alpha}}{\operatorname{Vol}(G)}\log_{2}\frac{V_{\alpha}}{V_{\alpha^{-}}},\vspace{-0.03in}$ (2)
where the scalar $g_{\alpha}$ is the total weights of graph edges with exactly one endpoint in the $T_{\alpha}$, and $V_{\alpha}$ is the volume of the module $T_{\alpha}$. The $H$-dimensional structural information of $G$ with respect to the partitioning tree $\mathcal{T}$ is written as follows,
$\mathcal{H}^{\mathcal{T}}(G)=\sum\nolimits_{\alpha\in\mathcal{T},\alpha\neq\lambda}\mathcal{H}^{\mathcal{T}}(G;\alpha).\vspace{-0.03in}$ (3)
Traversing all possible partitioning trees of $G$ with height $H$, the $H$-dimensional structural entropy of $G$ is defined as
$\mathcal{H}^{H}(G)=\min_{\mathcal{T}}\mathcal{H}^{\mathcal{T}}(G),\quad\mathcal{T}^{*}=\arg_{\mathcal{T}}\min\mathcal{H}^{\mathcal{T}}(G),\vspace{-0.03in}$ (4)
where $\mathcal{T}^{*}$ is the optimal partitioning tree of $G$ that minimizes the uncertainty of graph, and encodes the self-organization of node clusters in the tree structure.
II-3 Hyperbolic Space & Lorentz Group
Riemannian geometry offers an elegant framework to study complex structures. A Riemannian manifold is described as a smooth manifold $\mathbb{M}$ endowed with a Riemannian metric. Each point $\boldsymbol{x}$ in the manifold is associated with the tangent space $\mathbb{T}_{\boldsymbol{x}}\mathbb{M}$ where the Riemannian metric is defined. Mapping between the tangent space and manifold is done via the exponential and logarithmic maps. The notion of curvature $\kappa$ describes the extent of how a manifold derivatives from being “flat”. Accordingly, there exist three types of isotropic manifold: hyperbolic space of negative curvature, hyperspherical space of positive curvature and the special case of zero-curvature Euclidean space with “flat” geometry. In Riemannian geometry, a structure is related to certain manifold and tree structure is aligned with hyperbolic space. Here, we visit the Lorentz model of hyperbolic space to study the partitioning tree of structural entropy. Without loss of generality, a $d-$dimensional Lorentz model $\mathbb{L}^{\kappa,d}$ with curvature $\kappa$ is defined on the smooth manifold embedded in $\mathbb{R}^{d+1}$ space,
$\mathbb{L}^{\kappa,d}=\{\boldsymbol{x}\in\mathbb{R}^{d+1}:\langle\boldsymbol{x},\boldsymbol{x}\rangle_{\mathbb{L}}=\frac{1}{\kappa},x_{0}>0\},\vspace{-0.03in}$ (5)
where the Minkowski inner product is given as follows,
$\langle\boldsymbol{x},\boldsymbol{y}\rangle_{\mathbb{L}}=\boldsymbol{x}\mathbf{R}\boldsymbol{y},\quad\mathbf{R}=\operatorname{diag}(-1,1,\cdots,1),\vspace{-0.03in}$ (6)
$\mathbf{R}\in\mathbb{R}^{(d+1)\times(d+1)}$ describes the Minkowski metric in $\mathbb{R}^{d+1}$ space. $x_{0}$ is the first element of point coordinates. For arbitrary two points in the Lorentz model, $\forall\boldsymbol{x},\boldsymbol{y}$, the induced distance is given as $d_{\mathbb{L}}(\boldsymbol{x},\boldsymbol{y})=\sqrt{-\kappa}\operatorname{arccosh}(-\langle\boldsymbol{x},\boldsymbol{y}\rangle_{\mathbb{L}})$. $\mathbb{T}_{\boldsymbol{x}}\mathbb{L}^{\kappa,d}$ denotes the tangent space at $\boldsymbol{x}$, and $\lVert\boldsymbol{u}\rVert_{\mathbb{L}}=\sqrt{\langle\boldsymbol{u},\boldsymbol{u}\rangle_{\mathbb{L}}}$ is known as Lorentz norm for $\boldsymbol{u}\in\mathbb{T}_{\boldsymbol{x}}\mathbb{L}^{\kappa,d}$.
In group theory, the Lorentz group studies the Lorentz model of hyperbolic space. To be specific, Lorentz group is a (generalized) orthogonal transformation group that collects the Lorentz transformation, preserving the isometry regarding the Minkowski metric. With polar decomposition, any Lorentz transformation can be decomposed into a combination of Lorentz boost and Lorentz rotation [21, 22]. Both of them are given as special orthogonal matrices, posing a significant challenge to parameterization. Please refer to Appendix B for further facts.
II-4 Notations
Throughout this paper, $\mathbb{L}$ and $\mathcal{T}$ denote the Lorentz model of hyperbolic space and the partitioning tree, respectively. The vector and matrix are written in the lowercase boldfaced $\boldsymbol{x}$ and uppercase $\mathbf{X}$, respectively. The notation table is provided in Appendix I.
In this paper, we introduce the novel Differentiable Structural Information (DSI) and present Isometric Structural Entropy Learning (IsoSEL) for deep graph clustering. The key innovation of DSI lies in constructing a deep partitioning tree in hyperbolic space through gradient backpropagation. Additionally, we propose Lorentz tree contrastive learning with theoretical guarantees to further refine the deep partitioning tree, enhancing cluster assignment.
III Differentiable Structural Information
In this section, we establish the theory of differentiable structural information, which generalizes the classic discrete formalism to the continuous realm. The main theoretical results are dual: the proposed formulation emerges as a new objective for deep graph clustering without $K$; the corresponding partitioning tree can be created via gradient-based deep learning, different from the previous heuristic algorithm [16].
III-A New Formulation
The structural information in Eq. (1) cannot be optimized via gradient-based methods. To bridge this gap, we present a new formulation of structural information with learnable level-wise assignment, which is shown to be equivalent to the classic formulation in Eq. (3).
Definition III.1 (Level-wise Assignment).
[…]
Definition III.2 (H-dimensional Structural Information).
For a graph $G$ and its partitioning tree $\mathcal{T}$ in Definition III.1, we rewrite the formula of $H$-dimensional structural information of $G$ with respect to $\mathcal{T}$ at height $h$ as
$\mathcal{H}^{\mathcal{T}}(G;h)=-\frac{1}{V}\sum\limits_{k=1}^{N_{h}}(V^{h}_{k}-\sum\limits_{(i,j)\in\mathcal{E}}S^{h}_{ik}S^{h}_{jk}w_{ij})\log_{2}\frac{V^{h}_{k}}{V^{h-1}_{k^{-}}}\vspace{-0.03in}$ (7)
where $V=\operatorname{Vol}(G)$ is the volume of $G$. For the $k$-th node in height $h$, $V^{h}_{k}$ and $V^{h-1}_{k^{-}}$ are the volume of graph node sets $T_{k}$ and $T_{k^{-}}$, respectively. Thus we have
$\mathbf{S}^{h}$ $=\prod\nolimits_{k=H+1}^{h+1}\mathbf{C}^{k},\quad\mathbf{C}^{H+1}=\mathbf{I}_{N},$ (8)
$V^{h}_{k}$ $=\sum\nolimits_{i=1}^{N}S^{h}_{ik}d_{i},\quad V^{h-1}_{k^{-}}=\sum\nolimits_{k^{\prime}=1}^{N_{h-1}}\mathbf{C}_{kk^{\prime}}^{h}V^{h-1}_{k^{\prime}}.\vspace{-0.03in}$ (9)
Then, the $H$-dimensional structural information of $G$ is $\mathcal{H}^{\mathcal{T}}(G)=\sum\nolimits_{h=1}^{H}H^{\mathcal{T}}(G;h)$.
{mymath}
Theorem III.3 (Equivalence).
The formula $\mathcal{H}^{\mathcal{T}}(G)$ in Definition III.2 is equivalent to Eq. (3).
Proof.
Please refer to Appendix II, Theorem III.3. ∎
III-B Connecting Structural Information and Graph Clustering
This part shows some general properties of the proposed formulation and establishes the connection between structural entropy and graph clustering. We first give an arithmetic property regarding Definition III.2 to support the following claim on graph clustering. The proofs of the lemma/theorems are detailed in Appendix II.
Lemma III.4 (Additivity).
The $1$-dimensional structural entropy of $G$ can be decomposed as follows
$\mathcal{H}^{1}(G)=\sum_{h=1}^{H}\sum_{j=1}^{N_{h-1}}\frac{V^{h-1}_{j}}{V}E([\frac{C^{h}_{kj}V^{h}_{k}}{V^{h-1}_{j}}]_{k=1,...,N_{h}}),\vspace{-0.03in}$ (10)
where $E(p_{1},...,p_{n})=-\sum_{i=1}^{n}p_{i}\log_{2}p_{i}$ is the entropy.
{mymath}
Theorem III.5 (Connection Graph Clustering).
Given a graph $G=(\mathcal{V},\mathcal{E})$ with $w$, the normalized $H$-structural entropy of graph $G$ is defined as $\tau(G;H)={\mathcal{H}^{H}(G)}/{\mathcal{H}^{1}(G)}$, and $\Phi(G)$ is the graph structural conductance. With the additivity (Lemma III.4), the following inequality holds,
$\tau(G;H)\geq\Phi(G).\vspace{-0.03in}$ (11)
Proof.
We show the key equations here, and further details are given in Appendix II Theorem III.5. Without loss of generality, we assume $\min\{V^{h}_{k},V-V^{h}_{k}\}=V^{h}_{k}$. From Definition III.2,
$\mathcal{H}^{\mathcal{T}}(G)$ $=-\frac{1}{V}\sum_{h=1}^{H}\sum_{k=1}^{N_{h}}\phi_{h,k}V^{h}_{k}\log_{2}\frac{V^{h}_{k}}{V^{h-1}_{k^{-}}}$
$\geq-\frac{\Phi(G)}{V}\sum_{h=1}^{H}\sum_{k=1}^{N_{h}}V^{h}_{k}\log_{2}\frac{V^{h}_{k}}{V^{h-1}_{k^{-}}},$ (12)
Let the $k$-th tree node in height $h$ denoted as $\alpha$, then $\phi_{h,k}$ is the conductance of graph node subset $T_{\alpha}$, and is defined as $\frac{\sum_{i\in T_{\alpha},j\in\bar{T}_{\alpha}}w_{ij}}{\min\{\operatorname{Vol}(T_{\alpha}),\operatorname{Vol}(\bar{T}_{\alpha})\}}$. where $\bar{T}_{\alpha}$ is the complement set of $T_{\alpha}$. The graph conductance is given as $\Phi(G)=\min_{h,k}\{\phi_{h,k}\}$. Next, we use Eq. (9) to obtain the following inequality
$\mathcal{H}^{\mathcal{T}}(G)$ $\geq-\frac{\Phi(G)}{V}\sum_{h=1}^{H}\sum_{k=1}^{N_{h}}\sum_{j=1}^{N_{h-1}}C^{h}_{kj}V^{h}_{k}\log_{2}\frac{V^{h}_{k}}{V^{h-1}_{k^{-}}}$
$\mathcal{H}^{\mathcal{T}}(G)$ $\geq-\frac{\Phi(G)}{V}\sum_{h=1}^{H}\sum_{j=1}^{N_{h-1}}V^{h-1}_{j}\sum_{k=1}^{N_{h}}\frac{C^{h}_{kj}V^{h}_{k}}{V^{h-1}_{j}}\log_{2}\frac{C^{h}_{kj}V^{h}_{k}}{V^{h-1}_{j}}$
$=\Phi(G)\sum_{h=1}^{H}\sum_{j=1}^{N_{h-1}}\frac{V^{h-1}_{j}}{V}E([\frac{C^{h}_{kj}V^{h}_{k}}{V^{h-1}_{j}}]_{k=1,...,N_{h}})$
$=\Phi(G)\mathcal{H}^{1}(G).\vspace{-0.03in}$ (14)
Since $\tau(G;\mathcal{T})=\frac{\mathcal{H}^{\mathcal{T}}(G)}{\mathcal{H}^{1}(G)}\geq\Phi(G)$ holds for every $H$-height partitioning tree $\mathcal{T}$ of $G$, $\tau(G)\geq\Phi(G)$ holds. ∎
The conductance is a well-defined objective for graph clustering [23]. $\forall G$, the one-dimensional structural entropy $\mathcal{H}^{1}(G)=-\frac{1}{\operatorname{Vol}(G)}\sum_{i=1}^{N}d_{i}\log_{2}\frac{d_{i}}{\operatorname{Vol}(G)}$ is yielded as a constant. Minimizing Eq. (7) is thus equivalent to minimizing conductance, grouping nodes into clusters. Also, Eq. (7) needs no knowledge of cluster number in the graph $G$. In short, our formulation is a differentiable objective function for graph clustering without predefined cluster numbers.
III-C Deep Partitioning Tree
Here, we elaborate on how to construct a deep partitioning tree through gradient backpropagation. Recall Eq. (7), our objective is differentiable over the parent assignment. If the assignment is relaxed to be the likelihood given by a neural net, our formulation supports learning a deep model that constructs the partitioning tree. For a graph $G$, we denote the deep partitioning tree as $\mathcal{T}_{\operatorname{net}}$ where the mapping from learnable embeddings $\mathbf{Z}$ to the assignment is done via a neural net. The differentiable structural information (DSI) is given as $\mathcal{H}^{\mathcal{T}_{\operatorname{net}}}(G;\mathbf{Z};\mathbf{\Theta})$. It takes the same form of Eq. (7) where the assignment is derived by a neural net with parameters $\Theta$. Given $G$, we consider DSI minimization as follows,
$\mathbf{Z}^{*},\mathbf{\Theta}^{*}=\arg_{\mathbf{Z},\mathbf{\Theta}}\min\mathcal{H}^{\mathcal{T}_{\operatorname{net}}}(G;\mathbf{Z};\mathbf{\Theta}).$ (15)
We learn node embeddings $\mathbf{Z}^{*}$ and parameter $\mathbf{\Theta}^{*}$ to derive the level-wise parent assignment for $G$. Then, the optimal partitioning tree $\mathcal{T}_{\operatorname{net}}^{*}$ of $G$ is constructed by the assignment, so that densely connected nodes have a higher probability to be assigned to the same parent in $\mathcal{T}_{\operatorname{net}}^{*}$, minimizing the conductance and presenting the cluster structure. Meanwhile, node embeddings jointly learned in the optimization provide geometric notions to refine clusters in representation space.
[…]
Theorem III.6 (Bound).
[…]
$\lvert\mathcal{H}^{\mathcal{T}^{*}}(G)-\mathcal{H}^{\mathcal{T}_{\operatorname{net}}^{*}}(G)\rvert\leq\mathcal{O}(\epsilon).$ (16)
Proof.
Refer to Appendix II, Theorem III.3. ∎
That is, the structural entropy calculated by our $\mathcal{T}_{\operatorname{net}}^{*}$ is well approximated to that of [16], and the difference is bounded under slight constraints. Thus, $\mathcal{T}_{\operatorname{net}}^{*}$ serves as an alternative to the optimal $\mathcal{T}^{*}$ for graph clustering.
IV Theory: Structural Information & Graph Clustering
In this section, we derive an equivalent form of differential structural information to reduce its complexity, and provide the new theoretical results on the advantages of structural information for graph clustering.
IV-A Reducing Complexity
Eq. (7) requires all edges to compute structural information at each level. Instead of enumerating the edges, we reduce the complexity by searching all the parent nodes whose amount is much less than that of edges. With the assignment matrices, we rewrite the differentiable structural information as follows,
$\mathcal{H}^{\mathcal{T}}(G^{h};G^{h-1},h)=-\frac{1}{V}\sum_{k=1}^{N_{h}}(d^{h}_{k}-w^{h}_{kk})\log\frac{d^{h}_{k}}{\sum_{l=1}^{N_{h-1}}C_{kl}^{h-1}d_{l}^{h-1}},$
[…]
Theorem IV.1 (Node-wise Formulation).
At each level in the partitioning tree, the structural information calculated in Eq. (17) is equal to that of Eq. (8).
Proof.
In Eq. (7), we analyze it term by term. First,
$g^{h}_{k}=V^{h}_{k}-\sum\limits_{(i,j)\in\mathcal{E}}S^{h}_{ik}S^{h}_{jk}w_{ij}=\sum\limits_{(i,j)\in\mathcal{E}}S^{h}_{ik}(1-S^{h}_{jk})w_{ij},$
then we set $w^{h}_{kl}:=\sum_{i,j}^{N}S^{h}_{ik}S^{h}_{jk}w_{ij}=\sum\limits_{(i,j)\in\mathcal{E}}S^{h}_{ik}S^{h}_{jk}w_{ij}$ to define $\hat{g}^{h}_{k}$ and $d^{h}_{k}$ by
$\hat{g}^{h}_{k}=\sum\limits_{l,l\neq k}w^{h}_{kl}=\sum_{l,l\neq k}\sum\limits_{(i,j)\in\mathcal{E}}S^{h}_{ik}S^{h}_{jk}w_{ij}=\sum\limits_{(i,j)\in\mathcal{E}}S^{h}_{ik}(1-S^{h}_{jk})w_{ij},$
$d^{h}_{k}=\sum\limits_{l}w^{h}_{kl}=\sum\limits_{(i,j)\in\mathcal{E}}S^{h}_{ik}w_{ij}=\hat{g^{h}_{k}}+w^{h}_{kk}.$
We observe that $g^{h}_{k}=\hat{g}^{h}_{k}$ and $d^{h}_{k}=V^{h}_{k}$, substituting them into Eq. (7), the proof is completed. ∎
While keeping the space complexity unchanged, we significantly reduce the time complexity of calculating structural information. Note that, the formulation of Eq. (17) achieves the linear complexity with respect to the number of nodes.
IV-B Flexiblity of Partitioning Tree
We show that the relaxation of the partitioning tree with redundant nodes does not affect the value of structural information under certain conditions, thanks to the logarithmic in the definition. The flexibility of partitioning tree is claimed in the following theorem. {mymath}
Theorem IV.2 (Flexiblity).
$\forall G$, given a partitioning tree $\mathcal{T}$ and adding a node $\beta$ to get a relaxed $\mathcal{T}^{\prime}$ , the structural information remains unchanged, $H^{\mathcal{T}}(G)=H^{\mathcal{T}^{\prime}}(G)$, if one of the following conditions holds:
  1. 1.
    $\beta$ as a leaf node, and its module $T_{\beta}$ is an empty set.
  2. 2.
    $\beta$ is inserted between node $\alpha$ and its children nodes so that the modules $T_{\beta}$ and $T_{\alpha}$ are equal.
Proof.
Refer to Appendix II, Theorem IV.2. ∎
The theorem above will guide the neural architecture design for constructing the deep partitioning tree $\mathcal{T}_{\operatorname{net}}^{*}$.
IV-C Handling Imbalanced Clusters
We real another interesting fact that the structural entropy objective is preferable for detecting minority clusters. To formally describe this, we consider a scenario where an imbalanced graph $G$ contains two majority clusters $V_{1}$, $V_{2}$ and one minority $V_{\varepsilon}$. Partitioning $V_{\varepsilon}$ further reduces the structural information of the partitioning tree, thereby enhancing the identifiability of the minority cluster. {mymath}
Theorem IV.3 (Identifiability).
Given an imbalanced graph $G$ of $N$ nodes consisting of the majority clusters $V_{1}$, $V_{2}$ and minority $V_{\varepsilon}$, such that $|V_{1}|\gg|V_{\varepsilon}|$, $|V_{2}|\gg|V_{\epsilon}|$ and $N\gg 3$, structural information minimization makes $V_{\varepsilon}$ distinguishable from the majority clusters given the following inequality
$\mathcal{H}(G,V_{2})+\mathcal{H}(G,V_{1}\cup V_{\varepsilon})>\sum\nolimits_{i\in\{1,2,\varepsilon\}}\mathcal{H}(G,V_{i})$ (21)
holds for the slight condition that $|V_{1}|\geq g_{\alpha}+1$ where $g_{\alpha}$ is the cut between clusters $V_{1}$ and $V_{2}$.
Proof.
Refer to Appendix II, Theorem IV.3. ∎
Nodes within the cluster are typically densely connected while there are fewer links connecting the nodes of different clusters. The condition is easy to be satisfied in real graphs.
That is, we obtain the graph clustering objective with linear time complexity, the differential structural information in Eq. (17), which no longer requires the knowledge of the number of clusters and is suitable to identify minority clusters in real-world graphs often exhibiting skewed cluster distributions.
V Isometric Graph Augmentation
The structural information solely handles the graph structure, while node attributes are often equally important to graph clustering. To fill this gap, we propose isometric augmentation which integrates attribute information into graph structure, generalizing existing methods from representation perspective.
V-A Isometry between Manifolds
We demonstrate that arbitrary embedding induces a Riemannian manifold of hypersurface, and the existence of this manifold is claimed in the following lemma.
Lemma V.1 (Existence).
Given a set of points $\{\boldsymbol{p}_{i}\}_{i=1,..,N}$ on $\mathbb{R}^{n+1}$ derived from certain embedding, there exists a hypersurface $\mathcal{M}$ such that $\boldsymbol{p}_{i}\in\mathcal{M}$ holds for all $i$.
Proof.
Refer to Appendix II, Lemma V.1. ∎
Now, we elaborate on the isometry between manifolds. Specifically, we leverage the notion of Gaussian curvature, which depicts the local geometry of the manifold. Considering a point $\boldsymbol{p}_{i}\in\mathcal{M}$ on the manifold, the Gaussian curvature at $\boldsymbol{p}$ is intrinsically defined by the rate at which the geodesics on the manifold of hypersurface, emanating from $\boldsymbol{p}$, spread apart and this curvature is invariant among coordinate systems. Two hypersurfaces are said to be isometric if their Gaussian curvature are preserved relative to each other according to Gauss’s Theorema Egregium [24]. We develop the geometric intuition: The isometry means that two hypersurfaces exhibit the same pattern on how it curves at each point.
To find an isometry between hypersurfaces, we present our theoretical result on Givens rotation matrix $\mathbf{G}_{\theta}$. To be specific, the element of $\mathbf{G}_{\theta}$ is defined as follows,
$\mathbf{G}_{\theta}[i,j]=\mathbf{I}[i,j]+\beta_{ii}(\cos\theta-1)+\beta_{ij}\sin\theta,\vspace{-0.03in}$ (22)
where $\beta_{ii}=e_{i}e_{i}^{\top}+e_{j}e_{j}^{\top}$, $\beta_{ij}=e_{i}e_{j}^{\top}+e_{j}e_{i}^{\top}$, $\mathbf{I}$ is the identity, and $e$’s denote the orthogonal bases of the hypersurface. The matrix $\mathbf{G}_{\theta}$ characterizes the rotation along the span of $e_{i}$ and $e_{j}$ with the rotation angle of $\theta$. We prove the following statement holds for arbitrary hypersurfaces. {mymath}
Theorem V.2 (Rotational Isometry).
Given two hypersurfaces $\mathcal{S}_{1},\mathcal{S}_{2}$ in $\mathbb{R}^{n+1}$, equipped with the usual inner product, the Givens rotation $\mathbf{G}_{\theta}$ can be an isometry, such that $\mathbf{G}_{\theta}:\mathcal{S}_{1}\rightarrow\mathcal{S}_{2}$, and vice versa.
Proof.
Refer to Appendix II, Theorem V.2. ∎
In other words, the Givens rotation instantiate an isometry between hypersurfaces in terms of Gauss’s Theorema Egregium.
V-B Augmentation with Rotational Isometry
Typical integration methods [25] create an augmented graph $\hat{G}(\hat{\mathbf{A}},\mathbf{X})$ that aggregates the adjacency matrix and the similarity matrix of node attributes, given a graph $G(\mathbf{A},\mathbf{X})$. The augmented $\hat{\mathbf{A}}$ is given by the adjacency matrix $\mathbf{A}$ and the similarity $\mathbf{S}$ derived from the attributes $\mathbf{X}$, e.g., the structure of K-nearest neighbor graph based on attribute similarity.
In IsoSEL, we propose a more generic approach (referred to as isometric augmentation) from the representation perspective. The key idea is to perform aggregation on the manifolds and derive the augmented structure with the notion of isometry. First, we induce the manifold of $G$ with message-passing embedding, taking the popular SGC [26] as an example. Hence two sets of embeddings are obtained from adjacency matrix and attribute similarity, formulated as $\mathbf{A}^{L}\mathbf{X}$ and $\mathbf{S}^{L}\mathbf{X}$, respectively. The message-passing is conducted $L$ times without loss of generality. Second, we select an aggregation function $f$ to construct the hypersurface whose existence is guaranteed given Lemma V.1. $f$ can be a weighted sum or product, and the corresponding embedding is written as $f(\mathbf{A},\mathbf{S},\mathbf{X};L)$. Next, we aim to derive an augmented graph whose induced hypersurface is isometric to that of $f(\mathbf{A},\mathbf{S},\mathbf{X};L)$ in terms of Gauss’s Theorema Egregium. However, specifying the relationship between the embedding of isometric hypersurfaces is challenging. Thanks to the discovery of Theorem V.2, the augmented $\hat{\mathbf{A}}$ can be formulated in the optimization problem as follows. {mymath}
Theorem V.3 (Optimization of Isometric Augmentation).
Given a graph with an adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ and a similarity matrix $\mathbf{S}\in\mathbb{R}^{N\times N}$ induced by the attributes $\mathbf{X}\in\mathbb{R}^{N\times d}$, the adjacency matrix of augmented graph $\hat{\mathbf{A}}\in\mathbb{R}^{N\times N}$ considering the isometry in the $L$ convolution is the closed-form minimizer of the following optimization
$\min\ \|f(\mathbf{A},\mathbf{S},\mathbf{X};L)\mathbf{G}^{\top}_{\theta}-\hat{\mathbf{A}}^{L}\mathbf{X}\|^{2},\vspace{-0.05in}$ (23)
subject to the symmetry matrix cone of $\hat{\mathbf{A}}$, where $\mathbf{G}_{\theta}\in\mathbb{R}^{d\times d}$ denotes the Givens rotation, and $f(\mathbf{A},\mathbf{S},\mathbf{X};L)\in\mathbb{R}^{N\times d}$ is the aggregator of structural and attribute representations.
Proof.
Refer to Appendix II, Theorem V.3. ∎
Closed-form Expression. The construction above enjoys the existence of closed-form solution. In particular, the augmented graph is expressed as follows,
$\hat{\mathbf{A}}=\mathbf{P}\operatorname{diag}(\lambda_{1}^{\frac{1}{L}},\cdots,\lambda_{d}^{\frac{1}{L}})\mathbf{P}^{\top},$ (24)
where $\mathbf{P}$ and $\lambda$’s are derived as the eigenvectors and eigenvalues of the following matrix
We have $\mathbf{B}=f(\mathbf{A},\mathbf{S},\mathbf{X};L)\mathbf{G}_{\theta}^{T}$, and $\mathbf{U}=\mathbf{X}\mathbf{X}^{T}+\omega\mathbf{I}$ with penalty coefficient $\omega$. Note that, the eigenvalue decomposition is well-established with efficient methods [27] while the augmentation is typically applied as a preprocessing. The advantage is that we improve the expressiveness of the polynomial over structural and attribute matrices by incorporating representation consideration, as shown in Ablation Study, while recovering previous methods [25] as special cases.
VI LSEnet: Learning Partitioning Tree in Hyperbolic Space
We design a novel Lorentz Structural Entropy Neural Network (LSEnet), which aims to learn the optimal partitioning tree $\mathcal{T}_{\operatorname{net}}^{*}$ in the Lorentz model of hyperbolic space. LSEnet constructs the partitioning tree in a bottom-up fashion where we first embed leaf nodes of the tree, and then recursively learn the parent nodes, leveraging the clustering objective of differential structural information. In particular, we learn the partitioning trees from both the original and augmented structures, which serve as the different views in Lorentz tree contrastive learning (Sec. VII).
VI-A Why Hyperbolic Space?
First, we demonstrate the reason why we opt for the hyperbolic space, rather than the traditional Euclidean space. We introduce a quantitative metric of embedding distortion which is defined as $\frac{1}{|\mathcal{V}|^{2}}\sum_{ij}\left|\frac{d_{G}(v_{i},v_{j})}{d(\mathbf{x}_{i},\mathbf{x}_{j})}-1\right|$, where each node $v_{i}\in\mathcal{V}$ is embedded as $\mathbf{x}_{i}$ in representation space. $d_{G}$ and $d$ denote the distance in the graph and the space, respectively.
Theorem VI.1 (Tree $\mathcal{T}$ and Hyperbolic Space [28]).
$\forall\mathcal{T}$, scaling all edges by a constant so that the edge length is bounded below $\nu\frac{1+\epsilon}{\epsilon}$, there exists an embedding in hyperbolic space that the distortion overall node pairs are bounded by $1+\epsilon$. ($\nu$ is a constant detailed in Appendix V)
That is, hyperbolic space is well suited to embed the partitioning tree. Note that, Theorem VI.1 does not hold for Euclidean space or other Riemannian manifold.
VI-B Embedding Leaf Nodes
We design a Lorentz convolution layer to learn leaf embeddings in hyperbolic space, $\operatorname{LConv}:\boldsymbol{x}_{i}\to\boldsymbol{z}^{H}_{i},\forall v_{i}$, where $\boldsymbol{x}_{i}\in\mathbb{L}^{\kappa,d_{0}}$ and $\boldsymbol{z}^{H}_{i}\in\mathbb{L}^{\kappa,d_{\mathcal{T}}}$ are the node feature and leaf embedding, respectively. In the partitioning tree of height $H$, the level of nodes is denoted by superscript of $h$ and we have $h=H,\cdots,1,0$ from leaf nodes (bottom) to root (top), correspondingly.
We adopt attentional aggregation in $\operatorname{LConv}$ and specify the operations as follows. First, dimension transform is done via a Lorentz linear operator [21]. For node $\boldsymbol{x}_{i}\in\mathbb{L}^{\kappa,d_{0}}$, the linear operator is given as
$\operatorname{LLinear}(\boldsymbol{x})=\begin{bmatrix}\sqrt{\lVert h(\mathbf{W}\boldsymbol{x},\boldsymbol{v})\rVert^{2}-\frac{1}{\kappa}}\ \\ h(\mathbf{W}\boldsymbol{x},\boldsymbol{v})\ \end{bmatrix}\in\mathbb{L}^{\kappa,d_{\mathcal{T}}},\vspace{-0.03in}$ (26)
where $h$ is a neural network, and $\mathbf{W}$ and $\boldsymbol{v}$ are parameters. Second, we derive attentional weights from the self-attention mechanism. Concretely, the attentional weight $\omega_{ij}$ between nodes $i$ and $j$ is calculated as
$a_{ij}$ $=\operatorname{LeakyReLU}(\mathbf{W}[\boldsymbol{q}_{i}||\boldsymbol{k}_{j})])$ (27)
$\omega_{ij}$ $=\operatorname{LAtt}(\mathbf{Q},\mathbf{K})=\frac{\exp(a_{ij})}{\sum_{(i,l)\in\mathcal{E}}\exp(a_{il})},\vspace{-0.03in}$ (28)
where $\mathbf{W}\in\mathbb{R}^{1\times 2d}$ is learnable parameters, [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] is concatenate operation, $\boldsymbol{q}$ and $\boldsymbol{k}$ are the query and key vector collected in row vector of $\mathbf{Q}$ and $\mathbf{K}$, respectively. The queries and keys are derived from node feature $\boldsymbol{x}$ via $\operatorname{LLinear}$ with different parameters. Third, weighted aggregation is considered as the arithmetic mean among manifold-valued vectors. Given a set of points $\{\boldsymbol{x}_{i}\}_{i=1,\cdots,N}$ in the Lorentz model $\mathbb{L}^{\kappa,d_{\mathcal{T}}}$,
$\operatorname{LAgg}(\boldsymbol{\omega},\mathbf{X})=\frac{1}{\sqrt{-\kappa}}\sum_{i=1}^{N}\frac{\omega_{i}}{\lvert\lVert\sum_{j=1}^{N}\omega_{j}\boldsymbol{x}_{j}\rVert_{\mathbb{L}}\rvert}\boldsymbol{x}_{i},\vspace{-0.03in}$ (29)
where $\boldsymbol{\omega}$ is the weight vector, and $\boldsymbol{x}$ are summarized in $\mathbf{X}$. The augmented form is that for $\mathbf{\Omega}=[\boldsymbol{\omega}_{1},\cdots,\boldsymbol{\omega}_{N}]^{\top}$, we have $\operatorname{LAgg}(\mathbf{\Omega},\mathbf{X})=[\boldsymbol{\mu}_{1},\cdots,\boldsymbol{\mu}_{N}]^{\top}$, where the weighted mean is $\boldsymbol{\mu}_{i}=\operatorname{LAgg}(\boldsymbol{\omega}_{i},\mathbf{X})$. Overall, the Lorentz convolution layer is formulated as,
where $\odot$ is the Hadamard product, masking the attentional weight if the corresponding edge does not exist in the graph.
VI-C Learning Parent Nodes
A primary challenge is that, in the partitioning tree, the node number at each internal level is unknown. To address this issue, we introduce a simple yet effective method, setting a large enough node number $N_{h}$ at the $h$-th level. A large $N_{h}$ may introduce redundant nodes and result in a relaxed partitioning tree. According to Theorem IV.2, the redundant nodes in the partitioning tree do not affect the value of structural entropy, and finally present as empty leaf nodes by optimizing our objective. Theoretically, if an internal level has insufficient nodes, the self-organization of the graph can still be described by multiple levels in the partitioning tree.
[…]
$\mathbf{C}^{h}=\operatorname{LAtt}(\mathbf{Q},\mathbf{K})\sigma(\operatorname{MLP}(\mathbf{Z}^{h})),\vspace{-0.03in}$ (31)
where $\mathbf{A}^{h}$ is the graph structure at $h$-th level of the tree. $\sigma$ denotes the row-wise $Softmax$ function.
[…]
$\boldsymbol{z}^{h-1}_{j}=\arg_{\boldsymbol{z}^{h-1}_{j}}\min\sum\nolimits_{i=1}^{N}c_{ij}d^{2}_{\mathbb{L}}(\boldsymbol{z}^{h-1}_{j},\boldsymbol{z}^{h}_{i}).\vspace{-0.03in}$ (32)
Solving the optimization constrained in the manifold $\mathbb{L}^{\kappa,d_{\mathcal{T}}}$, we derive the closed-form solution of parent node embeddings,
$\mathbf{Z}^{h-1}=\operatorname{LAgg}(\mathbf{C}^{h},\mathbf{Z}^{h})\in\mathbb{L}^{\kappa,d_{\mathcal{T}}},$ (33)
which is the augmented form of Eq. (29). {mymath}
Theorem VI.2 (Geometric Centroid).
In Lorentz model $\mathbb{L}^{\kappa,d_{\mathcal{T}}}$ of hyperbolic space, for any set of points $\{\boldsymbol{z}^{h}_{i}\}$, the arithmetic mean is given as follows,
$\boldsymbol{z}^{h-1}_{j}=\frac{1}{\sqrt{-\kappa}}\sum_{i=1}^{N}\frac{c_{ji}}{\lvert\lVert\sum_{l=1}^{n}c_{jl}\boldsymbol{z}^{h}_{l}\rVert_{\mathbb{L}}\rvert}\boldsymbol{z}^{h}_{i}\in\mathbb{L}^{\kappa,d_{\mathcal{T}}},\vspace{-0.03in}$ (34)
and is the closed-form solution of the geometric centroid defined in the minimization of Eq. 32.
Remark. In fact, the geometric centroid in Theorem VI.2 is also equivalent to the gyro-midpoint in Poincaré ball model of hyperbolic space, detailed in Appendix II, Theorem VI.2.
[…]
$\mathbf{A}^{h-1}$ $=(\mathbf{C}^{h})^{\top}\mathbf{A}^{h}\mathbf{C}^{h}.\vspace{-0.03in}$ (35)
We recursively utilize Eqs. (31), (33) and (35) to build the partitioning tree from bottom to top in hyperbolic space.
VI-D Hyperbolic Partitioning Tree
In the principle of structural information minimization, the optimal partitioning tree is constructed to describe the self-organization of the graph [16, 18]. In the continuous realm, LSEnet learns the optimal partitioning tree in hyperbolic space by minimizing the objective in Eq. (7). DSI is applied on all the level-wise assignment $\mathbf{C}$’s, which are parameterized by neural networks in Sec. VI-B and VI-C. We suggest to place the tree root at the origin of $\mathbb{L}^{\kappa,d_{\mathcal{T}}}$, so that the learnt tree enjoys the symmetry of Lorentz model. Consequently, the hyperbolic partitioning tree describes the graph’s self-organization and clustering structure in light of Theorem III.5.
VII Graph Clustering with
Lorentz Tree Contrastive Learning
In this section, we introduce the Lorentz Tree Contrastive Learning designed to contrast trees in the hyperbolic space, achieving two key objectives. First, it incorporates attribute information in a contrastive manner to overcome the limitations of structural entropy. Second, it refines the hyperbolic tree representations, thereby enhancing clustering assignment. The main components of Lorentz Tree Contrastive Learning are i) Lorentz projection head, ii) hyperbolic reweighed sampling and iii) the Lorentz contrastive loss.
VII-A Neural Transformation in Lorentz Group
The projection head for contrasting partitioning trees differs significantly from typical graph contrastive learning. The key distinctions are as follows. First, existing methods primarily explore similarity in Euclidean space, whereas the partitioning tree is aligned with hyperbolic space, presenting a manifold-preserving challenge. In other words, the projection head is required to connect two hyperbolic spaces. Second, in graph contrastive learning, the projection head operates on graph nodes (leafs) in graph contrastive learning, while in the tree contrastive learning, it manipulates tree nodes across all levels simultaneously. Notably, maintaining consistency between successor and ancestor nodes is crucial owing to the unity of the tree, as supported by empirical evidence in Sec. VIII-E. However, ensuring consistency in hyperbolic space remains a significant challenge.
Before elaborating on the proposed projection head, we revisit the notion of Lorentz group $G^{L}=\{L_{i}\}$. To be specific, Lorentz group is a set of Lorentz transformations $L$, and $L_{i}\circ L_{j}\in G^{L}$ is closed on the group for any $L_{i},L_{j}\in G^{L}$. For any $\boldsymbol{x}\in\mathbb{L}^{\kappa,d_{1}}$, the equality of $L(\boldsymbol{x})=\mathbf{L}\boldsymbol{x}\in\mathbb{L}^{\kappa,d_{2}}$ holds for the block diagonal $\mathbf{L}=\left[\begin{array}[]{cc}w&\boldsymbol{v}^{\top}\\ \boldsymbol{v}&\mathbf{W}\end{array}\right],$ provided that $\mathbf{W}$ characterizes a Lorentz boost or Lorentz rotation. The advantage of Lorentz transformation lies in its clean matrix-vector multiplication form. However, the condition on $\mathbf{W}$ makes the optimization problematic; for instance, Lorentz rotation is restricted to the orthogonal cone of $\mathbf{W}$.
To bridge this gap, we introduce a novel Lorentz-type group consisting of learnable transformations that (i) connect different hyperbolic spaces (manifold-preserving) and (ii) ensure consistency for tree contrastive learning. The transformation is written as $L(\boldsymbol{x})=\mathbf{L}\boldsymbol{x}\in\mathbb{L}^{\kappa,d_{2}}$ for arbitrary $\boldsymbol{x}\in\mathbb{L}^{\kappa,d_{1}}$, where $d_{1}$ and $d_{2}$ are not necessarily the same. $\mathbf{L}$ takes the form of $\mathbf{L}:=\begin{bmatrix}\mathbf{v}^{T}\\ \mathbf{W}\end{bmatrix}$ satisfying the equation of
$\mathbf{W}^{T}\mathbf{W}=\mathbf{v}\mathbf{v}^{T}+diag(-1,1,...,1)\vspace{-0.05in}$ (36)
with $\mathbf{W}\in\mathbb{R}^{d_{2}\times d_{1}}$, and $d_{2}>d_{1}$ in the projection head. Next, we show the theoretical results on our construction. {mymath}
Theorem VII.1 (Manifold-preserving).
Given the construction of $\mathbf{L}$ in Eq. (36) and scalars $a,b$, the following properties holds for any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{L}^{d_{1}}$.
  • 1)
    $L(a\boldsymbol{x}+b\boldsymbol{y})=aL(\boldsymbol{x})+bL(\boldsymbol{y})$;
  • 2)
    $\langle L(\boldsymbol{x}),L(\boldsymbol{y})\rangle_{\mathbb{L}}=\langle\boldsymbol{x},\boldsymbol{y}\rangle_{\mathbb{L}}$,
where $\langle\boldsymbol{x},\boldsymbol{y}\rangle_{\mathbb{L}}$ denotes the Minkowski inner product in Eq. (6).
Proof.
Refer to Appendix II, Theorem VII.1. ∎
Note that, the manifold-preserving property is given by the claim (2) in the theorem above. It is easy to verify that $G$ is closed on itself, i.e., $L_{i}\circ L_{j}\in G^{L}$, $\forall L_{i},L_{j}\in G^{L}$. These properties also induce the following result. {mymath}
Theorem VII.2 (Consistence Insurance).
Given a set of points $\{\mathbf{z}_{i}\}$ in the Lorentz model of hyperbolic space, $\mathbf{z}_{i}\in\mathbb{L}^{\kappa,d_{1}}$ and their midpoint $\mathbf{z}\in\mathbb{L}^{\kappa,d_{1}}$ is constructed by Eq. (29), then $L(\mathbf{z})\in\mathbb{L}^{\kappa,d_{2}}$ is the midpoint of $\{L(\mathbf{z}_{i})\}$, $L(\mathbf{z}_{i})\in\mathbb{L}^{\kappa,d_{2}}$, if $L$ is a transformation in the Lorentz group as stated in Theorem VII.1.
Proof.
Refer to Appendix II, Theorem VII.2. ∎
The theorem above establishes the operation that propagates the action on tree nodes to their ancestors. In other words, only leaf contrast is needed in tree contrastive learning, and the consistence is maintained throughout the entire partitioning tree thanks to the properties of the proposed transformations.
Closed-form Expression. Another advantage is that Eq. (36) induces the closed-form expression. Specifically, we parameterize $\mathbf{L}$ on $\mathbf{v}$, and the corresponding $\mathbf{W}$ is derived as follows,
$\mathbf{W}=\begin{bmatrix}\mathbf{P}diag(\sqrt{\lambda_{1}},..,\sqrt{\lambda_{d+1}})\mathbf{P}^{T}\\ \mathbf{0}\end{bmatrix},\vspace{-0.05in}$ (37)
where $\mathbf{P}$ and $\lambda$’s are obtained from eigenvalue decomposition of $\mathbf{V}=\mathbf{v}\mathbf{v}^{T}+diag(-1,1,...,1)$ which takes the form of
$\mathbf{V}$ $=\mathbf{P}diag(\lambda_{1},..,\lambda_{d+1})\mathbf{P}^{T},\vspace{-0.05in}$ (38)
with eigenvalues $\lambda_{1},..,\lambda_{d+1}$ sorted in descending order. $\mathbf{0}$ is the padding matrix of zeros for dimensionality consistence.
Remark. We notice that recent studies propose a series of manifold-preserving operations in the Lorentz model of hyperbolic space [21, 31], but none of them satisfies consistence insurance and thereby cannot be utilized for the tree contrastive learning in hyperbolic space.
Algorithm 1 Training IsoSEL
0:  A weighted graph $G=(\mathcal{V},\mathcal{E},\mathbf{X})$ and the height of deep partitioning tree $H$.
0:  The partitioning tree $\mathcal{T}$, and tree node embeddings $\{\mathbf{Z}^{h}\}_{h=1,...,H}$ at all the levels.
1:  Generate the augmented graph using Eq. 25;
2:  while not converged do
3:     for both original and augmented graph do
4:        Obtain leaf node embeddings $\mathbf{Z}^{H}=\operatorname{LConv}(\mathbf{X},\mathbf{A})$;
5:        Compute the level-wise assignment and parent node embeddings recursively with Eqs. 31, 33 and 35;
6:     end for
7:     Forward leaf node embeddings into Lorentz projection head with Eq. 37;
8:     Compute the overall objective $\mathcal{L}_{\operatorname{deeptree}}+\alpha\mathcal{L}_{\operatorname{contrast}}$;
9:     Update the parameters with gradient-descent optimizer;
10:  end while
11:  Construct the optimal partitioning tree of root node $\lambda$ with a queue $\mathcal{Q}$;
12:  while $\mathcal{Q}$ is not empty do
13:     Get first item $\alpha$ in $\mathcal{Q}$;
14:     Let $h=\alpha.h+1$ and search subsets $P$ from $\mathbf{S}^{h}$ defined with level-wise assignments in Eq. 8;
15:     Create nodes from $P$ and put into the queue $\mathcal{Q}$;
16:     Add theses nodes into $\alpha$’s children list;
17:  end while
18:  Return the partitioning tree $\mathcal{T}:=\lambda$.
TABLE I: Clustering results on Cora, Citerseer, AMAP, and Computer, Pubmed datasets in terms of NMI and ARI (%). OOM denotes Out-of-Memory on our hardware. The best results are highlighted in boldface, and runner-ups are underlined.
Cora Citeseer AMAP Computer Pubmed
NMI ARI NMI ARI NMI ARI NMI ARI NMI ARI
DEC 23.54$\pm$1.13 15.13$\pm$0.72 28.34$\pm$0.42 28.12$\pm$0.24 37.35$\pm$0.84 18.29$\pm$0.64 38.56$\pm$0.24 34.76$\pm$0.35 22.44$\pm$0.14 19.55$\pm$0.13
VGAE 43.40$\pm$1.62 37.50$\pm$2.13 32.70$\pm$0.30 33.10$\pm$0.55 66.01$\pm$0.84 56.24$\pm$0.22 37.62$\pm$0.24 22.16$\pm$0.16 30.61$\pm$1.71 30.15$\pm$1.23
ARGA 48.10$\pm$0.45 44.10$\pm$0.28 35.10$\pm$0.58 34.60$\pm$0.48 58.36$\pm$1.02 44.18$\pm$0.85 37.21$\pm$0.58 26.28$\pm$1.38 24.80$\pm$0.17 24.35$\pm$0.17
MVGRL 62.91$\pm$0.42 56.96$\pm$0.74 46.96$\pm$0.25 44.97$\pm$0.57 36.89$\pm$1.31 18.77$\pm$1.54 10.12$\pm$2.21     5.53$\pm$1.78 31.59$\pm$1.45 29.42$\pm$1.06
Sublime 54.20$\pm$0.28 50.32$\pm$0.32 44.10$\pm$0.27 43.91$\pm$0.35     6.37$\pm$0.54     5.36$\pm$0.24 39.16$\pm$1.82 24.15$\pm$0.63 32.01$\pm$1.24 30.36$\pm$0.85
ProGCL 41.02$\pm$1.64 30.71$\pm$1.38 39.59$\pm$1.58 36.16$\pm$2.21 39.56$\pm$0.25 34.18$\pm$0.58 35.50$\pm$2.06 26.08$\pm$1.91 30.24$\pm$1.21 28.98$\pm$1.06
GRACE 57.30$\pm$0.86 52.70$\pm$1.20 39.90$\pm$2.26 37.70$\pm$1.35 53.46$\pm$1.32 42.74$\pm$1.57 47.90$\pm$1.65 36.40$\pm$1.56 30.08$\pm$0.24 27.60$\pm$1.24
MCGC 44.90$\pm$1.56 37.80$\pm$1.24 40.14$\pm$1.44 38.00$\pm$0.85 61.54$\pm$0.29 52.10$\pm$0.27 53.17$\pm$1.29 39.02$\pm$0.53 33.39$\pm$0.02 29.25$\pm$0.01
DCRN 48.86$\pm$0.85 43.79$\pm$0.48 45.86$\pm$0.35 47.64$\pm$0.30 73.70$\pm$1.24 63.69$\pm$0.84 OOM OOM 32.20$\pm$0.08 31.41$\pm$0.12
FT-VGAE 61.03$\pm$0.52 58.22$\pm$1.27 44.50$\pm$0.13 46.71$\pm$0.75 69.76$\pm$1.06 59.30$\pm$0.81 51.36$\pm$0.92 40.07$\pm$2.13 33.05$\pm$0.24 35.50$\pm$0.12
gCooL 58.33$\pm$0.24 56.87$\pm$1.03 47.29$\pm$0.10 46.78$\pm$1.51 63.21$\pm$0.09 52.40$\pm$0.11 47.42$\pm$1.76 27.71$\pm$2.28 30.82$\pm$0.15 31.48$\pm$0.34
S3GC 58.90$\pm$1.81 54.40$\pm$2.52 44.12$\pm$0.90 44.80$\pm$0.65 59.78$\pm$0.45 56.13$\pm$0.58 54.80$\pm$1.22 29.93$\pm$0.22 33.32$\pm$0.34 34.50$\pm$0.57
Congregate 63.16$\pm$0.71 59.27$\pm$1.23 50.92$\pm$1.58 47.59$\pm$1.60 70.99$\pm$0.67 60.55$\pm$1.36 46.03$\pm$0.47 38.57$\pm$1.05 35.03$\pm$0.23 33.42$\pm$0.47
DinkNet 62.28$\pm$0.24 61.61$\pm$0.90 45.87$\pm$0.24 46.96$\pm$0.30 74.36$\pm$1.24 68.40$\pm$1.37 39.54$\pm$1.52 33.87$\pm$0.12 36.39$\pm$0.34 35.77$\pm$0.13
GC-Flow 62.15$\pm$1.35 63.14$\pm$0.80 40.50$\pm$1.32 42.62$\pm$1.44 36.45$\pm$1.21 37.24$\pm$1.26 41.10$\pm$1.06 35.60$\pm$1.82 36.82$\pm$0.28 37.31$\pm$0.04
GCSEE 53.02$\pm$0.41 51.22$\pm$0.88 35.66$\pm$0.42 36.29$\pm$0.37 64.15$\pm$0.68 56.76$\pm$1.24 39.03$\pm$1.58 35.24$\pm$1.56 26.88$\pm$0.21 26.15$\pm$0.24
FastDGC 53.03$\pm$0.03 51.25$\pm$0.05 45.85$\pm$0.08 44.12$\pm$0.05 69.76$\pm$0.03 65.71$\pm$0.04 45.26$\pm$0.52 40.23$\pm$0.18 25.85$\pm$0.86 25.08$\pm$0.73
MAGI 59.72$\pm$1.22 57.31$\pm$1.31 45.28$\pm$0.24 46.86$\pm$0.47 71.62$\pm$0.11 61.55$\pm$0.24 54.22$\pm$1.88 46.27$\pm$1.64 33.28$\pm$0.04 31.29$\pm$0.08
RGC 57.60$\pm$1.36 50.46$\pm$1.72 45.70$\pm$0.29 45.47$\pm$0.43 47.65$\pm$0.91 42.65$\pm$1.53 46.24$\pm$1.05 36.12$\pm$0.30 28.01$\pm$1.85 25.56$\pm$1.91
DGCluster 61.47$\pm$0.34 44.75$\pm$0.28 43.75$\pm$0.05 31.55$\pm$0.09 77.12$\pm$0.25 67.59$\pm$0.35 50.10$\pm$1.87 38.92$\pm$1.27 33.73$\pm$0.68 21.84$\pm$0.71
IsoSEL 64.56$\pm$0.72 64.02$\pm$0.70 53.18$\pm$1.11 50.45$\pm$1.41 74.54$\pm$0.44 69.24$\pm$0.52 55.93$\pm$0.53 43.67$\pm$0.35 36.91$\pm$0.33 38.52$\pm$0.23
Avg. Gain 3.51 8.58 7.70 5.45 5.82 8.91 1.36 1.60 1.89 7.77
VII-B Hyperbolic Reweighing for Imbalance Clusters
In deep clustering on imbalanced graphs, minority clusters of fewer nodes are typically underrepresented due to the limited number of samples, which often results in a biased boundary among imbalanced clusters. To address this issue, in Lorentz tree contrastive learning, we propose hyperbolic reweighed sampling to mitigate the bias in clustering boundary. The core idea of reweighed sampling is to refine the hyperbolic representations of the majority clusters in the pushing-pulling process, as the structural entropy has a promising ability to identify minority clusters, as shown in Theorem IV.3.
To be specific, we first feed the original and augmented graphs into LSEnet to generate two hyperbolic views with their representations denoted as $\boldsymbol{z}$ and $\hat{\boldsymbol{z}}$, respectively. Second, we consider the contrast among the leaf nodes as the unity of the partitioning tree is ensured by the Lorentz projection head (Theorem VII.2). For each leaf node $\boldsymbol{z}^{H}_{i}$ in the original view, the positive sample is given as the corresponding $\hat{\boldsymbol{z}}^{H}_{i}$ in the augmented view. As for the negative samples, we draw $m$ leaf nodes of the same ancestor in the partitioning tree using a weighting function defined as follows,
$w_{ij}=\exp\left(-d_{\mathbb{L}}(L(\boldsymbol{z}^{H}_{i}),L(\hat{\boldsymbol{z}}^{H}_{j}))\right).\vspace{-0.05in}$ (39)
Accordingly, the Lorentz tree contrastive loss is formulated as
where $\tau$ is the temperature to control the strictness for negative samples, and $L$ denotes the Lorentz projection head. Note that, the reweighed sampling alleviates the over aggregation of majority cluster representations while preserving the closeness of similar nodes in hyperbolic space.
VII-C Overall Objective for Deep Graph Clustering
Finally, the overall objective of IsoSEL is formulated as
$\mathcal{J}=\mathcal{L}_{\operatorname{deeptree}}+\alpha\mathcal{L}_{\operatorname{contrast}},\vspace{-0.05in}$ (41)
where $\mathcal{L}_{\operatorname{deeptree}}$ is defined by Eq. (17), and $\alpha$ denotes the weighting coefficient. The training procedure is described in Algorithm 1, where Lorentz tree contrastive learning refines the deep partitioning tree created by differentiable structural information in hyperbolic space to generate promising node clusters. Specifically, we first learn the partitioning tree representation in a bottom-up fashion, as outlined in Lines $3$-$6$ of Algorithm 1. Then, we recover the tree through a top-down process in Lines $10$-$17$, ensuring that empty leaf nodes are avoided and single chains are fixed in the relaxation. Further details of the recovery procedure are provided in Algorithm 2 given in Appendix III.
Computational Complexity. In IsoSEL, the computational complexity of learning the deep partitioning tree is $\mathcal{O}(N)$ thanks to the node-wise formulation of differential structural entropy in Eq. (17). The Lorentz tree contrastive learning costs $\mathcal{O}(mN)$, where $N$ and $m$ denotes the number of nodes and negative samples, respectively. Accordingly, the overall complexity is yielded as $\mathcal{O}(N+mN)$, which is significantly less than that of our previous work of $\mathcal{O}(N|\mathcal{E}|)$ [20]. Note that, though the number of clusters $K$ is no longer required, IsoSEL presents competitive complexity to typical deep clustering models of $\mathcal{O}(KN)$.
In summary, IsoSEL takes the advantages of both hyperbolic geometry and differential structural entropy to effectively discover the cluster structure without $K$, while alleviating the clustering bias among imbalanced clusters.
VIII Experiments
In the experiment, we compare with $20$ strong baselines on $5$ commonly used datasets, aiming to answer the research questions as follows: RQ1 (Main Results): How does the proposed IsoSEL perform in terms of both effectiveness and efficiency? RQ2: How does IsoSEL outperform the classic structural entropy [16] and our previous model [20]? RQ3 (Ablation Study): How does each proposed module contribute to the success of IsoSEL, including the choice of hyperbolic space, augmentation strategy and tree contrastive learning? RQ4 (Hyperparameter Sensitivity): How do the hyperparameters affect the performance? RQ5: How do the imbalance clusters impact the clustering results? Besides, we examine the embedding expressiveness of IsoSEL, and include a visualization of the clusters learnt from IsoSEL. Additional results and further details are provided in Appendix VII. Codes and data are publicly available at https://github.com/RiemannGraph/DSE_clustering.
VIII-A Experimental Setups
VIII-A1 Datasets & Baselines
The experiments are conducted on a diversity of datasets including Cora, Citeseer, Pubmed [14], Amozon Photo, and a larger Computer dataset [32]. In this work, we focus on deep graph clustering and compare with $20$ strong deep baselines categorized into two groups. The first group includes a classic DEC [33] and $14$ state-of-the-art deep graph clustering models: GRACE [34], MCGC [35], DCRN [36], FT-VGAE [37], gCooL [38], S3GC [5], Congregate [39], DinkNet [40], GC-Flow [6], RGC [8], GCSEE [41], FastDGC [42], MAGI [43] and DGCluster [44]. The second group consists of $5$ self-supervised graph models: VGAE [45], ARGA [46], MVGRL [47], Sublime [25] and ProGCL [48]. The embedding, learned from the self-supervised models, is then followed up with $K$-Means algorithm to generate node clusters. Among the baselines, Congregate is the only Riemannian baseline that considers the curvature heterogeneity underlying the graph itself, while we explore the hyperbolic geometry of the partitioning tree. We specify that previous deep clustering models cannot be instantiated without the cluster number $K$, except RGC and DGCluster. Distinguishing from the previous models, we are dedicated to uncovering the cluster structure in hyperbolic space from the information theory perspective. (Detailed dataset and baseline descriptions are given in Appendix VI.)
TABLE II: Accuracy results on Cora, Citeseer, AMAP, Computer, and Pubmed datasets in terms of percentage (%). The best results are highlighted in boldface, and runner-ups are underlined.
VGAE ARGA MVGRL Sublime GRACE DCRN gCooL S3GC DinkNet GC-Flow DGCluster IsoSEL Avg. Gain
Cora 64.72 68.40 80.03 71.30 73.87 66.91 77.03 74.27 78.11 78.92 76.45 81.77 7.22
Citeseer 61.00 60.72 61.23 68.53 63.12 70.86 71.35 68.80 70.36 70.06 70.14 71.37 6.26
AMAP 74.26 69.28 41.07 47.22 67.66 77.94 68.52 75.15 79.23 72.54 78.92 78.68 7.51
Computer 42.44 45.67 29.53 45.52 46.62 30.42 44.50 39.03 43.52 43.20 43.65 47.34 5.14
Pubmed 68.31 64.82 67.01 66.62 63.72 69.63 68.24 71.30 69.42 73.42 66.39 73.53 6.26
TABLE III: Comparison of time complexity and time cost for baselines.
Method Time Complexity Inference Time (s)
DGI $\mathcal{O}(ED+Nd^{2})$ 20.15
MVGRL $\mathcal{O}(N^{2}d+Nd^{2})$ 176.35
S3GC $\mathcal{O}(NSd^{2})$ 520.41
Dink-Net $\mathcal{O}(BKd+K^{2}d+Bd)$ 37.32
DGCluster $\mathcal{O}(K^{2}N+P^{2}N)$ 8.56
IsoSEL (Ours) $\mathcal{O}(N+mN)$ 7.53
VIII-A2 Evaluation Metrics
We evaluate the node clustering results with three popular metrics: Normalized Mutual Information (NMI) [47], Adjusted Rand Index (ARI) [49] and Accuracy (ACC) [47, 38]. Concretely, ACC quantifies the accuracy of predicted cluster assignment. NMI evaluates the amount of information shared between predicted and true distribution. ARI considers the correct and wrong assignment. In the experiment, each case undergoes $10$ independent runs to ensure statistical robustness and fair comparison. We report the mean value along with the standard deviation. All experiments are done on a server of an NVIDIA GeForce RTX 4090D GPU (24GB VRAM) utilizing CUDA 11.8, an AMD EPYC 9754 128-Core Processor (18 vCPUs allocated), 60GB of RAM.
VIII-A3 Model Configuration & Reproducibility
We instantiate the proposed IsoSEL in the Lorentz model of hyperbolic space with a standard curvature of $\kappa=-1$. The model configuration by default is specified as follows. We consider the partitioning tree of height $2$ by default and hyperparameter $N_{1}$ of network architecture is set as $10$. The augmentation ratio $\alpha$ is $0.2$. We stack the Lorentz convolution once in LSEnet, and set the number of negative samples to $30$ in Lorentz tree contrastive learning. The parameters are optimized by Adam and further details are provided in implementation note of Appendix VI. As for the baselines, the hyperparameter setting is the same as that of original papers.
VIII-B Main Results of Node Clustering
VIII-B1 Effectiveness
We report the clustering results in terms of NMI and ARI on the five real datasets in Table I, and ACC in Table II. Deep clustering models require cluster number $K$ in training phase, except DGCluster, RGC and our IsoSEL. For clustering models without $K$, DGCluster conducts self-supervised representation learning with modularity [44], and RGC searches the optimal $K$ with reinforcement learning [8]. For Riemannian models, i.e., Congregate and IsoSEL, we project Euclidean attributes onto the respective manifold, and the projection is given by exponential map with the reference point of the origin. DCRN does not scale well to the larger Computer dataset, and runs into Out-Of-Memory (OOM) error on our hardware. In Table I, the proposed IsoSEL has the best clustering results except two cases, NMI on AMAP and ARI on Computer. Note that, IsoSEL outperforms $14$ recent deep clustering models by $+1.3\%$ NMI on average, and generally achieves the best ACC in Table II. Also, we observe that IsoSEL shows more consistent performance over different datasets, e.g., DGCluster obtain the best NMI on AMAP but is unsatisfactory on Computer.
VIII-B2 Efficiency
Table III collects the theoretical time complexity and empirical inference time on Cora. The inference time is given in seconds, and we are presented as the fastest deep model in the table. As for theoretical complexity, $N$ and $E$ denote the number of nodes and edges, respectively. $B$ is the batch size of DinkNet. $S$ is the average degree of nodes and thereby the complexity of S3GC is related to the number of edges. In DGCluster, $P$ is a dimensionality of auxiliary matrix, $P\ll N$. In IsoSEL, $m$ is the number of negative samples, and it yields the linear complexity thanks to the reformulation of differential structural information and hyperbolic reweighing of Lorentz tree contrastive learning. It is noteworthy to mention that, compared to recent deep clustering model such as S3GC, we achieve superior clustering effectiveness while removing the prior knowledge of cluster number without computational overhead. The following parts will demonstrate how the differential formulation, hyperbolic space and Lorentz tree contrastive learning contribute to the success of the proposed model.
TABLE IV: Ablation study results of different variants of IsoSEL on graph clustering tasks in terms of clustering performance (%).
Cora Citeseer AMAP Photo Pubmed
NMI ARI NMI ARI NMI ARI NMI ARI NMI ARI
LSEnet 63.97$\pm$0.67 63.35$\pm$0.56 52.26$\pm$1.09 48.01$\pm$1.25 71.12$\pm$1.30 65.08$\pm$0.73 55.03$\pm$0.79 42.15$\pm$1.02 34.25$\pm$0.24 35.16$\pm$0.30
w/o CIncorp 63.45$\pm$0.92 62.78$\pm$0.88 51.96$\pm$1.45 48.27$\pm$1.65 72.65$\pm$1.92 64.37$\pm$1.35 54.78$\pm$1.03 42.86$\pm$1.62 36.23$\pm$0.50 37.12$\pm$0.69
w/o IsoAug 62.89$\pm$1.05 62.15$\pm$0.97 51.42$\pm$1.58 47.93$\pm$1.80 72.23$\pm$2.01 64.01$\pm$1.50 54.34$\pm$1.12 42.54$\pm$1.79 35.97$\pm$0.64 36.85$\pm$0.82
w/o TRefine 62.34$\pm$1.20 61.78$\pm$1.10 50.87$\pm$1.70 47.42$\pm$1.95 71.89$\pm$2.14 63.78$\pm$1.68 53.98$\pm$1.25 42.21$\pm$1.98 35.78$\pm$0.81 36.64$\pm$0.95
w/o LGroup 61.78$\pm$1.32 61.23$\pm$1.18 50.31$\pm$1.85 46.89$\pm$2.10 71.54$\pm$2.28 63.45$\pm$1.80 53.56$\pm$1.38 41.89$\pm$2.12 35.56$\pm$0.95 36.42$\pm$1.12
IsoSEL 64.56$\pm$0.72 64.02$\pm$0.70 53.18$\pm$1.11 50.45$\pm$1.41 74.54$\pm$0.44 69.24$\pm$0.52 55.93$\pm$0.53 43.67$\pm$0.35 36.91$\pm$0.33 38.52$\pm$0.23
TABLE V: Comparison between LSEnet and CSE on Cora, Citerseer, Football and Karate datasets in terms of NMI and ARI (%).
Cora Citeseer Football Karate
NMI ARI NMI ARI NMI ARI NMI ARI
CSE 60.82$\pm$0.30 59.36$\pm$0.17 47.12$\pm$1.01 45.67$\pm$0.67 79.23$\pm$0.12 54.06$\pm$0.05 81.92$\pm$0.29 69.77$\pm$0.02
IsoSEL 64.56$\pm$0.72 64.02$\pm$0.70 53.18$\pm$1.11 50.45$\pm$1.41 82.35$\pm$0.53 57.36$\pm$0.35 83.28$\pm$0.72 74.15$\pm$0.55
Performance Gain 3.74 4.66 6.06 4.78 3.12 3.30 1.36 4.38
Refer to captionRefer to caption
Figure 2: Parameter sensitivity on dimension of structural entropy.
VIII-C Comparison with Classic Structural Entropy [16]
In this section, we analyze the strength of the proposed IsoSEL against the classic structural entropy [16] (referred to as CSE for short). (1) Time Complexity. We report the running time under different heights of partitioning trees in Fig 2. IsoSEL shows better scalability than CSE. Concretely, CSE has competitive time cost to IsoSEL on the small dataset (Football), but performs undesirably on larger datasets, e.g., at least 13×13$\times$13 × time cost to IsoSEL on Cora. In addition, the linear complexity of IsoSEL is regardless of the height of partitioning tree, while CSE scales exponentially over the height, e.g., $\mathcal{O}(N^{3}\log^{2}N)$ for height $3$, making it prohibited on large graphs. (2) Clustering Effectiveness. Note that, CSE itself neither generates cluster assignment nor node representations. We invoke the heuristic algorithm to construct a partitioning tree of height $2$, where the leaf-parent relationship is regarded as cluster assignment, and thus we instantiate IsoSEL to generate a deep tree of the same height for fair comparison. We summarize the results on Cora, Citeseer, Football and Karate in Table V. IsoSEL consistently outperforms CSE for all the cases. It suggests that cluster assignment is improved by the hyperbolic geometry and attribute incorporation of IsoSEL.
VIII-D Comparison with Our Previous LSEent [20]
This subsection demonstrates the significant improvement compared to LSEnet in the conference version. LSEnet constructs the deep partitioning tree in hyperbolic space, while IsoSEL leverages Lorentz tree contrastive learning to refine the deep tree of LSEnet. First, we compare the clustering effectiveness and report the results on Cora, Citeseer, AMAP, Photo and Pubmed datasets in Table IV. IsoSEL consistently achieves better performance in both NMI and ARI. The performance gain is attributed to hyperbolic tree representation refinement and attribute incorporation via Lorentz tree contrastive learning. Second, we examine the running time under different tree heights in Fig. 2. IsoSEL outperforms LSEnet on both datasets owing to the reduced complexity in the proposed reformulation and the sampling strategy of contrastive learning. In other words, IsoSEL in this paper exhibits superior clustering ability to the previous LSEnet, with minimal computational overhead. (The superiority in expressiveness and imbalanced clusters is elaborated in Sec. VIII-G and Sec. VIII-H, respectively.)
TABLE VI: Ablation study results of different representation space of partitioning tree of IsoSEL.
Cora Citeseer Pubmed
NMI ARI NMI ARI NMI ARI
Hyperbolic 64.56 64.02 53.18 50.45 36.91 38.52
Hyperspherical 50.53 51.23 43.53 40.09 27.35 27.26
Euclidean 56.33 53.78 48.92 46.42 30.07 30.19
VIII-E Ablation Study
This subsection introduce several variants to examine the effectiveness of each proposed module of IsoSEL, and comparison results are summarized in Table IV.
VIII-E1 Contrastive Incorporation vs Input Incorporation
We discuss the strategy for incorporating attribute information into partitioning tree. In IsoSEL, we propose a contrastive approach that incorporates attribute information through mutual information maximization between the original and augmented graphs. Correspondingly, a variant named $\operatorname{w/oCIncorp}$ is introduced to perform input incorporation by directly constructing the partitioning tree from the augmented graph with node attributes. As shown in Table IV, IsoSEL consistently outperforms $\operatorname{w/oCIncorp}$ in all cases. The superiority of contrastive incorporation lies in its ability to not only inject the attribute information but also to refine the tree representations in hyperbolic space, thereby improving cluster assignment.
VIII-E2 Isometric Augmentation vs Weighted-Sum Augmentation
We study how to effectively integrate node attributes into graph structure in the augmented view. In IsoSEL, we derive the isometric augmentation in Eq. (24) which conducts the integration from the representation perspective with the notion of isometry. To evaluate the effectiveness of the proposed formulation, we introduce the $\operatorname{w/oIsoAug}$ variant that generates the augmented structure by a weighted sum of original structure and attribute similarity. In Table IV, we find that the weighted-sum augmentation of $\operatorname{w/oIsoAug}$ leads to inferior performance, demonstrating the effectiveness of isometric augmentation.
VIII-E3 Representation Refinement by Tree Contrastive Learning
This part verifies that contrastive learning itself contributes to the success of clustering. To isolate the impact of attribute information, we introduce the $\operatorname{w/oTRefine}$variant, which is a degenerated version of IsoSEL. Specifically, we replace isometric augmentation with random structure perturbation, where the augmented view is generated by randomly adding or removing edges from the original structure. In Table IV, we find that $\operatorname{w/oTRefine}$ generally achieves a performance gain compared to LSEnet. The contrastive learning refines the partitioning tree representation in hyperbolic space, which is important to clustering performance.
VIII-E4 Neural Lorentz Group for Tree Contrastive Learning
To investigate its effectiveness, we design a variant named $\operatorname{w/oLGroup}$, which replaces proposed Lorentz projection head of neural Lorentz group by the Lorentz layer in [21]. Specifically, the Lorentz layer preserves the manifold structure in the Lorentz model but lacks the linearity of the proposed transformation. We emphasize that linearity ensures consistency between leaf and parent nodes in the partitioning tree, as elaborated in Theorem VII.2. In Table IV, IsoSEL achieves better clustering results than $\operatorname{w/oLGroup}$, demonstrating the importance of consistency in tree contrastive learning.
VIII-E5 Representation Space of Partitioning Tree
We conduct a geometric ablation to study the alignment between hyperbolic space and the partitioning tree. Note that, we are interested in the promising space of the deep tree, while the underlying manifold of input graph is orthogonal to our focus. Theoretically, Theorem VI.1 shows the connection between the tree and hyperbolic space. Empirically, we instantiate the proposed IsoSEL in hyperbolic, hyperspherical and Euclidean space without loss of generality. The clustering results on Cora, Citeseer and Pubmed are given in Table VI. These results show that hyperbolic space is well suited to model the partitioning tree and achieves the best performance among constant curvature spaces.
Figure 3: Sensitivity Analysis of $\alpha$
(a) NMI
(b) ARI
Figure 4: Parameter sensitivity on dimension of structural entropy.
VIII-F Hyperparameter Sensitivity
In this subsection, we discuss the hyperparameter sensitivity of the proposed IsoSEL: attribute augmentation ratio $\alpha$ and height of partitioning tree $H$.
VIII-F1 Attribute Augmentation Ratio
In the proposed isometric augmentation, the hyperparameter $\alpha$ controls the relative importance over attribute and structural information. To evaluate the sensitivity, we vary $\alpha$ in the range of $[0.2,0.8]$, while fixing other hyperparameters. We report the corresponding performance on Cora, Citeseer, AMAP and Computer datasets in Fig. 3. Note that, only slight changes of clustering results is observed under different ratios, and IsoSEL exhibits robustness regarding $\alpha$.
VIII-F2 Height of Partitioning Tree
The hyperparameter $H$ determines the height of the partitioning tree. We collect the clustering results with different tree heights in Fig 4, where $H$ takes the values in $\{2,3,4,5\}$. There are two findings from the empirical results: First, the height of partitioning tree affects the performance of IsoSEL. Second, we notice that IsoSEL achieves the best results on Cora at the height of $2$ while the optimal height of AMAP is $3$. We argue that the optimal height is depend on the inherent complexity of the graph.
VIII-G Embedding Expressiveness of IsoSEL
IsoSEL generates cluster assignment as well as node embedding in hyperbolic space. Here, we evaluate the embedding expressiveness with link prediction. To this end, IsoSEL is compared with the Euclidean models (including GCN [3], SAGE [50], GAT [4] and DGI [51]), hyperbolic models (including HGCN [52] and LGCN [53]), a recent $Q$GCN in ultra hyperbolic space [54] and our previous LSEnet. Popular metrics of AUC (Area Under the Curve) and AP (Average Precision) are employed to quantify the results. The link prediction performances on the five datasets are collected in Table VII, where the height of partitioning tree is set as $3$ for both LSEnet and IsoSEL. As shown in Table IV, Riemannian models generally outperforms the Euclidean counterparts, and IsoSEL achieves the state-of-the-art results. It suggests that Riemannian manifolds better model the complexity of graph structures and, more importantly, the partitioning tree encodes fruitful information for link prediction. Also, compared to LSEnet, IsoSEL consistently obtains the performance gain thanks to the proposed Lorentz tree contrastive learning.
TABLE VII: Embedding expressiveness regarding link prediction on Cora, Citeseer, AMAP, Computer, and Pubmed datasets. The best results are highlighted in boldface, and runner-ups are underlined.
Cora Citeseer AMAP Computer Pubmed
Model AUC AP AUC AP AUC AP AUC AP AUC AP
GCN 91.19$\pm$0.51 91.85$\pm$0.43 90.16$\pm$0.49 91.91$\pm$0.24 90.12$\pm$0.72 91.08$\pm$0.08 93.86$\pm$0.36 93.24$\pm$0.32 91.16$\pm$0.14 89.96$\pm$0.21
SAGE 86.02$\pm$0.55 88.24$\pm$0.87 88.18$\pm$0.22 89.15$\pm$0.36 98.02$\pm$0.32 97.51$\pm$0.21 92.82$\pm$0.20 93.12$\pm$0.14 89.22$\pm$0.87 89.44$\pm$0.82
GAT 92.55$\pm$0.49 90.27$\pm$0.23 89.32$\pm$0.36 90.08$\pm$0.06 98.67$\pm$0.08 97.21$\pm$0.34 95.93$\pm$0.19 94.57$\pm$0.24 91.60$\pm$0.37 90.37$\pm$0.21
DGI 90.02$\pm$0.80 90.61$\pm$1.00 94.53$\pm$0.40 95.72$\pm$0.10 98.24$\pm$0.13 96.84$\pm$0.24 95.84$\pm$0.05 95.22$\pm$0.36 91.24$\pm$0.60 92.23$\pm$0.50
HGCN 93.60$\pm$0.37 93.64$\pm$0.12 94.39$\pm$0.42 94.38$\pm$0.11 98.06$\pm$0.29 96.84$\pm$0.15 96.88$\pm$0.53 96.18$\pm$0.13 96.30$\pm$0.28 94.72$\pm$0.52
LGCN 92.69$\pm$0.26 91.09$\pm$0.37 93.49$\pm$1.11 94.96$\pm$0.18 97.08$\pm$0.08 96.73$\pm$0.52 96.37$\pm$0.70 95.66$\pm$0.24 95.40$\pm$0.23 93.12$\pm$0.24
$Q$GCN 95.22$\pm$0.29 94.12$\pm$0.51 94.31$\pm$0.73 94.08$\pm$0.34 95.17$\pm$0.45 93.94$\pm$0.18 95.10$\pm$0.03 94.28$\pm$0.33 96.20$\pm$0.34 94.24$\pm$0.31
LSEnet 95.51$\pm$0.60 95.40$\pm$0.04 94.42$\pm$1.51 94.69$\pm$0.05 98.75$\pm$0.67 97.04$\pm$0.13 97.06$\pm$1.02 97.06$\pm$0.05 96.41$\pm$0.10 95.46$\pm$0.10
IsoSEL 96.23$\pm$0.14 95.94$\pm$0.21 95.02$\pm$0.34 95.81$\pm$0.29 98.78$\pm$0.19 97.26$\pm$0.17 97.15$\pm$0.34 97.34$\pm$0.38 96.50$\pm$0.07 95.72$\pm$0.09
(a) gCool: true
(b) GDCluster: true
(c) RGC: true
(d) LSEnet: true
(e) IsoSEL: true
(f) gCool: prediction
(g) GDCluster: prediction
(h) RGC: prediction
(i) LSEnet: prediction
(j) IsoSEL:prediction
Figure 5: Visualization of real and predicted data.
VIII-H Case Study on Imbalanced Clusters
We examine the clustering boundaries among imbalanced clusters. Real graphs often exhibit skewed class distributions, and for this case study, we use the Photo dataset. The proportion of $8$ clusters in Photo is $25\%$, $22\%$, $11\%$, $11\%$, $10\%$, $9\%$, $4\%$ and $4\%$. We compare IsoSEL with gCool, DGCluster, RGC and LSEnet, and visualize the node embedding in Fig. 5 where different clusters are denoted by different colors. The distributions of true labels and predicted labels are given in the top and bottom subfigures, respectively. We invoke the normal t-sne for gCool, DGCluster, RGC, while employing the hyperbolic metric for LSEnet and IsoSEL, so that all the visualizations are finally presented in 2D Euclidean space. In Fig. 5, we obtain three findings as follows. (i) Clustering model without $K$ may fail to uncover the correct cluster number, e.g., RGC outputs $7$ node clusters on Photo dataset, but IsoSEL gives the true cluster number $8$. (ii) We observe that deep clustering models tend to generate biased clustering boundary. For example, in RGC, the members of violet cluster and the minority gray cluster (at the upper left in Fig. 5c) are assimilated into the majority cluster colored in purple, comparing the true label in Fig. 5h. The reason lies in that the deep models typically perform clustering in the representation space and minority clusters are underrepresented in the self-supervised learning owing to limited samples. (iii) Different from other deep models, the minority clusters are easier to be identified and discovered in both LSEnet and IsoSEL. We highlight that, the differentiable structural information objective effectively discovers node clusters without $K$ and is suitable to identify minority clusters in imbalanced graphs. In addition, compared to LSEnet, IsoSEL further mitigates the bias in clustering boundary with the hyperbolic reweighing proposed in Lorentz tree contrastive learning.
(a) Karate: Results
(b) Football: Results
(c) Karate: Groundtruth
(d) Football: Groundtruth
Figure 6: Visualization of hyperbolic partitioning trees.
VIII-I Visualizing Hyperbolic Partitioning Tree
We visualize the hyperbolic partitioning trees of IsoSEL. In IsoSEL, the partitioning tree is created in the Lorentz model of hyperbolic space. Note the fact that the 2D Lorentz model is represented in 3D Euclidean space and is hard to be visualized. To bridge this gap, we perform the stereographic projection $\Psi$ [55] on the Lorentz model and obtain the equivalent Poincaré disc $\mathbb{B}^{2}$, a 2D open disc that is more preferable for visualization. Given the standard curvature $\kappa=-1$, for any point in the Lorentz model $\boldsymbol{x}=[x_{1}\ \boldsymbol{x}_{s}]^{\top}\in\mathbb{L}^{2}$, we derive the corresponding point in Poincaré disc via $\Psi(\boldsymbol{x})=\frac{\boldsymbol{x}_{s}}{x_{1}+1}\in\mathbb{B}^{2}$. With $\Psi$, we study the clustering results on two toy examples of Karate and Football datasets, and show the visualization in the 2D Poincaré disc in Fig. 6. The partitioning tree with learnt clusters are plotted in Fig. 6 (a) and (b), while true labels are given in Fig. 6 (c) and (d). Two important findings are summarized as follows: (i) The clustering results of IsoSEL are generally aligned with the ground truth distribution. (ii) The learnt embeddings of each clusters are well separated, showing the effectiveness of IsoSEL.
IX Related Work
IX-A Deep Graph Clustering
In recent years, deep graph clustering achieves remarkable success thanks to the expressive representation learning. [^3] In this paper we study node clustering on the graph, and graph-level clustering is out of our scope. Typically, graph clustering requires the cluster number $K$ as the prior knowledge, and existing models can be roughly categorized into three groups. (i) Reconstructive methods learn node representation by recovering graph information [41], while generating node clusters via distance-based or model-based algorithms (e.g., k-means, Gaussian mixture model) [33, 56, 34]. (ii) Adversarial methods acquire the clustering results through the minmax game between generator and discriminator [57, 1, 42]. (iii) Contrastive methods obtain informative representation by pulling positive samples together while pushing negative samples apart [5, 35, 38, 43]. Recently, generative models, e.g., normalizing flows have been explored [6]. Self-supervised objective regarding Ricci curvature is designed for node clustering [39]. In the literature, research on deep graph clustering without $K$ is still limited. For instance, DGCluster generates node encodings by optimizing modularity and leverages Brich algorithm to obtain clustering results [44]. RGC automates the training-testing process to search the optimal $K$ via reinforcement learning [8]. Both of them have different focus from ours, and we aim at uncovering the inherent cluster structure from the information theory perspective. In parallel, efforts have been made to address the imbalance in graphs [12, 13], given that real-world graphs often exhibit skewed class distributions. Although resampling and reweighing techniques have been designed for imbalanced graphs [15, 58, 59], the clustering objective for handling imbalance clusters remains underexplored. Our work is dedicated to bridging this gap.
IX-B Structural Entropy
In the information theory, Shannon entropy measures the amount of information for unstructured data [60], but the information measure over graphs remains open. In the literature, efforts have been made to study the graph information measure. Von Neumann entropy [61] and Gibbs entropy [62] fall short in modeling the graph structure, while structural entropy addresses the absence of structural information and constructs a partitioning tree to reveal the self-organizing of graphs [16]. Recently, it has been successfully applied to biochemical analysis [63], dimensionality estimation [64], graph classification [65], adversarial attack [17], structure augmentation [19] and social event detection [66]. Theoretical advances include generalization to incremental setting [67] and incorporation with mutual information [68]. We observe that structural entropy is expressed as a discrete formula and its optimal partitioning tree has thus far been constructed using the heuristic algorithm [65], leading to high computational complexity that is impractical for real graphs. To address these challenges, this paper introduces the differential structural information, and theoretically demonstrates its effectiveness in clustering without requiring K, as well as in identifying minority clusters in imbalanced graphs. Beyond graph clustering, our work paves the way for learning structural information through gradient-based optimization.
IX-C Riemannian Deep Learning on Graphs
Euclidean space has been the workhorse of graph learning for decades [3, 4], while Riemannian manifolds have emerged as an exciting alternative in recent years. In Riemannian geometry, a structure is related to certain manifold with respect to its geometric property. Among Riemannian manifolds, hyperbolic space is well recognized for its alignment to hierarchical/tree-like structures. A series of hyperbolic GNNs (e.g., HGNN [69], HGCN [52]) show superior performance to their Euclidean counterparts. In the literature, hyperbolic manifold routinely find itself in graph representation learning, recommender system, knowledge graphs, etc., In fact, an arbitrary tree can be embedded in hyperbolic space with bounded distortion [28], but the connection between hyperbolic manifold and structural entropy has not yet been established, to the best of our knowledge. Beyond hyperbolic manifold, recent studies visit hyperspherical manifold [70, 71], gyrovector spaces [55, 72], product spaces [73, 74], quotient spaces [75, 54], SPD manifolds [76], as well as the notion of Ricci curvature for learning on graphs [77, 78]. However, previous Riemannian models study the underlying manifold of the complex graph structures, while we primarily focus on the hyperbolic space to create and refine the partitioning tree for graph clustering.
X Conclusion
Our work bridges structural information with deep learning and provides a principled solution for graph clustering without without requiring a predefined number of clusters while accounting for the imbalance in real-world graphs. Theoretically, we establish differentiable structural information in continuous realm, and prove its effectiveness in clustering without $K$ and detecting minority clusters. Accordingly, we present the IsoSEL framework where LSEnet is designed to learn a deep partitioning tree via gradient backpropagation in the Lorentz model of hyperbolic space. Furthermore, we propose Lorentz tree contrastive learning with isometric augmentation to refine the deep tree, thereby enhancing cluster assignment. Extensive experiments demonstrate that IsoSEL effectively generates node clusters and mitigates the bias in the clustering boundary among imbalanced clusters.

### v2

@@
-IsoSEL: Isometric Structural Entropy Learning for Deep Graph Clustering in Hyperbolic Space
+ASIL: Augmented Structural Information Learning for Deep Graph Clustering in Hyperbolic Space
 Abstract
@@
 In the literature, structural information has rarely been touched in deep clustering, and the classic definition falls short in its discrete formulation, neglecting node attributes and exhibiting prohibitive complexity.
-In this paper, we first establish a new Differentiable Structural Information, generalizing the discrete formalism to continuous realm, so that the optimal partitioning tree, revealing the cluster structure, can be created by the gradient backpropagation.
-Theoretically, we demonstrate its capability in clustering without requiring $K$ and identifying the minority clusters in imbalanced graphs, while reducing the time complexity to $O(N)$ w.r.t.
-the number of nodes.
-Subsequently, we present a novel IsoSEL framework for deep graph clustering, where we design a hyperbolic neural network to learn the partitioning tree in the Lorentz model of hyperbolic space, and further conduct Lorentz Tree Contrastive Learning with isometric augmentation.
-As a result, the partitioning tree incorporates node attributes via mutual information maximization, while the cluster assignment is refined by the proposed tree contrastive learning.
-Extensive experiments on five benchmark datasets show the IsoSEL outperforms $14$ recent baselines by an average of $+1.3\%$ in NMI.
-Index Terms:
-Graph Clustering, Deep Learning, Structural Entropy, Hyperbolic Space, Lorentz Group.
+In this paper, we first establish a differentiable structural information, generalizing the discrete formalism to continuous realm, so that we design a hyperbolic deep model (LSEnet) to learn the neural partitioning tree in the Lorentz model of hyperbolic space.
+Theoretically, we demonstrate its capability in clustering without requiring $K$ and identifying minority clusters in imbalanced graphs.
+Second, we refine hyperbolic representations of the partitioning tree, enhancing graph semantics, for better clustering.
+Contrastive learning for tree structures is non-trivial and costs quadratic complexity.
+Instead, we further advance our theory by discovering an interesting fact that structural entropy indeed bounds the tree contrastive loss.
+Finally, with an efficient reformulation, we approach graph clustering through a novel augmented structural information learning (ASIL), which offers a simple yet effective objective of augmented structural entropy to seamlessly integrates hyperbolic partitioning tree construction and contrastive learning.
+With a provable improvement in graph conductance, ASIL achieves effective debiased graph clustering in linear complexity with respect to the graph size.
+Extensive experiments show the ASIL outperforms $20$ strong baselines by an average of $12.42\%$ in NMI on Citeseer dataset.
 Codeshttps://github.com/RiemannGraph/DSE_clustering
@@
 While achieving remarkable success, deep graph clustering still faces two major challenges:
-  • The requirement of predefined cluster numbers $K$.
-So far, deep graph clustering cannot work without a predefined cluster number $K$, and thus one is required to correctly provide $K$ before clustering, which is often impractical in real cases.
+Challenge 1: The requirement of predefined cluster numbers $K$.
+So far, deep graph clustering typically relies on predefined cluster numbers K, and thus one is required to provide an accurate K prior to clustering, which is often impractical in real cases.
 On the one hand, it is challenging to estimate the number of node clusters.
-The methods, e.g., Bayesian information criterion [7] and recent RGC [8], repeatedly train and test the deep model to search the optimal $K$, and the searching process typically results in high computation overhead.
+The methods, e.g., Bayesian information criterion [49] and recent RGC [34], repeatedly train and test the deep model to search for the optimal $K$, and the searching process typically results in high computation overhead.
 On the other hand, existing solutions without graph structures (e.g., Bayesian non-parametric methods [18], DBSCAN [16], and hierarchical clustering [10]) cannot be directly applied owing to the inter-correlation among nodes.
-That is, deep graph clustering without $K$ is largely underexplored.
-  • Clustering the imbalanced graphs.
+Thus, deep graph clustering without $K$ remains largely underexplored.
+Challenge 2: Clustering the imbalanced graphs.
 The imbalance is prevalent in real-world graphs, where a large proportion of vertices belong to the majority clusters, while minority clusters contain significantly fewer vertices [23, 39].
@@
 Also, the challenge becomes even greater when the number of clusters is unknown.
-Figure 1: In IsoSEL, we learn the partitioning tree of structural information with the LSEnet in hyperbolic space, and then propose Lorentz tree contrastive learning to refine the partitioning tree for graph clustering without $K$.
-In this paper, we revisit deep graph clustering with an unknown $K$, considering the imbalance in real-world graphs, and approach this problem from a fresh perspective based on information theory.
-Stemming from Shannon entropy, structural information [16] has been proposed to measure the uncertainty in graphs.
-According to structural information minimization, the optimal partitioning tree can be constructed to describe a graph’s self-organization without requiring $K$.
-While this sheds light on the targeted problem, it presents several significant challenges to deep clustering: (1) Connection to Clustering.
-Structural information has recently garnered increasing research attention [17, 18, 19], but its theoretical connection to clustering remains unclear.
-As a result, structural information has not yet been introduced to graph clustering.
-(2) Discrete Formulation.
-The discrete nature of structural information prevents gradient backpropagation, posing a major challenge in training deep graph models.
-(3) Attribute Ignorance.
-Structural information operates solely on the graph structure overlooking node attributes, which are often equally important to graph clustering.
-(4) Computational Complexity.
-Constructing a partitioning tree of height $3$ yields the complexity of $O(\lvert N\rvert^{3}\log^{2}\lvert N\rvert)$, where $N$ is the number of nodes.
-Note that, the complexity scales exponentially with tree height [16], making it infeasible for real-world graphs.
+Figure 1: Illustration of ASIL.
+With a novel objective of augmented structural entropy, ASIL seamlessly integrates partitioning tree construction and contrastive learning in the Lorentz model of hyperbolic space.
+It achieves debiased node clustering in linear complexity without requiring the number of clusters.
+[…]
 To tackle these challenges, we establish a new differentiable structural information (DSI), generalizing the discrete formalism to the continuous realm.
@@
 Theoretically, we show that the structural entropy of the optimal $\mathcal{T}_{\operatorname{net}}$ well approximates that of [27] under slight constraint.
-Next, we reduce the time complexity to the linear $\mathcal{O}(N)$ with respect to the number of nodes by exploring the equivalence in level-wise assignments.
-Furthermore, we discover another advantage of the deep partitioning tree that it is able to identify the minority clusters from the majority ones, which is suitable for clustering real graphs that often exhibit imbalanced cluster distributions.
-Grounded on DSI theory, we present a novel Isometric Structural Entropy Learning (referred to as IsoSEL) to learn and refine the deep partitioning tree in the hyperbolic space for graph clustering without $K$.
-The overall architecture is illustrated in Fig.
-1, consisting of the Lorentz Structural Entropy net (LSEnet) from the conference version [^1] The preliminary results have been published in the proceedings of the 41st International Conference on Machine Learning (ICML) as an Oral Paper [20]., and Lorentz tree contrastive learning proposed in this full version.
-As in the part B of Fig.
-1, LSEnet is designed to parameterize the deep partitioning tree in the Lorentz model of hyperbolic space.
-The novelty lies in exploring the alignment between hyperbolic space and hierarchical/tree-like structures.
-Specifically, we first embed the leaf nodes of the partitioning tree using Lorentz convolution over the graph.
-Then, we recursively learn the parent nodes from the bottom (leaf nodes) to the top (root node) using the proposed Lorentz assigner.
-In IsoSEL, another key innovation is the proposed Lorentz Tree Contrastive Learning (as shown in part C of Fig.
-1), which refines the deep partitioning tree in the hyperbolic space.
-To achieve this, we first formulate the isometric augmentation to generate an augmented graph, integrating attribute similarity into structural information from the representation perspective using the notion of Gauss curvature.
-Second, we design a projection head for tree contrastive learning.
-Unlike graph contrastive learning, the main challenge in contrasting trees lies in ensuring consistency between leaf and parent nodes in hyperbolic space, given the unified nature of the partitioning tree.
-To address this, we instantiate a neural Lorentz group in hyperbolic space.
-Specifically, we parameterize the transformation in the Lorentz group using a closed form matrix-vector multiplication, which supports gradient descent optimization.
-This parameterization ensures that the projection head preserves the hyperbolic manifold and maintains consistency between leaf and parent nodes, leveraging the linearity of the Lorentz group.
-Finally, we utilize the Minkowski inner product to contrast the deep partitioning trees, refining the hyperbolic representations and enhancing cluster assignment.
-As a result, IsoSEL seamlessly integrates the strengths of both structural entropy and hyperbolic space for graph clustering.
+it is able to identify the minority clusters from the majority ones, which is suitable for clustering real graphs that often exhibit imbalanced cluster distributions.
+Furthermore, we uncover another advantage of the deep partitioning tree—its capability to distinguish minority clusters from majority ones, making it suitable for clustering real-world graphs that often exhibit imbalanced distributions.
+[…]
+Going beyond, in this full version, we aim to enhance both effectiveness and efficiency of structural entropy clustering.
+For effectiveness, we refine the partitioning tree representation in the Lorentz model of hyperbolic space via contrastive learning.
+Different from graph contrastive learning, we derive a neural Lorentz boost (Theorem VI.1, VI.2) as a special projection head of consistency insurance—pushing leaf nodes together with the parent hierarchies as a unity—and formulate a tree contrastive loss considering partitioning tree branches for clustering assignment.
+However, its quadratic complexity prevents the clustering model scaling up.
+For efficiency, we further advance our established theory by an interesting new finding (Theorem VII.1) that structural entropy indeed bounds the contrastive loss.
+Accordingly, we connect the entropy (of structural perspective) and contrastive learning (of representational perspective) through a novel virtual graph constructed from the representations of the proposed Lorentz boost, maintaining consistency insurance, so that optimizing the virtual graph implicitly performs the tree contrastive learning, as shown in Fig.
+1B.
+Meanwhile, we equivalently reformulate structural entropy by exploring level-wise assignments, achieving linear time complexity.
+Finally, we propose an expressive objective of augmented structural entropy, fusing the virtual graph described above, which seamlessly integrates partitioning tree construction and contrastive learning.
+With a provable improvement in graph conductance (Theorem VII.2), ASIL achieves effective debiased graph clustering in linear complexity with respect to the graph size.
 Contribution Highlights. The key contributions are five-fold:
@@
   • Theory.
-We establish the differentiable structural information (DSI) in continuous realm, so that a deep partitioning tree with linear computational complexity is created by gradient backpropagation for graph clustering without $K$.
+We establish the differentiable structural information in continuous realm, so that a deep partitioning tree with linear computational complexity is created by gradient backpropagation for graph clustering without $K$.
+We further connect the theory of structural information to constrastive learning by proving a new bound.
   • Neural Architecture.
-We present a novel IsoSEL approach where the hyperbolic LSEnet is designed to learn the deep partitioning tree with Lorentz convolution and assigner, connecting DSI to hyperbolic geometry.
+We present a hyperbolic clustering architecture that first fuses a parameterized virtual graph and then leverages LSEnet to learn the partitioning tree in the Lorentz model of hyperbolic space.
   • Learning Paradigm.
-We propose a Lorentz tree contrastive learning with the isometric augmentation to refine the deep partitioning tree in the hyperbolic space, while incorporating the attribute information.
+We propose a novel Augmented Structural Information Learning that integrates partitioning tree construction and contrastive learning for more effectively and efficiently debiased graph clustering without requiring $K$.
   • Experiment.
-Extensive real world results show the superiority of IsoSEL for graph clustering, and we examine that biased boundary among imbalanced clusters is alleviated with case study and visualization.
+Extensive real world results show the superiority of ASIL for graph clustering, and we examine that biased boundary among imbalanced clusters is alleviated with case study and visualization.
 II Preliminaries and Notations
@@
 The one-dimensional structural entropy is defined as follows,
-$\mathcal{H}^{1}(G)=-\sum_{v\in\mathcal{V}}\frac{d_{v}}{\operatorname{Vol}(G)}\log_{2}\frac{d_{v}}{\operatorname{Vol}(G)},\vspace{-0.03in}$ (1)
+$\mathcal{H}^{1}(G)=-\sum_{v\in\mathcal{V}}\frac{d_{v}}{\operatorname{Vol}(G)}\log_{2}\frac{d_{v}}{\operatorname{Vol}(G)},\vskip-2.168pt$ (1)
 yielding as a constant for certain given graph.
 The $H$-dimensional structural information is accompanied with a partitioning tree $\mathcal{T}$ of height $H$.
-Concretely, in $\mathcal{T}$ with root node $\lambda$, each tree node $\alpha$ is associated with a subset of $\mathcal{V}$, referred to as module $T_{\alpha}\subset\mathcal{V}$ [^2] The vertex of partitioning tree is termed as “tree node” to distinguish from the node in graphs.
+Concretely, in $\mathcal{T}$ with root node $\lambda$, each tree node $\alpha$ is associated with a subset of $\mathcal{V}$, referred to as module $T_{\alpha}\subset\mathcal{V}$.
+[^3] The vertex of partitioning tree is termed as “tree node” to distinguish from the node in graphs.
 and its immediate predecessor is written as $\alpha^{-}$.
@@
 The structural information assigned to each non-root node $\alpha$ is defined as
-$\mathcal{H}^{\mathcal{T}}(G;\alpha)=-\frac{g_{\alpha}}{\operatorname{Vol}(G)}\log_{2}\frac{V_{\alpha}}{V_{\alpha^{-}}},\vspace{-0.03in}$ (2)
+$\mathcal{H}^{\mathcal{T}}(G;\alpha)=-\frac{g_{\alpha}}{\operatorname{Vol}(G)}\log_{2}\frac{V_{\alpha}}{V_{\alpha^{-}}},\vskip-2.168pt$ (2)
 where the scalar $g_{\alpha}$ is the total weights of graph edges with exactly one endpoint in the $T_{\alpha}$, and $V_{\alpha}$ is the volume of the module $T_{\alpha}$.
 The $H$-dimensional structural information of $G$ with respect to the partitioning tree $\mathcal{T}$ is written as follows,
-$\mathcal{H}^{\mathcal{T}}(G)=\sum\nolimits_{\alpha\in\mathcal{T},\alpha\neq\lambda}\mathcal{H}^{\mathcal{T}}(G;\alpha).\vspace{-0.03in}$ (3)
+$\mathcal{H}^{\mathcal{T}}(G)=\sum\nolimits_{\alpha\in\mathcal{T},\alpha\neq\lambda}\mathcal{H}^{\mathcal{T}}(G;\alpha).\vskip-2.168pt$ (3)
 Traversing all possible partitioning trees of $G$ with height $H$, the $H$-dimensional structural entropy of $G$ is defined as
-$\mathcal{H}^{H}(G)=\min_{\mathcal{T}}\mathcal{H}^{\mathcal{T}}(G),\quad\mathcal{T}^{*}=\arg_{\mathcal{T}}\min\mathcal{H}^{\mathcal{T}}(G),\vspace{-0.03in}$ (4)
+$\mathcal{H}^{H}(G)=\min_{\mathcal{T}}\mathcal{H}^{\mathcal{T}}(G),\quad\mathcal{T}^{*}=\arg_{\mathcal{T}}\min\mathcal{H}^{\mathcal{T}}(G),\vskip-2.168pt$ (4)
 where $\mathcal{T}^{*}$ is the optimal partitioning tree of $G$ that minimizes the uncertainty of graph, and encodes the self-organization of node clusters in the tree structure.
@@
 Without loss of generality, a $d$-dimensional Lorentz model $\mathbb{L}^{\kappa,d}$ with curvature $\kappa$ is defined on the smooth manifold embedded in $\mathbb{R}^{d+1}$ space,
-$\mathbb{L}^{\kappa,d}=\{\boldsymbol{x}\in\mathbb{R}^{d+1}:\langle\boldsymbol{x},\boldsymbol{x}\rangle_{\mathbb{L}}=\frac{1}{\kappa},x_{0}>0\},\vspace{-0.03in}$ (5)
+$\mathbb{L}^{\kappa,d}=\{\boldsymbol{x}\in\mathbb{R}^{d+1}:\langle\boldsymbol{x},\boldsymbol{x}\rangle_{\mathbb{L}}=\frac{1}{\kappa},x_{0}>0\},\vskip-5.05942pt$ (5)
 where the Minkowski inner product is given as follows,
-$\langle\boldsymbol{x},\boldsymbol{y}\rangle_{\mathbb{L}}=\boldsymbol{x}\mathbf{R}\boldsymbol{y},\quad\mathbf{R}=\operatorname{diag}(-1,1,\cdots,1),\vspace{-0.03in}$ (6)
+$\langle\boldsymbol{x},\boldsymbol{y}\rangle_{\mathbb{L}}=\boldsymbol{x}\mathbf{R}\boldsymbol{y},\quad\mathbf{R}=\operatorname{diag}(-1,1,\cdots,1),\vskip-5.05942pt$ (6)
 $\mathbf{R}\in\mathbb{R}^{(d+1)\times(d+1)}$ describes the Minkowski metric in $\mathbb{R}^{d+1}$ space.
@@
 In group theory, the Lorentz group studies the Lorentz model of hyperbolic space.
-To be specific, Lorentz group is a (generalized) orthogonal transformation group that collects the Lorentz transformation, preserving the isometry regarding the Minkowski metric.
-With polar decomposition, any Lorentz transformation can be decomposed into a combination of Lorentz boost and Lorentz rotation [21, 22].
+To be specific, Lorentz group is a (generalized) orthogonal transformation group that collects the Lorentz transformation, consisting of Lorentz boosts and Lorentz rotations, which preserves the isometry regarding the Minkowski metric [9, 47].
 Both of them are given as special orthogonal matrices, posing a significant challenge to parameterization.
-Please refer to Appendix B for further facts.
+Further facts on hyperbolic space and Lorentz group are provided in Appendix V.
 II-4 Notations
@@
 The notation table is provided in Appendix I.
-In this paper, we introduce the novel Differentiable Structural Information (DSI) and present Isometric Structural Entropy Learning (IsoSEL) for deep graph clustering.
-The key innovation of DSI lies in constructing a deep partitioning tree in hyperbolic space through gradient backpropagation.
-Additionally, we propose Lorentz tree contrastive learning with theoretical guarantees to further refine the deep partitioning tree, enhancing cluster assignment.
+In this paper, we establish the theory of differential structural information, demonstrating its capacity in graph clustering and connecting it to contrastive learning.
+Thus, we approach graph clustering through Augmented Structural Entropy Learning (ASIL), whose key innovation lies in the integration of partitioning tree construction and contrastive learning.
+Without requiring the number of clusters, ASIL effectively achieves debiased node clustering in linear complexity with respect to the graph size.
 III Differentiable Structural Information
@@
 The main theoretical results are dual: the proposed formulation emerges as a new objective for deep graph clustering without $K$; the corresponding partitioning tree can be created via gradient-based deep learning, different from the previous heuristic algorithm [27].
-III-A New Formulation
-The structural information in Eq.
-(1) cannot be optimized via gradient-based methods.
-To bridge this gap, we present a new formulation of structural information with learnable level-wise assignment, which is shown to be equivalent to the classic formulation in Eq.
-(3).
+III-A New Formulation: A Differentiable Equivalence
+We introduce a differentiable version of structural entropy where the hard indicator for calculating the volume is relaxed as soft level-wise assignments.
+It is shown to be equivalent to the classic formulation, while enabling the optimization via gradient backpropagation.
 Definition III.1 (Level-wise Assignment).
-[…]
+For a partitioning tree $\mathcal{T}$ with height $H$, assuming that the number of tree nodes at the $h$-th level is $N_{h}$, we define a level-wise parent assignment matrix $\mathbf{C}^{h}\in\{0,1\}^{N_{h}\times N_{h-1}}$ from $h$-th to (h1)(h-1)-th level, where $\mathbf{C}^{h}_{ij}=1$ means the $i$-th node of $\mathcal{T}$ at $h$-th level is the child node of $j$-th node at (h1)(h-1)-th level.
 Definition III.2 (H-dimensional Structural Information).
 For a graph $G$ and its partitioning tree $\mathcal{T}$ in Definition III.1, we rewrite the formula of $H$-dimensional structural information of $G$ with respect to $\mathcal{T}$ at height $h$ as
-$\mathcal{H}^{\mathcal{T}}(G;h)=-\frac{1}{V}\sum\limits_{k=1}^{N_{h}}(V^{h}_{k}-\sum\limits_{(i,j)\in\mathcal{E}}S^{h}_{ik}S^{h}_{jk}w_{ij})\log_{2}\frac{V^{h}_{k}}{V^{h-1}_{k^{-}}}\vspace{-0.03in}$ (7)
+$\mathcal{H}^{\mathcal{T}}(G;h)=-\frac{1}{V}\sum\limits_{k=1}^{N_{h}}(V^{h}_{k}-\sum\limits_{(i,j)\in\mathcal{E}}S^{h}_{ik}S^{h}_{jk}w_{ij})\log_{2}\frac{V^{h}_{k}}{V^{h-1}_{k^{-}}}\vskip-2.168pt$ (7)
 where $V=\operatorname{Vol}(G)$ is the volume of $G$.
@@
 $\mathbf{S}^{h}$ $=\prod\nolimits_{k=H+1}^{h+1}\mathbf{C}^{k},\quad\mathbf{C}^{H+1}=\mathbf{I}_{N},$ (8)
-$V^{h}_{k}$ $=\sum\nolimits_{i=1}^{N}S^{h}_{ik}d_{i},\quad V^{h-1}_{k^{-}}=\sum\nolimits_{k^{\prime}=1}^{N_{h-1}}\mathbf{C}_{kk^{\prime}}^{h}V^{h-1}_{k^{\prime}}.\vspace{-0.03in}$ (9)
+$V^{h}_{k}$ $=\sum\nolimits_{i=1}^{N}S^{h}_{ik}d_{i},\quad V^{h-1}_{k^{-}}=\sum\nolimits_{k^{\prime}=1}^{N_{h-1}}\mathbf{C}_{kk^{\prime}}^{h}V^{h-1}_{k^{\prime}}.\vskip-2.168pt$ (9)
 Then, the $H$-dimensional structural information of $G$ is $\mathcal{H}^{\mathcal{T}}(G)=\sum\nolimits_{h=1}^{H}H^{\mathcal{T}}(G;h)$.
@@
 The $1$-dimensional structural entropy of $G$ can be decomposed as follows
-$\mathcal{H}^{1}(G)=\sum_{h=1}^{H}\sum_{j=1}^{N_{h-1}}\frac{V^{h-1}_{j}}{V}E([\frac{C^{h}_{kj}V^{h}_{k}}{V^{h-1}_{j}}]_{k=1,...,N_{h}}),\vspace{-0.03in}$ (10)
+$\mathcal{H}^{1}(G)=\sum_{h=1}^{H}\sum_{j=1}^{N_{h-1}}\frac{V^{h-1}_{j}}{V}E([\frac{C^{h}_{kj}V^{h}_{k}}{V^{h-1}_{j}}]_{k=1,...,N_{h}}),\vskip-2.168pt$ (10)
 where $E(p_{1},...,p_{n})=-\sum_{i=1}^{n}p_{i}\log_{2}p_{i}$ is the entropy.
 {mymath}
-Theorem III.5 (Connection Graph Clustering).
+Theorem III.5 (Connection to Graph Clustering).
 Given a graph $G=(\mathcal{V},\mathcal{E})$ with $w$, the normalized $H$-structural entropy of graph $G$ is defined as $\tau(G;H)={\mathcal{H}^{H}(G)}/{\mathcal{H}^{1}(G)}$, and $\Phi(G)$ is the graph structural conductance.
 With the additivity (Lemma III.4), the following inequality holds,
-$\tau(G;H)\geq\Phi(G).\vspace{-0.03in}$ (11)
+$\tau(G;H)\geq\Phi(G).\vskip-2.168pt$ (11)
 Proof.
@@
 $=\Phi(G)\sum_{h=1}^{H}\sum_{j=1}^{N_{h-1}}\frac{V^{h-1}_{j}}{V}E([\frac{C^{h}_{kj}V^{h}_{k}}{V^{h-1}_{j}}]_{k=1,...,N_{h}})$
-$=\Phi(G)\mathcal{H}^{1}(G).\vspace{-0.03in}$ (14)
+$=\Phi(G)\mathcal{H}^{1}(G).\vskip-2.168pt$ (14)
 Since $\tau(G;\mathcal{T})=\frac{\mathcal{H}^{\mathcal{T}}(G)}{\mathcal{H}^{1}(G)}\geq\Phi(G)$ holds for every $H$-height partitioning tree $\mathcal{T}$ of $G$, $\tau(G)\geq\Phi(G)$ holds. ∎
@@
 In short, our formulation is a differentiable objective function for graph clustering without predefined cluster numbers.
+Remark.
+Different from classical graph entropies measuring the uncertainty on graph, structural entropy also models the hierarchical self-organization of graphs through a partitioning tree.
+This enables it to serve as a clustering objective without requiring the number of clusters, a property absent in traditional graph entropy measures.
 III-C Deep Partitioning Tree
@@
 IV Theory: Structural Information & Graph Clustering
-In this section, we derive an equivalent form of differential structural information to reduce its complexity, and provide the new theoretical results on the advantages of structural information for graph clustering.
+In this section, we derive a more efficient formulation of differential structural information, achieving linear complexity, and provide the new theoretical results on the advantages of structural information for graph clustering.
 IV-A Reducing Complexity
@@
 $\mathcal{H}^{\mathcal{T}}(G^{h};G^{h-1},h)=-\frac{1}{V}\sum_{k=1}^{N_{h}}(d^{h}_{k}-w^{h}_{kk})\log\frac{d^{h}_{k}}{\sum_{l=1}^{N_{h-1}}C_{kl}^{h-1}d_{l}^{h-1}},$
-[…]
+where $G^{h-1}$ is induced by the assignment matrix $\mathbf{S}$ in Eq.
+(8) at the (h1)(h-1) level, and $d_{i}^{h}$ is the degree of $i$-th node in graph $G^{h}$.
+The equivalence and variables in Eq.
+(17) are stated in the following theorem.
+{mymath}
 Theorem IV.1 (Node-wise Formulation).
@@
 Refer to Appendix II, Theorem IV.3. ∎
-Nodes within the cluster are typically densely connected while there are fewer links connecting the nodes of different clusters. The condition is easy to be satisfied in real graphs.
+Nodes within the cluster are typically densely connected while there are fewer links connecting the nodes of different clusters. The condition is easy to satisfy in real graphs.
 That is, we obtain the graph clustering objective with linear time complexity, the differential structural information in Eq.
 (17), which no longer requires the knowledge of the number of clusters and is suitable to identify minority clusters in real-world graphs often exhibiting skewed cluster distributions.
-V Isometric Graph Augmentation
-The structural information solely handles the graph structure, while node attributes are often equally important to graph clustering.
-To fill this gap, we propose isometric augmentation which integrates attribute information into graph structure, generalizing existing methods from representation perspective.
-V-A Isometry between Manifolds
-We demonstrate that arbitrary embedding induces a Riemannian manifold of hypersurface, and the existence of this manifold is claimed in the following lemma.
-Lemma V.1 (Existence).
-Given a set of points $\{\boldsymbol{p}_{i}\}_{i=1,..,N}$ on $\mathbb{R}^{n+1}$ derived from certain embedding, there exists a hypersurface $\mathcal{M}$ such that $\boldsymbol{p}_{i}\in\mathcal{M}$ holds for all $i$.
-Proof.
-Refer to Appendix II, Lemma V.1. ∎
-Now, we elaborate on the isometry between manifolds.
-Specifically, we leverage the notion of Gaussian curvature, which depicts the local geometry of the manifold.
-Considering a point $\boldsymbol{p}_{i}\in\mathcal{M}$ on the manifold, the Gaussian curvature at $\boldsymbol{p}$ is intrinsically defined by the rate at which the geodesics on the manifold of hypersurface, emanating from $\boldsymbol{p}$, spread apart and this curvature is invariant among coordinate systems.
-Two hypersurfaces are said to be isometric if their Gaussian curvature are preserved relative to each other according to Gauss’s Theorema Egregium [24].
-We develop the geometric intuition: The isometry means that two hypersurfaces exhibit the same pattern on how it curves at each point.
-To find an isometry between hypersurfaces, we present our theoretical result on Givens rotation matrix $\mathbf{G}_{\theta}$.
-To be specific, the element of $\mathbf{G}_{\theta}$ is defined as follows,
-$\mathbf{G}_{\theta}[i,j]=\mathbf{I}[i,j]+\beta_{ii}(\cos\theta-1)+\beta_{ij}\sin\theta,\vspace{-0.03in}$ (22)
-where $\beta_{ii}=e_{i}e_{i}^{\top}+e_{j}e_{j}^{\top}$, $\beta_{ij}=e_{i}e_{j}^{\top}+e_{j}e_{i}^{\top}$, $\mathbf{I}$ is the identity, and $e$’s denote the orthogonal bases of the hypersurface.
-The matrix $\mathbf{G}_{\theta}$ characterizes the rotation along the span of $e_{i}$ and $e_{j}$ with the rotation angle of $\theta$.
-We prove the following statement holds for arbitrary hypersurfaces.
-{mymath}
-Theorem V.2 (Rotational Isometry).
-Given two hypersurfaces $\mathcal{S}_{1},\mathcal{S}_{2}$ in $\mathbb{R}^{n+1}$, equipped with the usual inner product, the Givens rotation $\mathbf{G}_{\theta}$ can be an isometry, such that $\mathbf{G}_{\theta}:\mathcal{S}_{1}\rightarrow\mathcal{S}_{2}$, and vice versa.
-Proof.
-Refer to Appendix II, Theorem V.2. ∎
-In other words, the Givens rotation instantiate an isometry between hypersurfaces in terms of Gauss’s Theorema Egregium.
-V-B Augmentation with Rotational Isometry
-Typical integration methods [25] create an augmented graph $\hat{G}(\hat{\mathbf{A}},\mathbf{X})$ that aggregates the adjacency matrix and the similarity matrix of node attributes, given a graph $G(\mathbf{A},\mathbf{X})$.
-The augmented $\hat{\mathbf{A}}$ is given by the adjacency matrix $\mathbf{A}$ and the similarity $\mathbf{S}$ derived from the attributes $\mathbf{X}$, e.g., the structure of K-nearest neighbor graph based on attribute similarity.
-In IsoSEL, we propose a more generic approach (referred to as isometric augmentation) from the representation perspective.
-The key idea is to perform aggregation on the manifolds and derive the augmented structure with the notion of isometry.
-First, we induce the manifold of $G$ with message-passing embedding, taking the popular SGC [26] as an example.
-Hence two sets of embeddings are obtained from adjacency matrix and attribute similarity, formulated as $\mathbf{A}^{L}\mathbf{X}$ and $\mathbf{S}^{L}\mathbf{X}$, respectively.
-The message-passing is conducted $L$ times without loss of generality.
-Second, we select an aggregation function $f$ to construct the hypersurface whose existence is guaranteed given Lemma V.1.
-$f$ can be a weighted sum or product, and the corresponding embedding is written as $f(\mathbf{A},\mathbf{S},\mathbf{X};L)$.
-Next, we aim to derive an augmented graph whose induced hypersurface is isometric to that of $f(\mathbf{A},\mathbf{S},\mathbf{X};L)$ in terms of Gauss’s Theorema Egregium.
-However, specifying the relationship between the embedding of isometric hypersurfaces is challenging.
-Thanks to the discovery of Theorem V.2, the augmented $\hat{\mathbf{A}}$ can be formulated in the optimization problem as follows.
-{mymath}
-Theorem V.3 (Optimization of Isometric Augmentation).
-Given a graph with an adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ and a similarity matrix $\mathbf{S}\in\mathbb{R}^{N\times N}$ induced by the attributes $\mathbf{X}\in\mathbb{R}^{N\times d}$, the adjacency matrix of augmented graph $\hat{\mathbf{A}}\in\mathbb{R}^{N\times N}$ considering the isometry in the $L$ convolution is the closed-form minimizer of the following optimization
-$\min\ \|f(\mathbf{A},\mathbf{S},\mathbf{X};L)\mathbf{G}^{\top}_{\theta}-\hat{\mathbf{A}}^{L}\mathbf{X}\|^{2},\vspace{-0.05in}$ (23)
-subject to the symmetry matrix cone of $\hat{\mathbf{A}}$, where $\mathbf{G}_{\theta}\in\mathbb{R}^{d\times d}$ denotes the Givens rotation, and $f(\mathbf{A},\mathbf{S},\mathbf{X};L)\in\mathbb{R}^{N\times d}$ is the aggregator of structural and attribute representations.
-Proof.
-Refer to Appendix II, Theorem V.3. ∎
-Closed-form Expression. The construction above enjoys the existence of closed-form solution. In particular, the augmented graph is expressed as follows,
-$\hat{\mathbf{A}}=\mathbf{P}\operatorname{diag}(\lambda_{1}^{\frac{1}{L}},\cdots,\lambda_{d}^{\frac{1}{L}})\mathbf{P}^{\top},$ (24)
-where $\mathbf{P}$ and $\lambda$’s are derived as the eigenvectors and eigenvalues of the following matrix
-We have $\mathbf{B}=f(\mathbf{A},\mathbf{S},\mathbf{X};L)\mathbf{G}_{\theta}^{T}$, and $\mathbf{U}=\mathbf{X}\mathbf{X}^{T}+\omega\mathbf{I}$ with penalty coefficient $\omega$.
-Note that, the eigenvalue decomposition is well-established with efficient methods [27] while the augmentation is typically applied as a preprocessing.
-The advantage is that we improve the expressiveness of the polynomial over structural and attribute matrices by incorporating representation consideration, as shown in Ablation Study, while recovering previous methods [25] as special cases.
-VI LSEnet: Learning Partitioning Tree in Hyperbolic Space
+V LSEnet: Learning Partitioning Tree in Hyperbolic Space
 We design a novel Lorentz Structural Entropy Neural Network (LSEnet), which aims to learn the optimal partitioning tree $\mathcal{T}_{\operatorname{net}}^{*}$ in the Lorentz model of hyperbolic space.
@@
 In particular, we learn the partitioning trees from both the original and augmented structures, which serve as the different views in Lorentz tree contrastive learning (Sec.
-VII).
-VI-A Why Hyperbolic Space?
+VI).
+V-A Why Hyperbolic Space?
 First, we demonstrate the reason why we opt for the hyperbolic space, rather than the traditional Euclidean space.
@@
 $d_{G}$ and $d$ denote the distance in the graph and the space, respectively.
-Theorem VI.1 (Tree $\mathcal{T}$ and Hyperbolic Space [28]).
+Theorem V.1 (Tree $\mathcal{T}$ and Hyperbolic Space [48]).
 $\forall\mathcal{T}$, scaling all edges by a constant so that the edge length is bounded below $\nu\frac{1+\epsilon}{\epsilon}$, there exists an embedding in hyperbolic space that the distortion overall node pairs are bounded by $1+\epsilon$.
 ($\nu$ is a constant detailed in Appendix V)
-That is, hyperbolic space is well suited to embed the partitioning tree. Note that, Theorem VI.1 does not hold for Euclidean space or other Riemannian manifold.
-VI-B Embedding Leaf Nodes
+That is, hyperbolic space is well suited to embed the partitioning tree. Note that, Theorem V.1 does not hold for Euclidean space or other Riemannian manifold.
+V-B Embedding Leaf Nodes
 We design a Lorentz convolution layer to learn leaf embeddings in hyperbolic space, $\operatorname{LConv}:\boldsymbol{x}_{i}\to\boldsymbol{z}^{H}_{i},\forall v_{i}$, where $\boldsymbol{x}_{i}\in\mathbb{L}^{\kappa,d_{0}}$ and $\boldsymbol{z}^{H}_{i}\in\mathbb{L}^{\kappa,d_{\mathcal{T}}}$ are the node feature and leaf embedding, respectively.
@@
 For node $\boldsymbol{x}_{i}\in\mathbb{L}^{\kappa,d_{0}}$, the linear operator is given as
-$\operatorname{LLinear}(\boldsymbol{x})=\begin{bmatrix}\sqrt{\lVert h(\mathbf{W}\boldsymbol{x},\boldsymbol{v})\rVert^{2}-\frac{1}{\kappa}}\ \\ h(\mathbf{W}\boldsymbol{x},\boldsymbol{v})\ \end{bmatrix}\in\mathbb{L}^{\kappa,d_{\mathcal{T}}},\vspace{-0.03in}$ (26)
+$\operatorname{LLinear}(\boldsymbol{x})=\begin{bmatrix}\sqrt{\lVert h(\mathbf{W}\boldsymbol{x},\boldsymbol{v})\rVert^{2}-\frac{1}{\kappa}}\ \\ h(\mathbf{W}\boldsymbol{x},\boldsymbol{v})\ \end{bmatrix}\in\mathbb{L}^{\kappa,d_{\mathcal{T}}},\vskip-2.168pt$ (22)
 where $h$ is a neural network, and $\mathbf{W}$ and $\boldsymbol{v}$ are parameters.
@@
 $a_{ij}$ $=\operatorname{LeakyReLU}(\mathbf{W}[\boldsymbol{q}_{i}||\boldsymbol{k}_{j})])$ (23)
-$\omega_{ij}$ $=\operatorname{LAtt}(\mathbf{Q},\mathbf{K})=\frac{\exp(a_{ij})}{\sum_{(i,l)\in\mathcal{E}}\exp(a_{il})},\vspace{-0.03in}$ (28)
-where $\mathbf{W}\in\mathbb{R}^{1\times 2d}$ is learnable parameters, [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] is concatenate operation, $\boldsymbol{q}$ and $\boldsymbol{k}$ are the query and key vector collected in row vector of $\mathbf{Q}$ and $\mathbf{K}$, respectively.
+$\omega_{ij}$ $=\operatorname{LAtt}(\mathbf{Q},\mathbf{K})=\frac{\exp(a_{ij})}{\sum_{(i,l)\in\mathcal{E}}\exp(a_{il})},\vskip-2.168pt$ (24)
+where $\mathbf{W}\in\mathbb{R}^{1\times 2d}$ is learnable parameters, $\cdot$$\cdot$] is concatenate operation, $\boldsymbol{q}$ and $\boldsymbol{k}$ are the query and key vector collected in row vector of $\mathbf{Q}$ and $\mathbf{K}$, respectively.
 The queries and keys are derived from node feature $\boldsymbol{x}$ via $\operatorname{LLinear}$ with different parameters.
@@
 Given a set of points $\{\boldsymbol{x}_{i}\}_{i=1,\cdots,N}$ in the Lorentz model $\mathbb{L}^{\kappa,d_{\mathcal{T}}}$,
-$\operatorname{LAgg}(\boldsymbol{\omega},\mathbf{X})=\frac{1}{\sqrt{-\kappa}}\sum_{i=1}^{N}\frac{\omega_{i}}{\lvert\lVert\sum_{j=1}^{N}\omega_{j}\boldsymbol{x}_{j}\rVert_{\mathbb{L}}\rvert}\boldsymbol{x}_{i},\vspace{-0.03in}$ (29)
+$\operatorname{LAgg}(\boldsymbol{\omega},\mathbf{X})=\frac{1}{\sqrt{-\kappa}}\sum_{i=1}^{N}\frac{\omega_{i}}{\lvert\lVert\sum_{j=1}^{N}\omega_{j}\boldsymbol{x}_{j}\rVert_{\mathbb{L}}\rvert}\boldsymbol{x}_{i},\vskip-2.168pt$ (25)
 where $\boldsymbol{\omega}$ is the weight vector, and $\boldsymbol{x}$ are summarized in $\mathbf{X}$.
@@
 Overall, the Lorentz convolution layer is formulated as,
+$\operatorname{LConv}(\mathbf{X}|\mathbf{A})=\operatorname{LAgg}(\operatorname{LAtt}(\mathbf{Q},\mathbf{K})\odot\mathbf{A},\operatorname{LLinear}(\mathbf{X})),$
 where $\odot$ is the Hadamard product, masking the attentional weight if the corresponding edge does not exist in the graph.
-VI-C Learning Parent Nodes
+V-C Learning Parent Nodes
 A primary challenge is that, in the partitioning tree, the node number at each internal level is unknown.
@@
 Theoretically, if an internal level has insufficient nodes, the self-organization of the graph can still be described by multiple levels in the partitioning tree.
-[…]
-$\mathbf{C}^{h}=\operatorname{LAtt}(\mathbf{Q},\mathbf{K})\sigma(\operatorname{MLP}(\mathbf{Z}^{h})),\vspace{-0.03in}$ (31)
+Without loss of generality, we consider the assignment between $h$-th and (h1)(h-1)-th levels given node embeddings $\boldsymbol{z}^{h}\in\mathbb{L}^{\kappa,d_{\mathcal{T}}}$ at $h$-th level.
+Recalling Definition III.1, the $i$-th row of assignment $\mathbf{C}^{h}\in\mathbb{R}^{N_{h}\times N_{h-1}}$ describes the belonging of $i$-th node at $h$-th to the parent nodes at (h1)(h-1)-th.
+We design a Lorentz assigner following the intuition that neighborhood nodes in the graph tend to have similar assignments.
+Concretely, we leverage Multilayer Perceptron $\operatorname{MLP}$ to learn the assignment from embeddings, and meanwhile the similarity is parameterized by $\operatorname{LAtt}$ defined in Eq.
+24.
+Thus, the Lorentz assigner is formulated as follows,
+$\mathbf{C}^{h}=\operatorname{LAtt}(\mathbf{Q},\mathbf{K})\sigma(\operatorname{MLP}(\mathbf{Z}^{h})),\vskip-2.168pt$ (27)
 where $\mathbf{A}^{h}$ is the graph structure at $h$-th level of the tree. $\sigma$ denotes the row-wise $Softmax$ function.
-[…]
-$\boldsymbol{z}^{h-1}_{j}=\arg_{\boldsymbol{z}^{h-1}_{j}}\min\sum\nolimits_{i=1}^{N}c_{ij}d^{2}_{\mathbb{L}}(\boldsymbol{z}^{h-1}_{j},\boldsymbol{z}^{h}_{i}).\vspace{-0.03in}$ (32)
+The remaining task is to infer the node embeddings $\boldsymbol{z}^{h-1}\in\mathbb{L}^{\kappa,d_{\mathcal{T}}}$ at (h1)(h-1)-th level.
+As reported in [7], a parent node is located at the point that has the shortest path to all the child nodes at the $h$-th level, correspondingly, the parent node is the Fréchet mean of the child nodes in the manifold.
+The challenge here is that Fréchet mean regarding the canonical distance has no closed-form solution [40].
+Alternatively, we consider the geometric centroid for the squared Lorentz distance, where the weights are given by the soft assignment $\mathbf{C}^{h}$,
+$\boldsymbol{z}^{h-1}_{j}=\arg_{\boldsymbol{z}^{h-1}_{j}}\min\sum\nolimits_{i=1}^{N}c_{ij}d^{2}_{\mathbb{L}}(\boldsymbol{z}^{h-1}_{j},\boldsymbol{z}^{h}_{i}).\vskip-2.168pt$ (28)
 Solving the optimization constrained in the manifold $\mathbb{L}^{\kappa,d_{\mathcal{T}}}$, we derive the closed-form solution of parent node embeddings,
@@
 which is the augmented form of Eq. (25). {mymath}
-Theorem VI.2 (Geometric Centroid).
+Theorem V.2 (Geometric Centroid).
 In Lorentz model $\mathbb{L}^{\kappa,d_{\mathcal{T}}}$ of hyperbolic space, for any set of points $\{\boldsymbol{z}^{h}_{i}\}$, the arithmetic mean is given as follows,
-$\boldsymbol{z}^{h-1}_{j}=\frac{1}{\sqrt{-\kappa}}\sum_{i=1}^{N}\frac{c_{ji}}{\lvert\lVert\sum_{l=1}^{n}c_{jl}\boldsymbol{z}^{h}_{l}\rVert_{\mathbb{L}}\rvert}\boldsymbol{z}^{h}_{i}\in\mathbb{L}^{\kappa,d_{\mathcal{T}}},\vspace{-0.03in}$ (34)
+$\boldsymbol{z}^{h-1}_{j}=\frac{1}{\sqrt{-\kappa}}\sum_{i=1}^{N}\frac{c_{ji}}{\lvert\lVert\sum_{l=1}^{n}c_{jl}\boldsymbol{z}^{h}_{l}\rVert_{\mathbb{L}}\rvert}\boldsymbol{z}^{h}_{i}\in\mathbb{L}^{\kappa,d_{\mathcal{T}}},\vskip-2.168pt$ (30)
 and is the closed-form solution of the geometric centroid defined in the minimization of Eq. 28.
-Remark. In fact, the geometric centroid in Theorem VI.2 is also equivalent to the gyro-midpoint in Poincaré ball model of hyperbolic space, detailed in Appendix II, Theorem VI.2.
-[…]
-$\mathbf{A}^{h-1}$ $=(\mathbf{C}^{h})^{\top}\mathbf{A}^{h}\mathbf{C}^{h}.\vspace{-0.03in}$ (35)
+Remark.
+In fact, the geometric centroid in Theorem V.2 is also equivalent to the gyro-midpoint in Poincaré ball model of hyperbolic space, detailed in Appendix II, Theorem VI.2.
+Note that attentive aggregation for parent node embedding is uni-directional, different from that in typical graph neural nets.
+According to Eq.
+(27), it requires the graph structure $\mathbf{A}^{h-1}$ among the parent nodes at (h1)(h-1)-th level to derive the assignment.
+We give the adjacency matrix of the (h1)(h-1)-th level graph as follows,
+$\mathbf{A}^{h-1}$ $=(\mathbf{C}^{h})^{\top}\mathbf{A}^{h}\mathbf{C}^{h}.\vskip-2.168pt$ (31)
 We recursively utilize Eqs. (27), (29) and (31) to build the partitioning tree from bottom to top in hyperbolic space.
-VI-D Hyperbolic Partitioning Tree
+V-D Hyperbolic Partitioning Tree
 In the principle of structural information minimization, the optimal partitioning tree is constructed to describe the self-organization of the graph [27, 63].
@@
 DSI is applied on all the level-wise assignment $\mathbf{C}$’s, which are parameterized by neural networks in Sec.
-VI-B and VI-C.
+V-B and V-C.
 We suggest to place the tree root at the origin of $\mathbb{L}^{\kappa,d_{\mathcal{T}}}$, so that the learnt tree enjoys the symmetry of Lorentz model.
 Consequently, the hyperbolic partitioning tree describes the graph’s self-organization and clustering structure in light of Theorem III.5.
-VII Graph Clustering with
-Lorentz Tree Contrastive Learning
-In this section, we introduce the Lorentz Tree Contrastive Learning designed to contrast trees in the hyperbolic space, achieving two key objectives.
-First, it incorporates attribute information in a contrastive manner to overcome the limitations of structural entropy.
-Second, it refines the hyperbolic tree representations, thereby enhancing clustering assignment.
-The main components of Lorentz Tree Contrastive Learning are i) Lorentz projection head, ii) hyperbolic reweighed sampling and iii) the Lorentz contrastive loss.
-VII-A Neural Transformation in Lorentz Group
-The projection head for contrasting partitioning trees differs significantly from typical graph contrastive learning.
-The key distinctions are as follows.
-First, existing methods primarily explore similarity in Euclidean space, whereas the partitioning tree is aligned with hyperbolic space, presenting a manifold-preserving challenge.
-In other words, the projection head is required to connect two hyperbolic spaces.
-Second, in graph contrastive learning, the projection head operates on graph nodes (leafs) in graph contrastive learning, while in the tree contrastive learning, it manipulates tree nodes across all levels simultaneously.
-Notably, maintaining consistency between successor and ancestor nodes is crucial owing to the unity of the tree, as supported by empirical evidence in Sec.
-VIII-E.
-However, ensuring consistency in hyperbolic space remains a significant challenge.
-Before elaborating on the proposed projection head, we revisit the notion of Lorentz group $G^{L}=\{L_{i}\}$.
-To be specific, Lorentz group is a set of Lorentz transformations $L$, and $L_{i}\circ L_{j}\in G^{L}$ is closed on the group for any $L_{i},L_{j}\in G^{L}$.
-For any $\boldsymbol{x}\in\mathbb{L}^{\kappa,d_{1}}$, the equality of $L(\boldsymbol{x})=\mathbf{L}\boldsymbol{x}\in\mathbb{L}^{\kappa,d_{2}}$ holds for the block diagonal $\mathbf{L}=\left[\begin{array}[]{cc}w&\boldsymbol{v}^{\top}\\ \boldsymbol{v}&\mathbf{W}\end{array}\right],$ provided that $\mathbf{W}$ characterizes a Lorentz boost or Lorentz rotation.
-The advantage of Lorentz transformation lies in its clean matrix-vector multiplication form.
-However, the condition on $\mathbf{W}$ makes the optimization problematic; for instance, Lorentz rotation is restricted to the orthogonal cone of $\mathbf{W}$.
-To bridge this gap, we introduce a novel Lorentz-type group consisting of learnable transformations that (i) connect different hyperbolic spaces (manifold-preserving) and (ii) ensure consistency for tree contrastive learning.
-The transformation is written as $L(\boldsymbol{x})=\mathbf{L}\boldsymbol{x}\in\mathbb{L}^{\kappa,d_{2}}$ for arbitrary $\boldsymbol{x}\in\mathbb{L}^{\kappa,d_{1}}$, where $d_{1}$ and $d_{2}$ are not necessarily the same.
-$\mathbf{L}$ takes the form of $\mathbf{L}:=\begin{bmatrix}\mathbf{v}^{T}\\ \mathbf{W}\end{bmatrix}$ satisfying the equation of
-$\mathbf{W}^{T}\mathbf{W}=\mathbf{v}\mathbf{v}^{T}+diag(-1,1,...,1)\vspace{-0.05in}$ (36)
-with $\mathbf{W}\in\mathbb{R}^{d_{2}\times d_{1}}$, and $d_{2}>d_{1}$ in the projection head. Next, we show the theoretical results on our construction. {mymath}
-Theorem VII.1 (Manifold-preserving).
-Given the construction of $\mathbf{L}$ in Eq. (36) and scalars $a,b$, the following properties holds for any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{L}^{d_{1}}$.
+VI Lorentz Tree Contrastive Learning
+The partitioning tree above resides in the hyperbolic space, and thus we propose to refine its hyperbolic representation through Lorentz tree contrastive learning, enhancing the graph semantics from a representational perspective.
+In this contrastive learning framework, we develop a novel neural Lorentz boost for consistency insurance and a tree contrastive loss considering the tree branches in light of clustering assignment.
+VI-A Neural Lorentz Boost
+Different from graph contrastive learning, the projection head in our case further requires the consistency insurance—maintaining the parent hierarchies of the partitioning tree as a unity.
+That is, the level-wise child-parent assignment cannot be disrupted by this projection head.
+To achieve this, we leverage the Lorentz group $G^{L}=\{L_{i}\}$, whose elements are isometries over Lorentz model.
+For any $\boldsymbol{x}\in\mathbb{L}^{\kappa,d}$, a Lorentz transformation can be written as follows,
+[…]
+where $\mathbf{W}\in\mathbb{R}^{d\times d}$ is a matrix, $w$ is a scaling scalar, and $\boldsymbol{v}\in\mathbb{R}^{d}$ is a boost vector.
+Note that Lorentz rotations act tangentially to the radial direction.
+They rotate the partitioning tree and do not affect representation refinement due to rotational symmetry.
+Thus, we consider the Lorentz boosts, which serve as a linear transformation to operate on the partitioning tree in the Lorentz model of hyperbolic space.
+We derive a closed-form Lorentz boost parameterized by $\boldsymbol{\beta}\in\mathbb{R}^{d}$, maintaining the consistency insurance in partitioning tree contrasting (as shown in Theorems VI.1 and VI.2).
+{mymath}
+Theorem VI.1 (Closed-from Lorentz Boost).
+Given the construction in Eq. (32) and a parameter vector $\boldsymbol{\beta}\in\mathbb{R}^{d}$, Lorentz boosts have a closed form with components given as
+$w$ $=\frac{1}{\sqrt{1-\|\boldsymbol{\beta}\|^{2}_{\mathbb{L}}}}$ (33)
+$\boldsymbol{v}$ $=-w\cdot\boldsymbol{\beta}$ (34)
+$\mathbf{W}$ $=\frac{(w-1)}{\|\boldsymbol{\beta}\|^{2}_{\mathbb{L}}}\cdot\boldsymbol{\beta}\boldsymbol{\beta}^{\top}+\mathbf{I}_{d}$ (35)
+Proof.
+Refer to Appendix II.G, Theorem VI.1. ∎
+{mymath}
+Theorem VI.2 (Manifold Preserving).
+Given the construction of $\mathbf{L}$ in Theorem VI.1, and scalars $a,b$, the following properties holds for any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{L}^{d_{1}}$.
   • 1)
@@
     $\langle L(\boldsymbol{x}),L(\boldsymbol{y})\rangle_{\mathbb{L}}=\langle\boldsymbol{x},\boldsymbol{y}\rangle_{\mathbb{L}}$,
-where $\langle\boldsymbol{x},\boldsymbol{y}\rangle_{\mathbb{L}}$ denotes the Minkowski inner product in Eq. (6).
+where $\langle\boldsymbol{x},\boldsymbol{y}\rangle_{\mathbb{L}}$ denotes the Minkowski inner product.
 Proof.
-Refer to Appendix II, Theorem VII.1. ∎
-Note that, the manifold-preserving property is given by the claim (2) in the theorem above.
+Refer to Appendix II.G, Theorem VI.2. ∎
+Note that, the manifold-preserving property is given by claim (2) in the theorem above.
 It is easy to verify that $G$ is closed on itself, i.e., $L_{i}\circ L_{j}\in G^{L}$, $\forall L_{i},L_{j}\in G^{L}$.
@@
 {mymath}
-Theorem VII.2 (Consistence Insurance).
-Given a set of points $\{\mathbf{z}_{i}\}$ in the Lorentz model of hyperbolic space, $\mathbf{z}_{i}\in\mathbb{L}^{\kappa,d_{1}}$ and their midpoint $\mathbf{z}\in\mathbb{L}^{\kappa,d_{1}}$ is constructed by Eq.
-(29), then $L(\mathbf{z})\in\mathbb{L}^{\kappa,d_{2}}$ is the midpoint of $\{L(\mathbf{z}_{i})\}$, $L(\mathbf{z}_{i})\in\mathbb{L}^{\kappa,d_{2}}$, if $L$ is a transformation in the Lorentz group as stated in Theorem VII.1.
+Theorem VI.3 (Consistence Insurance).
+Given a set of points $\{\boldsymbol{z}_{i}\}$ in the Lorentz model of hyperbolic space, $\boldsymbol{z}_{i}\in\mathbb{L}^{\kappa,d_{1}}$ and their midpoint $\boldsymbol{z}\in\mathbb{L}^{\kappa,d_{1}}$ is constructed by Eq.
+(25), then $L(\boldsymbol{z})\in\mathbb{L}^{\kappa,d_{2}}$ is the midpoint of $\{L(\boldsymbol{z}_{i})\}$, $L(\boldsymbol{z}_{i})\in\mathbb{L}^{\kappa,d_{2}}$, if $L$ is a transformation in the Lorentz group as stated in Theorem VI.2.
 Proof.
-Refer to Appendix II, Theorem VII.2. ∎
+Refer to Appendix II.H, Theorem VI.3. ∎
 The theorem above establishes the operation that propagates the action on tree nodes to their ancestors.
 In other words, only leaf contrast is needed in tree contrastive learning, and the consistence is maintained throughout the entire partitioning tree thanks to the properties of the proposed transformations.
-Closed-form Expression.
-Another advantage is that Eq.
-(36) induces the closed-form expression.
-Specifically, we parameterize $\mathbf{L}$ on $\mathbf{v}$, and the corresponding $\mathbf{W}$ is derived as follows,
-$\mathbf{W}=\begin{bmatrix}\mathbf{P}diag(\sqrt{\lambda_{1}},..,\sqrt{\lambda_{d+1}})\mathbf{P}^{T}\\ \mathbf{0}\end{bmatrix},\vspace{-0.05in}$ (37)
-where $\mathbf{P}$ and $\lambda$’s are obtained from eigenvalue decomposition of $\mathbf{V}=\mathbf{v}\mathbf{v}^{T}+diag(-1,1,...,1)$ which takes the form of
-$\mathbf{V}$ $=\mathbf{P}diag(\lambda_{1},..,\lambda_{d+1})\mathbf{P}^{T},\vspace{-0.05in}$ (38)
-with eigenvalues $\lambda_{1},..,\lambda_{d+1}$ sorted in descending order. $\mathbf{0}$ is the padding matrix of zeros for dimensionality consistence.
 Remark.
 We notice that recent studies propose a series of manifold-preserving operations in the Lorentz model of hyperbolic space [9, 57], but none of them satisfies consistence insurance and thereby cannot be utilized for the tree contrastive learning in hyperbolic space.
-Algorithm 1 Training IsoSEL
-0:  A weighted graph $G=(\mathcal{V},\mathcal{E},\mathbf{X})$ and the height of deep partitioning tree $H$.
-0:  The partitioning tree $\mathcal{T}$, and tree node embeddings $\{\mathbf{Z}^{h}\}_{h=1,...,H}$ at all the levels.
-1:  Generate the augmented graph using Eq. 25;
-2:  while not converged do
-3:     for both original and augmented graph do
-4:        Obtain leaf node embeddings $\mathbf{Z}^{H}=\operatorname{LConv}(\mathbf{X},\mathbf{A})$;
-5:        Compute the level-wise assignment and parent node embeddings recursively with Eqs. 31, 33 and 35;
-6:     end for
-7:     Forward leaf node embeddings into Lorentz projection head with Eq. 37;
-8:     Compute the overall objective $\mathcal{L}_{\operatorname{deeptree}}+\alpha\mathcal{L}_{\operatorname{contrast}}$;
+VI-B Tree Contrastive Loss
+In this part, we introduce the loss function for contrasting partitioning trees in the Lorentz model of hyperbolic space.
+For each node, the corresponding node in the counterpart view is treated as positive sample; otherwise, negative samples.
+Accordingly, Tree Contrastive Loss (TCL) is formulated as
+$\mathcal{L}_{\text{TCL}}(\mathbf{Z},\mathbf{Z^{\prime}})=\frac{1}{N}\sum_{i=1}^{N}\frac{\exp(d_{\mathbb{L}}(L(\boldsymbol{z_{i}}),L(\boldsymbol{z^{\prime}_{i}})))}{\sum_{j=1}^{N}w_{ij}\exp(d_{\mathbb{L}}(L(\boldsymbol{z_{i}}),L(\boldsymbol{z^{\prime}_{j}})))},\vskip-5.05942pt$ (36)
+where the reweighting factor is defined with the level-wise assignment, $w_{ij}=\sum_{k=1}^{N_{H-1}}S^{H}_{ik}S^{H}_{jk}$.
+That is, TCL considers the (hierarchical) tree branches in the partitioning trees, essentially differing itself from graph contrastive learning.
+Differences to Graph Contrastive Learning (GCL).
+GCL refines representations through exploring the similarity among all nodes in the graph.
+On the contrary, we contrast the partitioning trees of the graph—nodes and their parent hierarchies—under consistency insurance, so that we propose the Neural Lorentz Boost with linearity and isometry, maintaining the partitioning tree as a whole, and Tree Contrastive Loss, considering the partitioning structures.
+Also, GCL is typically conducted in Euclidean space, while we operate in hyperbolic space given its alignment to tree structures.
+Remaining Issues.
+Although tree contrastive learning refines node representations in hyperbolic space, it presents the following two limitations: On the one hand, view generation in this case is non-trivial.
+A straightforward way is to perform graph augmentation followed by LSEnet to generate a new hyperbolic view, but it may introduce noise that undermines model expressiveness as evidenced in Ablation Study.
+On the other hand, computing the loss function of Eq.
+(36) yields a quadratic complexity with respect to $|\mathcal{V}|$, the number of nodes, preventing the model scaling up to large graphs.
+VII Augmented Structural Information Learning
+In this section, we first elaborate on a theoretical insight that the structural entropy bounds contrastive learning, allowing for efficient and implicit tree contrastive learning.
+Integrating the theoretical establishment and neural net designs, we finally propose a novel deep graph clustering approach, Augmented Structural Information Learning (ASIL).
+VII-A Theoretical Insight: Connection between Structural Entropy and Contrastive Learning
+We uncover an inherent connection between structural entropy and contrastive learning, while simultaneously addressing the two aforementioned limitations.
+Specifically, we demonstrate that the objective of structural entropy aligns closely with that of contrastive learning in terms of graph conductance, and this connection is formally established by introducing a virtual graph constructed from node representations.
+First, prior to the mathematical formulation, we provide an intuitive discussion of their relationship: in hyperbolic space, the tree contrastive learning pulls together node representations belonging to the same partition (i.e., tree branch), while pushing apart those from different partitions; the structural entropy operates directly on the graph structure—namely, the adjacency matrix—and generates an optimal partitioning tree primarily based on the node connectivity.
+Thus, both structural entropy and contrastive learning explore the node similarity, but from structural and representational perspectives, respectively.
+Second, we introduce a virtual graph constructed from node representations to bridge these two viewpoints.
+In particular, we perform the Lorentz Boost Construction defined as follows,
+$\tilde{\mathbf{A}}_{ij}=\exp\left(-d_{\mathbb{L}}\left(L(\boldsymbol{z}_{i}),L(\boldsymbol{z}_{j})\right)/t\right),\vskip-5.78172pt$ (37)
+where $d_{\mathbb{L}}$ denotes the induced distance of Minkowski metric in the hyperbolic space, and $t$ is the temperature hyperparameter.
+The transformation $L$ corresponds to the Lorentz boost within the tree contrastive learning, and $\boldsymbol{z}$ refers to the optimized node representation.
+In practice, we sparsify the virtual adjacency matrix by retaining $k$-nearest neighbors regarding $d_{\mathbb{L}}$.
+An interesting finding is formally stated in the theorem below.
+{mymath}
+Theorem VII.1 (Connection between Structural Entropy and Tree Contrastive Learning).
+[…]
+Proof.
+Refer to Appendix II.I, Theorem VII.1. ∎
+The theorem above claims that structural entropy indeed bounds the tree contrastive learning defined in Eq.
+(36).
+The key point is Eq.
+(36) has been proven to align with minimizing the graph conductance, which is bounded by structural entropy as formalized in Theorem III.5.
+Our theoretical insight offers a promising alternative for tree contrastive learning, addressing its limitation in view generation and computational complexity.
+Concretely, it is preferable to perform structural information minimization over the virtual graph, which implicitly contrasts the partitioning trees in the hyperbolic space.
+The effect of contrasting two hyperbolic views is captured in the virtual graph of Lorentz boost construction in a parameterized fashion, and thus view generation is no longer required.
+As for computational complexity, we achieve the linear complexity regarding $|\mathcal{V}|$ with the reformulation in Eq.
+(17), compared to the quadratic complexity of contrastive learning in Eq.
+(36).
+VII-B Augmented Structural Entropy: Integrating Partitioning Tree Construction and Contrastive Learning
+This part presents our deep graph clustering approach —Augmented Structural Information Learning, referred to as ASIL.
+Based on the established theory of differential structural information and findings on its connection to tree contrastive learning, ASIL seamlessly integrates partitioning tree construction and (implicit) contrastive learning in the hyperbolic space, allowing for effective and efficient node clustering without the knowledge of the number of clusters.
+Specifically, ASIL first fuses the structural and representational views, and then feed them into LSEnet (Sec.
+V) to learn the hyperbolic partitioning tree for clustering.
+The overall architecture is sketched in Fig.
+1.
+The key design is a simple yet expressive clustering objective—Augmented Structural Entropy, which is formalized below
+$\min\nolimits_{\mathbf{\Theta}}\ \mathcal{H}^{\mathcal{T}_{\text{net}}}(G^{\gamma};\mathbf{Z};\mathbf{\Theta}),$ (38)
+where $G^{\gamma}$ is a fusion of the original graph and its representational view, the parameterized virtual graph (Sec.
+VII-A) that is constructed from node representations derived by Lorentz convolution in practice.
+Thus, the adjacency matrix is given with a balancing factor $\gamma$ as follows,
+$\mathbf{A}^{\gamma}=(1-\gamma)\mathbf{A}+\gamma\tilde{\mathbf{A}},\quad\gamma>0.$ (39)
+Note that the objective of Eq.
+(38) simultaneously learns an enhanced partitioning tree with a virtual graph, in which optimizing the virtual graph refines node representation via implicitly Lorentz tree contrastive learning (Sec.
+VI).
+Interestingly, this enhanced partitioning tree enjoys superior clustering capability to the deep tree of $\min\nolimits_{\mathbf{\Theta}}\ \mathcal{H}^{\mathcal{T}_{\text{net}}}(G;\mathbf{Z};\mathbf{\Theta})$, which is proved in light of graph conductance.
+{mymath}
+Theorem VII.2 (Graph Conductance Improvement with Graph Fusion).
+Given a graph $G=(\mathcal{V},\mathcal{E})$ and a virtual graph $\tilde{G}$ as in Theorem VII.1, the fused graph $G^{\gamma}$ is derived as in Eq.
+(39).
+For any nontrivial subset $\mathcal{S}\subset\mathcal{V}$ with $0<\mathrm{vol}_{A}(\mathcal{S})<\mathrm{vol}_{A}(V)$, denote the conductance of $\mathcal{S}$ under a graph $G$ as
+$\phi_{G}(\mathcal{S})=\frac{\mathrm{cut}_{G}(\mathcal{S},\mathcal{V}\setminus\mathcal{S})}{\min\{\mathrm{vol}_{G}(\mathcal{S}),\,\mathrm{vol}_{G}(\mathcal{V}\setminus\mathcal{S})\}},$
+where $\mathrm{cut}_{G}(\mathcal{S},\mathcal{P})=\sum_{i\in\mathcal{S},j\in\mathcal{P}}w_{ij}$ and $\mathrm{vol}_{G}(\mathcal{S})=\sum_{i\in\mathcal{S}}\sum_{j}w_{ij}$.
+Suppose that for every $\mathcal{S}\subset\mathcal{V}$,
+$\phi_{\tilde{G}}(\mathcal{S})\leq\phi_{G}(\mathcal{S}).$ (40)
+Then, the conductance of the fused graph satisfies
+$\phi_{G^{\alpha}}(\mathcal{S})\leq\phi_{G}(\mathcal{S}),\quad\forall\mathcal{S}\subset V.$
+Moreover, if there exists a subset $\mathcal{S}^{\star}$ such that $\phi_{\tilde{G}}(\mathcal{S}^{\star})<\phi_{G}(\mathcal{S}^{\star})$, then $\phi_{G^{\alpha}}(\mathcal{S}^{\star})<\phi_{G}(\mathcal{S}^{\star})$ for all $\alpha\in(0,1]$.
+Proof.
+Refer to Appendix II.I, Theorem VII.2. ∎
+Algorithm 1 Training ASIL
+0: A weighted graph $G=(\mathcal{V},\mathcal{E},\mathbf{X})$ and the height of deep partitioning tree $H$, fusion parameter $\gamma\in(0,1]$, nearset neighbors number $k$.
+0: The partitioning tree $\mathcal{T}$, and tree node embeddings $\{\mathbf{Z}_{\gamma}^{h}\}_{h=1,...,H}$ at all the levels.
+1:while not converged do
+2:  Obtain initial embeddings $\mathbf{Z}^{H}=\operatorname{LConv}(\mathbf{X},\mathbf{A})$;
+3:  Infer virtual graph $\tilde{\mathbf{A}}_{ij}=\exp(-d_{\mathbb{L}}(L(\boldsymbol{z}_{i}),L(\boldsymbol{z}_{j}))/\tau)$ with Lorentz boost $L$ in Eq. 32;
+4:  Sparsify the virtual graph with $k$-nearest neighbors;
+5:  Conduct graph fusion by $\mathbf{A}^{\gamma}=\mathbf{A}+\gamma\tilde{\mathbf{A}}$;
+6:  Obtain leaf node embeddings $\mathbf{Z}_{\gamma}^{H}=\operatorname{LConv}(\mathbf{X},\mathbf{A}^{\gamma})$;
+7:  Compute the level-wise assignment and parent node embeddings $\{\mathbf{Z}_{\gamma}^{h}\}_{h=1,...,H-1}$ recursively with Eqs. 27, 29 and 31;
+8:  Compute the overall objective $\mathcal{H}^{\mathcal{T}_{\text{net}}}(G^{\gamma};\mathbf{Z};\mathbf{\Theta})$;
 9:  Update the parameters with gradient-descent optimizer;
@@
 18: Return the partitioning tree $\mathcal{T}:=\lambda$.
+VII-C Algorithm: ASIL for Graph Clustering
+The overall procedure of ASIL is summarized in Algorithm 3.
+The first phase (Line 1-10) is to learn hyperbolic partitioning tree representation in a bottom-up manner.
+Specifically, in Line 2-4, a parameterized virtual graph first is constructed from node representations with a neural Lorentz boost, implicitly capturing tree contrastive learning.
+In Line 2-4, we forward a fused graph into LSEnet, recursively inducing parent node embeddings.
+In Line 8-9, the partitioning tree is optimized by a novel objective of augmented structural entropy.
+The second phase (Line 11-18) is to recover the partitioning tree structure through a top-down decoding process, which prunes empty leaves and resolves redundant chains in accordance with Theorem IV.2.
+Clustering structures are naturally uncovered in the partitioning tree, whose procedure are provided in Algorithm 2 given in Appendix III.
+Computational Complexity.
+First, we analyze the complexity of computing differentiable structural information.
+Thanks to the node-wise formulation in Eq.
+(17), this computation no longer requires enumerating all edges at each level.
+Instead, it operates solely on the coarsened graphs $\{G^{h}\}_{h=1}^{H}$, whose node counts $\{N_{h}\}$ satisfy $N_{h}\ll N_{h-1}$ and $N_{H}=N$.
+Second, given that the induced graph $G^{h-1}$ is sparse, $w^{h}_{kl}$ is computed via sparse matrix multiplication $\mathbf{W}^{h}=(\mathbf{S}^{h})^{\top}\mathbf{A}\mathbf{S}^{h}$, yielding the computational cost of $\mathcal{O}(|\mathcal{E}|)$ at leaf level, where $\mathcal{E}$ is the edge set of graph $G$.
+Third, we leverage $k$-nearest neighbors to construct the virtual adjacency matrix, which costs $\mathcal{O}(k|\mathcal{V}|)$, where $\mathcal{V}$ is the node set.
+Finally, the objective of augmented structural entropy regarding $\mathbf{A}^{\gamma}=(1-\gamma)\mathbf{A}+\gamma\tilde{\mathbf{A}}$ costs $\mathcal{O}(|\mathcal{E}|+k|\mathcal{V}|)$ in total.
+That is, the proposed ASIL discovers node clusters with a more preferable linear complexity in graph size, unknowing the number of clusters, while previous methods (e.g., MVGRL [21]) typically undergo a clustering procedure costing $\mathcal{O}(|\mathcal{N}|^{2})$.
+In summary, we present a hyperbolic deep model ASIL built on the established theory of differential structural entropy that allows for effective and efficient debiased clustering in linear complexity, without requiring the number of clusters.
 TABLE I: Clustering results on Cora, Citerseer, AMAP, and Computer, Pubmed datasets in terms of NMI and ARI (%).
@@
 NMI ARI NMI ARI NMI ARI NMI ARI NMI ARI
-DEC 23.54$\pm$1.13 15.13$\pm$0.72 28.34$\pm$0.42 28.12$\pm$0.24 37.35$\pm$0.84 18.29$\pm$0.64 38.56$\pm$0.24 34.76$\pm$0.35 22.44$\pm$0.14 19.55$\pm$0.13
-VGAE 43.40$\pm$1.62 37.50$\pm$2.13 32.70$\pm$0.30 33.10$\pm$0.55 66.01$\pm$0.84 56.24$\pm$0.22 37.62$\pm$0.24 22.16$\pm$0.16 30.61$\pm$1.71 30.15$\pm$1.23
-ARGA 48.10$\pm$0.45 44.10$\pm$0.28 35.10$\pm$0.58 34.60$\pm$0.48 58.36$\pm$1.02 44.18$\pm$0.85 37.21$\pm$0.58 26.28$\pm$1.38 24.80$\pm$0.17 24.35$\pm$0.17
-MVGRL 62.91$\pm$0.42 56.96$\pm$0.74 46.96$\pm$0.25 44.97$\pm$0.57 36.89$\pm$1.31 18.77$\pm$1.54 10.12$\pm$2.21     5.53$\pm$1.78 31.59$\pm$1.45 29.42$\pm$1.06
-Sublime 54.20$\pm$0.28 50.32$\pm$0.32 44.10$\pm$0.27 43.91$\pm$0.35     6.37$\pm$0.54     5.36$\pm$0.24 39.16$\pm$1.82 24.15$\pm$0.63 32.01$\pm$1.24 30.36$\pm$0.85
-ProGCL 41.02$\pm$1.64 30.71$\pm$1.38 39.59$\pm$1.58 36.16$\pm$2.21 39.56$\pm$0.25 34.18$\pm$0.58 35.50$\pm$2.06 26.08$\pm$1.91 30.24$\pm$1.21 28.98$\pm$1.06
-GRACE 57.30$\pm$0.86 52.70$\pm$1.20 39.90$\pm$2.26 37.70$\pm$1.35 53.46$\pm$1.32 42.74$\pm$1.57 47.90$\pm$1.65 36.40$\pm$1.56 30.08$\pm$0.24 27.60$\pm$1.24
-MCGC 44.90$\pm$1.56 37.80$\pm$1.24 40.14$\pm$1.44 38.00$\pm$0.85 61.54$\pm$0.29 52.10$\pm$0.27 53.17$\pm$1.29 39.02$\pm$0.53 33.39$\pm$0.02 29.25$\pm$0.01
-DCRN 48.86$\pm$0.85 43.79$\pm$0.48 45.86$\pm$0.35 47.64$\pm$0.30 73.70$\pm$1.24 63.69$\pm$0.84 OOM OOM 32.20$\pm$0.08 31.41$\pm$0.12
-FT-VGAE 61.03$\pm$0.52 58.22$\pm$1.27 44.50$\pm$0.13 46.71$\pm$0.75 69.76$\pm$1.06 59.30$\pm$0.81 51.36$\pm$0.92 40.07$\pm$2.13 33.05$\pm$0.24 35.50$\pm$0.12
-gCooL 58.33$\pm$0.24 56.87$\pm$1.03 47.29$\pm$0.10 46.78$\pm$1.51 63.21$\pm$0.09 52.40$\pm$0.11 47.42$\pm$1.76 27.71$\pm$2.28 30.82$\pm$0.15 31.48$\pm$0.34
-S3GC 58.90$\pm$1.81 54.40$\pm$2.52 44.12$\pm$0.90 44.80$\pm$0.65 59.78$\pm$0.45 56.13$\pm$0.58 54.80$\pm$1.22 29.93$\pm$0.22 33.32$\pm$0.34 34.50$\pm$0.57
-Congregate 63.16$\pm$0.71 59.27$\pm$1.23 50.92$\pm$1.58 47.59$\pm$1.60 70.99$\pm$0.67 60.55$\pm$1.36 46.03$\pm$0.47 38.57$\pm$1.05 35.03$\pm$0.23 33.42$\pm$0.47
-DinkNet 62.28$\pm$0.24 61.61$\pm$0.90 45.87$\pm$0.24 46.96$\pm$0.30 74.36$\pm$1.24 68.40$\pm$1.37 39.54$\pm$1.52 33.87$\pm$0.12 36.39$\pm$0.34 35.77$\pm$0.13
-GC-Flow 62.15$\pm$1.35 63.14$\pm$0.80 40.50$\pm$1.32 42.62$\pm$1.44 36.45$\pm$1.21 37.24$\pm$1.26 41.10$\pm$1.06 35.60$\pm$1.82 36.82$\pm$0.28 37.31$\pm$0.04
-GCSEE 53.02$\pm$0.41 51.22$\pm$0.88 35.66$\pm$0.42 36.29$\pm$0.37 64.15$\pm$0.68 56.76$\pm$1.24 39.03$\pm$1.58 35.24$\pm$1.56 26.88$\pm$0.21 26.15$\pm$0.24
-FastDGC 53.03$\pm$0.03 51.25$\pm$0.05 45.85$\pm$0.08 44.12$\pm$0.05 69.76$\pm$0.03 65.71$\pm$0.04 45.26$\pm$0.52 40.23$\pm$0.18 25.85$\pm$0.86 25.08$\pm$0.73
-MAGI 59.72$\pm$1.22 57.31$\pm$1.31 45.28$\pm$0.24 46.86$\pm$0.47 71.62$\pm$0.11 61.55$\pm$0.24 54.22$\pm$1.88 46.27$\pm$1.64 33.28$\pm$0.04 31.29$\pm$0.08
-RGC 57.60$\pm$1.36 50.46$\pm$1.72 45.70$\pm$0.29 45.47$\pm$0.43 47.65$\pm$0.91 42.65$\pm$1.53 46.24$\pm$1.05 36.12$\pm$0.30 28.01$\pm$1.85 25.56$\pm$1.91
-DGCluster 61.47$\pm$0.34 44.75$\pm$0.28 43.75$\pm$0.05 31.55$\pm$0.09 77.12$\pm$0.25 67.59$\pm$0.35 50.10$\pm$1.87 38.92$\pm$1.27 33.73$\pm$0.68 21.84$\pm$0.71
-IsoSEL 64.56$\pm$0.72 64.02$\pm$0.70 53.18$\pm$1.11 50.45$\pm$1.41 74.54$\pm$0.44 69.24$\pm$0.52 55.93$\pm$0.53 43.67$\pm$0.35 36.91$\pm$0.33 38.52$\pm$0.23
-Avg. Gain 3.51 8.58 7.70 5.45 5.82 8.91 1.36 1.60 1.89 7.77
-VII-B Hyperbolic Reweighing for Imbalance Clusters
-In deep clustering on imbalanced graphs, minority clusters of fewer nodes are typically underrepresented due to the limited number of samples, which often results in a biased boundary among imbalanced clusters.
-To address this issue, in Lorentz tree contrastive learning, we propose hyperbolic reweighed sampling to mitigate the bias in clustering boundary.
-The core idea of reweighed sampling is to refine the hyperbolic representations of the majority clusters in the pushing-pulling process, as the structural entropy has a promising ability to identify minority clusters, as shown in Theorem IV.3.
-To be specific, we first feed the original and augmented graphs into LSEnet to generate two hyperbolic views with their representations denoted as $\boldsymbol{z}$ and $\hat{\boldsymbol{z}}$, respectively.
-Second, we consider the contrast among the leaf nodes as the unity of the partitioning tree is ensured by the Lorentz projection head (Theorem VII.2).
-For each leaf node $\boldsymbol{z}^{H}_{i}$ in the original view, the positive sample is given as the corresponding $\hat{\boldsymbol{z}}^{H}_{i}$ in the augmented view.
-As for the negative samples, we draw $m$ leaf nodes of the same ancestor in the partitioning tree using a weighting function defined as follows,
-$w_{ij}=\exp\left(-d_{\mathbb{L}}(L(\boldsymbol{z}^{H}_{i}),L(\hat{\boldsymbol{z}}^{H}_{j}))\right).\vspace{-0.05in}$ (39)
-Accordingly, the Lorentz tree contrastive loss is formulated as
-where $\tau$ is the temperature to control the strictness for negative samples, and $L$ denotes the Lorentz projection head.
-Note that, the reweighed sampling alleviates the over aggregation of majority cluster representations while preserving the closeness of similar nodes in hyperbolic space.
-VII-C Overall Objective for Deep Graph Clustering
-Finally, the overall objective of IsoSEL is formulated as
-$\mathcal{J}=\mathcal{L}_{\operatorname{deeptree}}+\alpha\mathcal{L}_{\operatorname{contrast}},\vspace{-0.05in}$ (41)
-where $\mathcal{L}_{\operatorname{deeptree}}$ is defined by Eq.
-(17), and $\alpha$ denotes the weighting coefficient.
-The training procedure is described in Algorithm 1, where Lorentz tree contrastive learning refines the deep partitioning tree created by differentiable structural information in hyperbolic space to generate promising node clusters.
-Specifically, we first learn the partitioning tree representation in a bottom-up fashion, as outlined in Lines $3$-$6$ of Algorithm 1.
-Then, we recover the tree through a top-down process in Lines $10$-$17$, ensuring that empty leaf nodes are avoided and single chains are fixed in the relaxation.
-Further details of the recovery procedure are provided in Algorithm 2 given in Appendix III.
-Computational Complexity.
-In IsoSEL, the computational complexity of learning the deep partitioning tree is $\mathcal{O}(N)$ thanks to the node-wise formulation of differential structural entropy in Eq.
-(17).
-The Lorentz tree contrastive learning costs $\mathcal{O}(mN)$, where $N$ and $m$ denotes the number of nodes and negative samples, respectively.
-Accordingly, the overall complexity is yielded as $\mathcal{O}(N+mN)$, which is significantly less than that of our previous work of $\mathcal{O}(N|\mathcal{E}|)$ [20].
-Note that, though the number of clusters $K$ is no longer required, IsoSEL presents competitive complexity to typical deep clustering models of $\mathcal{O}(KN)$.
-In summary, IsoSEL takes the advantages of both hyperbolic geometry and differential structural entropy to effectively discover the cluster structure without $K$, while alleviating the clustering bias among imbalanced clusters.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Avg. Gain 11.55 15.16 12.42 10.25 18.20 20.52 12.84 12.82 6.95 10.02
 VIII Experiments
-In the experiment, we compare with $20$ strong baselines on $5$ commonly used datasets, aiming to answer the research questions as follows: RQ1 (Main Results): How does the proposed IsoSEL perform in terms of both effectiveness and efficiency?
-RQ2: How does IsoSEL outperform the classic structural entropy [16] and our previous model [20]?
-RQ3 (Ablation Study): How does each proposed module contribute to the success of IsoSEL, including the choice of hyperbolic space, augmentation strategy and tree contrastive learning?
-RQ4 (Hyperparameter Sensitivity): How do the hyperparameters affect the performance?
+In the experiment, we compare with $20$ strong baselines on $5$ commonly used datasets, aiming to answer the research questions as follows: RQ1 (Main Results): How does the proposed ASIL perform in terms of both effectiveness and efficiency, especially compared to vallina LSEnet?
+RQ2: How does ASIL outperform the classic structural entropy [27] and our previous model [54]?
+RQ3 (Ablation Study): How does each proposed module contribute to the success of ASIL, including the choice of hyperbolic space, Neural Lorentz Boost Construction and Implicit Tree Contrastive Learning?
+RQ4 (Hyperparameter Sensitivity): How do the hyperparameters affect the performance in each experimental datasets?
 RQ5: How do the imbalance clusters impact the clustering results?
-Besides, we examine the embedding expressiveness of IsoSEL, and include a visualization of the clusters learnt from IsoSEL.
+Besides, we examine the embedding expressiveness of ASIL, and include a visualization of the clusters learnt from ASIL.
 Additional results and further details are provided in Appendix VII.
@@
 TABLE II: Accuracy results on Cora, Citeseer, AMAP, Computer, and Pubmed datasets in terms of percentage (%). The best results are highlighted in boldface, and runner-ups are underlined.
-VGAE ARGA MVGRL Sublime GRACE DCRN gCooL S3GC DinkNet GC-Flow DGCluster IsoSEL Avg. Gain
-Cora 64.72 68.40 80.03 71.30 73.87 66.91 77.03 74.27 78.11 78.92 76.45 81.77 7.22
-Citeseer 61.00 60.72 61.23 68.53 63.12 70.86 71.35 68.80 70.36 70.06 70.14 71.37 6.26
-AMAP 74.26 69.28 41.07 47.22 67.66 77.94 68.52 75.15 79.23 72.54 78.92 78.68 7.51
-Computer 42.44 45.67 29.53 45.52 46.62 30.42 44.50 39.03 43.52 43.20 43.65 47.34 5.14
-Pubmed 68.31 64.82 67.01 66.62 63.72 69.63 68.24 71.30 69.42 73.42 66.39 73.53 6.26
-TABLE III: Comparison of time complexity and time cost for baselines.
-Method Time Complexity Inference Time (s)
-DGI $\mathcal{O}(ED+Nd^{2})$ 20.15
-MVGRL $\mathcal{O}(N^{2}d+Nd^{2})$ 176.35
-S3GC $\mathcal{O}(NSd^{2})$ 520.41
-Dink-Net $\mathcal{O}(BKd+K^{2}d+Bd)$ 37.32
-DGCluster $\mathcal{O}(K^{2}N+P^{2}N)$ 8.56
-IsoSEL (Ours) $\mathcal{O}(N+mN)$ 7.53
+VGAE ARGA MVGRL Sublime GRACE DCRN gCooL S3GC DinkNet GC-Flow DGCluster ASIL Avg. Gain
+Cora 64.72 68.40 80.03 71.30 73.87 66.91 77.03 74.27 78.11 78.92 76.45 82.14 8.50
+Citeseer 61.00 60.72 61.23 68.53 63.12 70.86 71.35 68.80 70.36 70.06 70.14 72.73 5.81
+AMAP 74.26 69.28 41.07 47.22 67.66 77.94 68.52 75.15 79.23 72.54 78.92 79.31 10.97
+Computer 42.44 45.67 29.53 45.52 46.62 30.42 44.50 39.03 43.52 43.20 43.65 49.03 7.75
+Pubmed 68.31 64.82 67.01 66.62 63.72 69.63 68.24 71.30 69.42 73.42 66.39 74.47 6.39
 VIII-A2 Evaluation Metrics
@@
 VIII-A3 Model Configuration & Reproducibility
-We instantiate the proposed IsoSEL in the Lorentz model of hyperbolic space with a standard curvature of $\kappa=-1$.
+We instantiate the proposed ASIL in the Lorentz model of hyperbolic space with a standard curvature of $\kappa=-1$.
 The model configuration by default is specified as follows.
 We consider the partitioning tree of height $2$ by default and hyperparameter $N_{1}$ of network architecture is set as $10$.
-The augmentation ratio $\alpha$ is $0.2$.
-We stack the Lorentz convolution once in LSEnet, and set the number of negative samples to $30$ in Lorentz tree contrastive learning.
-The parameters are optimized by Adam and further details are provided in implementation note of Appendix VI.
-As for the baselines, the hyperparameter setting is the same as that of original papers.
+The augmentation ratio $\gamma$ is $0.01$, the $k$ nearest neighbors for sparsifying $\tilde{\mathbf{A}}$ is set to $8$.
+We stack the Lorentz convolution once in LSEnet, The parameters are optimized by Adam, the further details and the hyperparameter of the baselines are provided in implementation note of Appendix VI.
 VIII-B Main Results of Node Clustering
@@
 We report the clustering results in terms of NMI and ARI on the five real datasets in Table I, and ACC in Table II.
-Deep clustering models require cluster number $K$ in training phase, except DGCluster, RGC and our IsoSEL.
+Deep clustering models require cluster number $K$ in training phase, except DGCluster, RGC and our ASIL.
 For clustering models without $K$, DGCluster conducts self-supervised representation learning with modularity [3], and RGC searches the optimal $K$ with reinforcement learning [34].
-For Riemannian models, i.e., Congregate and IsoSEL, we project Euclidean attributes onto the respective manifold, and the projection is given by exponential map with the reference point of the origin.
+For Riemannian models, i.e., Congregate and ASIL, we project Euclidean attributes onto the respective manifold, and the projection is given by exponential map with the reference point of the origin.
 DCRN does not scale well to the larger Computer dataset, and runs into Out-Of-Memory (OOM) error on our hardware.
-In Table I, the proposed IsoSEL has the best clustering results except two cases, NMI on AMAP and ARI on Computer.
-Note that, IsoSEL outperforms $14$ recent deep clustering models by $+1.3\%$ NMI on average, and generally achieves the best ACC in Table II.
-Also, we observe that IsoSEL shows more consistent performance over different datasets, e.g., DGCluster obtain the best NMI on AMAP but is unsatisfactory on Computer.
+In Table I, the proposed ASIL has the best clustering results except two cases, NMI on AMAP and ARI on Computer.
+Note that, ASIL outperforms $20$ recent deep clustering models by an average of $12.42\%$ NMI on Citeseer, and generally achieves the best ACC in Table II.
+Also, we observe that ASIL shows more consistent performance over different datasets, e.g., DGCluster obtain the best NMI on AMAP but is unsatisfactory on Computer.
 VIII-B2 Efficiency
-Table III collects the theoretical time complexity and empirical inference time on Cora.
-The inference time is given in seconds, and we are presented as the fastest deep model in the table.
-As for theoretical complexity, $N$ and $E$ denote the number of nodes and edges, respectively.
+Table IV collects the theoretical time complexity and empirical running time on Citeseer dataset.
+The running time is given in seconds, and we are presented as the fastest deep model in the table.
+As for theoretical complexity, $N$ and $|E|$ denote the number of nodes and edges, respectively.
+$d$ is the hidden dimension of models.
 $B$ is the batch size of DinkNet.
@@
 In DGCluster, $P$ is a dimensionality of auxiliary matrix, $P\ll N$.
-In IsoSEL, $m$ is the number of negative samples, and it yields the linear complexity thanks to the reformulation of differential structural information and hyperbolic reweighing of Lorentz tree contrastive learning.
+In ASIL, $k$ is the number of nearest neighbors for sparsifying $\tilde{A}$, and it yields the linear complexity thanks to the reformulation of differential structural information and adjacency sparsifying.
 It is noteworthy to mention that, compared to recent deep clustering model such as S3GC, we achieve superior clustering effectiveness while removing the prior knowledge of cluster number without computational overhead.
-The following parts will demonstrate how the differential formulation, hyperbolic space and Lorentz tree contrastive learning contribute to the success of the proposed model.
-TABLE IV: Ablation study results of different variants of IsoSEL on graph clustering tasks in terms of clustering performance (%).
-Cora Citeseer AMAP Photo Pubmed
+The following parts will demonstrate how the differential formulation, hyperbolic space and neural Loretnz boost contribute to the success of the proposed model.
+TABLE III: Comparison to LSEnet and CSE on Cora, Citerseer, Computer, AMAP and PubMed datasets in terms of NMI and ARI (%).
+Cora Citeseer Computer AMAP PubMed
 NMI ARI NMI ARI NMI ARI NMI ARI NMI ARI
-LSEnet 63.97$\pm$0.67 63.35$\pm$0.56 52.26$\pm$1.09 48.01$\pm$1.25 71.12$\pm$1.30 65.08$\pm$0.73 55.03$\pm$0.79 42.15$\pm$1.02 34.25$\pm$0.24 35.16$\pm$0.30
-w/o CIncorp 63.45$\pm$0.92 62.78$\pm$0.88 51.96$\pm$1.45 48.27$\pm$1.65 72.65$\pm$1.92 64.37$\pm$1.35 54.78$\pm$1.03 42.86$\pm$1.62 36.23$\pm$0.50 37.12$\pm$0.69
-w/o IsoAug 62.89$\pm$1.05 62.15$\pm$0.97 51.42$\pm$1.58 47.93$\pm$1.80 72.23$\pm$2.01 64.01$\pm$1.50 54.34$\pm$1.12 42.54$\pm$1.79 35.97$\pm$0.64 36.85$\pm$0.82
-w/o TRefine 62.34$\pm$1.20 61.78$\pm$1.10 50.87$\pm$1.70 47.42$\pm$1.95 71.89$\pm$2.14 63.78$\pm$1.68 53.98$\pm$1.25 42.21$\pm$1.98 35.78$\pm$0.81 36.64$\pm$0.95
-w/o LGroup 61.78$\pm$1.32 61.23$\pm$1.18 50.31$\pm$1.85 46.89$\pm$2.10 71.54$\pm$2.28 63.45$\pm$1.80 53.56$\pm$1.38 41.89$\pm$2.12 35.56$\pm$0.95 36.42$\pm$1.12
-IsoSEL 64.56$\pm$0.72 64.02$\pm$0.70 53.18$\pm$1.11 50.45$\pm$1.41 74.54$\pm$0.44 69.24$\pm$0.52 55.93$\pm$0.53 43.67$\pm$0.35 36.91$\pm$0.33 38.52$\pm$0.23
-TABLE V: Comparison between LSEnet and CSE on Cora, Citerseer, Football and Karate datasets in terms of NMI and ARI (%).
-Cora Citeseer Football Karate
-NMI ARI NMI ARI NMI ARI NMI ARI
-CSE 60.82$\pm$0.30 59.36$\pm$0.17 47.12$\pm$1.01 45.67$\pm$0.67 79.23$\pm$0.12 54.06$\pm$0.05 81.92$\pm$0.29 69.77$\pm$0.02
-IsoSEL 64.56$\pm$0.72 64.02$\pm$0.70 53.18$\pm$1.11 50.45$\pm$1.41 82.35$\pm$0.53 57.36$\pm$0.35 83.28$\pm$0.72 74.15$\pm$0.55
-Performance Gain 3.74 4.66 6.06 4.78 3.12 3.30 1.36 4.38
-Refer to captionRefer to caption
-Figure 2: Parameter sensitivity on dimension of structural entropy.
+[…]
+[…]
+[…]
+Average Gain 3.39 3.37 5.32 4.93 4.69 4.44 1.28 3.62 3.88 4.58
+TABLE IV: Comparison of time complexity and time cost for baselines.
+Method Time Complexity Running Time (s)
+DGI $\mathcal{O}(|E|d+Nd^{2})$ 100.7
+MVGRL $\mathcal{O}(N^{2}d+Nd^{2})$ 829.8
+S3GC $\mathcal{O}(NSd^{2})$ 1247.2
+Dink-Net $\mathcal{O}(BKd+K^{2}d+Bd)$ 217.7
+DGCluster $\mathcal{O}(K^{2}Nd+P^{2}Nd^{2})$ 79.3
+ASIL (Ours) $\mathcal{O}(|E|d+kNd^{2})$ 22.1
 VIII-C Comparison with Classic Structural Entropy [27]
-In this section, we analyze the strength of the proposed IsoSEL against the classic structural entropy [16] (referred to as CSE for short).
+In this section, we analyze the strength of the proposed ASIL against the classic structural entropy [27] (referred to as CSE for short), and LSEnet [55].
 (1) Time Complexity.
 We report the running time under different heights of partitioning trees in Fig 2.
-IsoSEL shows better scalability than CSE.
-Concretely, CSE has competitive time cost to IsoSEL on the small dataset (Football), but performs undesirably on larger datasets, e.g., at least 13×13$\times$13 × time cost to IsoSEL on Cora.
-In addition, the linear complexity of IsoSEL is regardless of the height of partitioning tree, while CSE scales exponentially over the height, e.g., $\mathcal{O}(N^{3}\log^{2}N)$ for height $3$, making it prohibited on large graphs.
+ASIL shows better scalability than CSE.
+Concretely, CSE has competitive time cost to ASIL on the small dataset (Football, right image of Fig.
+2), but performs undesirably on larger datasets, e.g., an average of $30\times$ time cost to ASIL on Cora (left image of Fig.
+2).
+In addition, the linear complexity of ASIL is regardless of the height of partitioning tree, while CSE scales exponentially over the height, e.g., $\mathcal{O}(N^{3}\log^{2}N)$ for height $3$, making it prohibited on large graphs.
 (2) Clustering Effectiveness.
 Note that, CSE itself neither generates cluster assignment nor node representations.
-We invoke the heuristic algorithm to construct a partitioning tree of height $2$, where the leaf-parent relationship is regarded as cluster assignment, and thus we instantiate IsoSEL to generate a deep tree of the same height for fair comparison.
-We summarize the results on Cora, Citeseer, Football and Karate in Table V.
-IsoSEL consistently outperforms CSE for all the cases.
-It suggests that cluster assignment is improved by the hyperbolic geometry and attribute incorporation of IsoSEL.
+We invoke the heuristic algorithm to construct a partitioning tree of height $2$, where the leaf-parent relationship is regarded as cluster assignment, and thus we instantiate ASIL to generate a deep tree of the same height for fair comparison.
+We summarize the results on Cora, Citeseer, AMAP, Computers and PubMed in Table III.
+ASIL consistently outperforms CSE and LSEnet for all the cases.
+It suggests that cluster assignment is improved by the hyperbolic geometry, attribute incorporation and neural Lorentz boost construction of ASIL.
+Refer to captionRefer to caption
+Figure 2: Running time comparison between ASIL, LSEnet and CSE.
 VIII-D Comparison with Our Previous LSEent [54]
 This subsection demonstrates the significant improvement compared to LSEnet in the conference version.
-LSEnet constructs the deep partitioning tree in hyperbolic space, while IsoSEL leverages Lorentz tree contrastive learning to refine the deep tree of LSEnet.
-First, we compare the clustering effectiveness and report the results on Cora, Citeseer, AMAP, Photo and Pubmed datasets in Table IV.
-IsoSEL consistently achieves better performance in both NMI and ARI.
-The performance gain is attributed to hyperbolic tree representation refinement and attribute incorporation via Lorentz tree contrastive learning.
+LSEnet constructs the deep partitioning tree in hyperbolic space, while ASIL leverages neural Lorentz boost construction to refine the deep tree of LSEnet.
+First, we compare the clustering effectiveness and report the results on Cora, Citeseer, AMAP, Photo and Pubmed datasets in Table V.
+ASIL consistently achieves better performance in both NMI and ARI.
+The performance gain is attributed to hyperbolic tree representation refinement and attribute incorporation via Neural Lorentz boost construction.
 Second, we examine the running time under different tree heights in Fig.
 2.
-IsoSEL outperforms LSEnet on both datasets owing to the reduced complexity in the proposed reformulation and the sampling strategy of contrastive learning.
-In other words, IsoSEL in this paper exhibits superior clustering ability to the previous LSEnet, with minimal computational overhead.
+ASIL outperforms LSEnet on both datasets owing to the reduced complexity in the proposed reformulation and the sparsifying process in the neural Lorentz boost construction.
+That is, ASIL significantly outperforms LSEnet in both effectiveness and efficiency on the benchmark datasets.
 (The superiority in expressiveness and imbalanced clusters is elaborated in Sec.
@@
 VIII-H, respectively.)
-TABLE VI: Ablation study results of different representation space of partitioning tree of IsoSEL.
+Figure 3: Sensitivity analysis on model hyperparameters $\gamma$, $k$ and $H$.
+TABLE V: Ablation study results of different variants of ASIL on graph clustering tasks in terms of clustering performance (%).
+Cora Citeseer AMAP Computers PubMed
+NMI ARI NMI ARI NMI ARI NMI ARI NMI ARI
+[…]
+[…]
+[…]
+[…]
+TABLE VI: Ablation study results of different representation space of partitioning tree of ASIL.
 Cora Citeseer PubMed
 NMI ARI NMI ARI NMI ARI
-Hyperbolic 64.56 64.02 53.18 50.45 36.91 38.52
-Hyperspherical 50.53 51.23 43.53 40.09 27.35 27.26
-Euclidean 56.33 53.78 48.92 46.42 30.07 30.19
+Hyperbolic 65.78 64.72 55.01 51.77 38.13 39.74
+Hyperspherical 52.39 51.75 44.93 40.77 28.05 28.17
+Euclidean 58.39 55.77 49.72 47.92 31.51 31.69
 VIII-E Ablation Study
-This subsection introduce several variants to examine the effectiveness of each proposed module of IsoSEL, and comparison results are summarized in Table IV.
-VIII-E1 Contrastive Incorporation vs Input Incorporation
-We discuss the strategy for incorporating attribute information into partitioning tree.
-In IsoSEL, we propose a contrastive approach that incorporates attribute information through mutual information maximization between the original and augmented graphs.
-Correspondingly, a variant named $\operatorname{w/oCIncorp}$ is introduced to perform input incorporation by directly constructing the partitioning tree from the augmented graph with node attributes.
-As shown in Table IV, IsoSEL consistently outperforms $\operatorname{w/oCIncorp}$ in all cases.
-The superiority of contrastive incorporation lies in its ability to not only inject the attribute information but also to refine the tree representations in hyperbolic space, thereby improving cluster assignment.
-VIII-E2 Isometric Augmentation vs Weighted-Sum Augmentation
-We study how to effectively integrate node attributes into graph structure in the augmented view.
-In IsoSEL, we derive the isometric augmentation in Eq.
-(24) which conducts the integration from the representation perspective with the notion of isometry.
-To evaluate the effectiveness of the proposed formulation, we introduce the $\operatorname{w/oIsoAug}$ variant that generates the augmented structure by a weighted sum of original structure and attribute similarity.
-In Table IV, we find that the weighted-sum augmentation of $\operatorname{w/oIsoAug}$ leads to inferior performance, demonstrating the effectiveness of isometric augmentation.
-VIII-E3 Representation Refinement by Tree Contrastive Learning
-This part verifies that contrastive learning itself contributes to the success of clustering.
-To isolate the impact of attribute information, we introduce the $\operatorname{w/oTRefine}$variant, which is a degenerated version of IsoSEL.
-Specifically, we replace isometric augmentation with random structure perturbation, where the augmented view is generated by randomly adding or removing edges from the original structure.
-In Table IV, we find that $\operatorname{w/oTRefine}$ generally achieves a performance gain compared to LSEnet.
-The contrastive learning refines the partitioning tree representation in hyperbolic space, which is important to clustering performance.
-VIII-E4 Neural Lorentz Group for Tree Contrastive Learning
-To investigate its effectiveness, we design a variant named $\operatorname{w/oLGroup}$, which replaces proposed Lorentz projection head of neural Lorentz group by the Lorentz layer in [21].
-Specifically, the Lorentz layer preserves the manifold structure in the Lorentz model but lacks the linearity of the proposed transformation.
-We emphasize that linearity ensures consistency between leaf and parent nodes in the partitioning tree, as elaborated in Theorem VII.2.
-In Table IV, IsoSEL achieves better clustering results than $\operatorname{w/oLGroup}$, demonstrating the importance of consistency in tree contrastive learning.
-VIII-E5 Representation Space of Partitioning Tree
-We conduct a geometric ablation to study the alignment between hyperbolic space and the partitioning tree.
-Note that, we are interested in the promising space of the deep tree, while the underlying manifold of input graph is orthogonal to our focus.
-Theoretically, Theorem VI.1 shows the connection between the tree and hyperbolic space.
-Empirically, we instantiate the proposed IsoSEL in hyperbolic, hyperspherical and Euclidean space without loss of generality.
-The clustering results on Cora, Citeseer and Pubmed are given in Table VI.
-These results show that hyperbolic space is well suited to model the partitioning tree and achieves the best performance among constant curvature spaces.
-Figure 3: Sensitivity Analysis of $\alpha$
-(a) NMI
-(b) ARI
-Figure 4: Parameter sensitivity on dimension of structural entropy.
-VIII-F Hyperparameter Sensitivity
-In this subsection, we discuss the hyperparameter sensitivity of the proposed IsoSEL: attribute augmentation ratio $\alpha$ and height of partitioning tree $H$.
-VIII-F1 Attribute Augmentation Ratio
-In the proposed isometric augmentation, the hyperparameter $\alpha$ controls the relative importance over attribute and structural information.
-To evaluate the sensitivity, we vary $\alpha$ in the range of $[0.2,0.8]$, while fixing other hyperparameters.
-We report the corresponding performance on Cora, Citeseer, AMAP and Computer datasets in Fig.
-3.
-Note that, only slight changes of clustering results is observed under different ratios, and IsoSEL exhibits robustness regarding $\alpha$.
-VIII-F2 Height of Partitioning Tree
-The hyperparameter $H$ determines the height of the partitioning tree.
-We collect the clustering results with different tree heights in Fig 4, where $H$ takes the values in $\{2,3,4,5\}$.
-There are two findings from the empirical results: First, the height of partitioning tree affects the performance of IsoSEL.
-Second, we notice that IsoSEL achieves the best results on Cora at the height of $2$ while the optimal height of AMAP is $3$.
-We argue that the optimal height is depend on the inherent complexity of the graph.
-VIII-G Embedding Expressiveness of IsoSEL
-IsoSEL generates cluster assignment as well as node embedding in hyperbolic space.
-Here, we evaluate the embedding expressiveness with link prediction.
-To this end, IsoSEL is compared with the Euclidean models (including GCN [3], SAGE [50], GAT [4] and DGI [51]), hyperbolic models (including HGCN [52] and LGCN [53]), a recent $Q$GCN in ultra hyperbolic space [54] and our previous LSEnet.
-Popular metrics of AUC (Area Under the Curve) and AP (Average Precision) are employed to quantify the results.
-The link prediction performances on the five datasets are collected in Table VII, where the height of partitioning tree is set as $3$ for both LSEnet and IsoSEL.
-As shown in Table IV, Riemannian models generally outperforms the Euclidean counterparts, and IsoSEL achieves the state-of-the-art results.
-It suggests that Riemannian manifolds better model the complexity of graph structures and, more importantly, the partitioning tree encodes fruitful information for link prediction.
-Also, compared to LSEnet, IsoSEL consistently obtains the performance gain thanks to the proposed Lorentz tree contrastive learning.
+This subsection introduces several variants to examine the effectiveness of each proposed module of ASIL, and comparison results are summarized in Table V.
+VIII-E1 Importance of fusing a virtual graph
+We examine the performance gain achieved by fusing a virtual graph $\mathbf{A}^{\gamma}$ constructed from hyperbolic node representations.
+To this end, we introduce a variant that disables the construction of $\tilde{\mathbf{A}}$ and instead uses only the original topology $\mathbf{A}$, which indeed reduces to LSEnet.
+ASIL significantly outperforms LSEnet on all datasets, particularly on AMAP that with imbalanced classes, where semantic are crucial for distinguishing topological noise.
+This validates that the inclusion of a virtual enables ASIL to discover clusters enhanced by both structural and semantic perspectives.
+VIII-E2 Importance of consistency insurance, achieved by neural Lorentz boost
+To investigate the role of consistency insurance in tree contrastive learning, we design a variant named w/oLBoost, which replaces our neural Lorentz group layer with the standard Lorentz linear layer from [9].
+While the baseline layer preserves the manifold structure, it lacks the isometric linearity that ensures geometric consistency between leaf and internal nodes in the partitioning tree, as formalized in Theorem VI.3.
+Table V shows that ASIL achieves significantly better clustering than w/oLBoost, demonstrating that consistency insurance are critical for high-fidelity tree contrastive learning.
+VIII-E3 Implicit or explicit tree contrastive learning
+In this part, we demonstrate that the implicit tree contrastive learning by a parameterized virtual graph outperforms explicit tree contrastive learning.
+To this end, we give another variant named as ExplicitTCL.
+Specifically, we use an fixed isometric augmentation (details in Appendix X) to generate an augmentation view of original graph, and then use neural Lorentz boost to compute tree contrastive loss explicitly as in Eq.
+36.
+Thus, the total loss is given as $\mathcal{H}^{\mathcal{T}_{\text{net}}}(G^{\gamma})+\lambda\mathcal{L}_{\text{TCL}}$, where $\lambda=0.01$ is a coefficient.
+The results are shown in the third line of Table V.
+It shows that the proposed implicit approach consistently achieves better performance on the five datasets, which verifies the effectiveness of augmented structural entropy.
+VIII-E4 Representation Space of the Partitioning Tree
+We conduct a geometric ablation to study the alignment between the underlying manifold and the hierarchical partitioning tree.
+Note that our focus is on the geometry of the learned tree, not the input graph manifold.
+Theoretically, Theorem V.1 establishes the low-distortion embeddability of trees in hyperbolic space.
+Empirically, we instantiate ASIL in three constant-curvature spaces: hyperbolic ($\kappa<0$), hyperspherical ($\kappa>0$), and Euclidean ($\kappa=0$), keeping all other components identical.
+The clustering results on Cora, Citeseer, and Pubmed are given in Table III.
+These results confirm that hyperbolic space is uniquely suited for modeling hierarchical partitioning trees, consistently achieving the best performance across datasets.
 TABLE VII: Embedding expressiveness regarding link prediction on Cora, Citeseer, AMAP, Computer, and Pubmed datasets. The best results are highlighted in boldface, and runner-ups are underlined.
@@
 Model AUC AP AUC AP AUC AP AUC AP AUC AP
-GCN 91.19$\pm$0.51 91.85$\pm$0.43 90.16$\pm$0.49 91.91$\pm$0.24 90.12$\pm$0.72 91.08$\pm$0.08 93.86$\pm$0.36 93.24$\pm$0.32 91.16$\pm$0.14 89.96$\pm$0.21
-SAGE 86.02$\pm$0.55 88.24$\pm$0.87 88.18$\pm$0.22 89.15$\pm$0.36 98.02$\pm$0.32 97.51$\pm$0.21 92.82$\pm$0.20 93.12$\pm$0.14 89.22$\pm$0.87 89.44$\pm$0.82
-GAT 92.55$\pm$0.49 90.27$\pm$0.23 89.32$\pm$0.36 90.08$\pm$0.06 98.67$\pm$0.08 97.21$\pm$0.34 95.93$\pm$0.19 94.57$\pm$0.24 91.60$\pm$0.37 90.37$\pm$0.21
-DGI 90.02$\pm$0.80 90.61$\pm$1.00 94.53$\pm$0.40 95.72$\pm$0.10 98.24$\pm$0.13 96.84$\pm$0.24 95.84$\pm$0.05 95.22$\pm$0.36 91.24$\pm$0.60 92.23$\pm$0.50
-HGCN 93.60$\pm$0.37 93.64$\pm$0.12 94.39$\pm$0.42 94.38$\pm$0.11 98.06$\pm$0.29 96.84$\pm$0.15 96.88$\pm$0.53 96.18$\pm$0.13 96.30$\pm$0.28 94.72$\pm$0.52
-LGCN 92.69$\pm$0.26 91.09$\pm$0.37 93.49$\pm$1.11 94.96$\pm$0.18 97.08$\pm$0.08 96.73$\pm$0.52 96.37$\pm$0.70 95.66$\pm$0.24 95.40$\pm$0.23 93.12$\pm$0.24
-$Q$GCN 95.22$\pm$0.29 94.12$\pm$0.51 94.31$\pm$0.73 94.08$\pm$0.34 95.17$\pm$0.45 93.94$\pm$0.18 95.10$\pm$0.03 94.28$\pm$0.33 96.20$\pm$0.34 94.24$\pm$0.31
-LSEnet 95.51$\pm$0.60 95.40$\pm$0.04 94.42$\pm$1.51 94.69$\pm$0.05 98.75$\pm$0.67 97.04$\pm$0.13 97.06$\pm$1.02 97.06$\pm$0.05 96.41$\pm$0.10 95.46$\pm$0.10
-IsoSEL 96.23$\pm$0.14 95.94$\pm$0.21 95.02$\pm$0.34 95.81$\pm$0.29 98.78$\pm$0.19 97.26$\pm$0.17 97.15$\pm$0.34 97.34$\pm$0.38 96.50$\pm$0.07 95.72$\pm$0.09
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+VIII-F Hyperparameter Sensitivity
+In this subsection, we analyze the sensitivity of ASIL to three key hyperparameters: (i) the number of neighbors $k$ for sparsifying $\tilde{\mathbf{A}}$, (ii) the coefficient $\gamma$ of Lorentz boost construction $\mathbf{A}^{\gamma}=(1-\gamma)\mathbf{A}+\gamma\tilde{\mathbf{A}}$, and (iii) the height $H$ of the partitioning tree (i.e., the dimension of structural entropy).
+VIII-F1 Coefficient for Lorentz Boost Construction
+The coefficient $\gamma$ controls the importance of representational perspective in the graph fusion.
+We test $\gamma\in\{0.01,0.05,0.1,0.5,1.0,2.0\}$.
+When $\gamma=0$, ASIL degenerates to the original LSEnet.
+As shown in Figure 3(a–e), when $\gamma$ gets larger from $0.5$ to $1.0$, the performance of ASIL gets worse sharply since too much information injection may be treated as noisy structure, especially when $\gamma=2.0$ that exceeds the scope defined in Theorem VII.2.
+VIII-F2 $k$-NN Neighborhood Size ($k$)
+The virtual graph $\tilde{\mathbf{A}}$ is constructed via $k$-nearest neighbors in Lorentz space.
+We vary $k\in\{5,8,10,15,30,60\}$ and fix other hyperparameters.
+As shown in Figure 3(f–j), a small $k$ (e.g., $k=5,8,10$) achieves well performance while a large $k$ (e.g., $k=30,60$) introduces noisy cross-cluster edges, destroying clustering effectiveness.
+VIII-F3 Height of Partitioning Tree ($H$)
+The hyperparameter $H$ determines the depth of the hierarchical partitioning tree.
+We evaluate $H\in\{2,3,4,5\}$ and report results in Figure 3(k–o).
+We observe that performance is sensitive to $H$, with optimal values depending on graph complexity: ASIL achieves peak NMI on Cora at $H=2$, while AMAP prefers $H=3$.
+This suggests that deeper trees benefit more complex graphs with hierarchical community structures.
+Notably, ASIL remains stable across a wide range of $H$, demonstrating robustness to model capacity.
+In summary, ASIL is generally robust to hyperparameter choices, and default values ($k=8$, $\alpha=0.01$, $H=2$) work well across datasets.
+VIII-G Embedding Expressiveness of ASIL
+ASIL generates cluster assignment as well as node embedding in hyperbolic space.
+Here, we evaluate the embedding expressiveness with link prediction.
+To this end, ASIL is compared with the Euclidean models (including GCN [25], SAGE [20], GAT [60] and DGI [61]), hyperbolic models (including HGCN [8] and LGCN [76]), a recent QQGCN in ultra hyperbolic space [67] and our previous LSEnet.
+Popular metrics of AUC (Area Under the Curve) and AP (Average Precision) are employed to quantify the results.
+The link prediction performances on the five datasets are collected in Table VII, where the height of partitioning tree is set as $3$ for both LSEnet and ASIL.
+As shown in Table VII, Riemannian models generally outperforms the Euclidean counterparts, and ASIL achieves the state-of-the-art results.
+It suggests that Riemannian manifolds better model the complexity of graph structures and, more importantly, the partitioning tree encodes fruitful information for link prediction.
+Also, compared to LSEnet, ASIL consistently obtains the performance gain thanks to the implicit Lorentz tree contrastive learning.
 (a) gCool: true
@@
 (d) LSEnet: true
-(e) IsoSEL: true
+(e) ASIL: true
 (f) gCool: prediction
@@
 (i) LSEnet: prediction
-(j) IsoSEL:prediction
-Figure 5: Visualization of real and predicted data.
+(j) ASIL:prediction
+Figure 4: Visualization of real and predicted data on Photo dataset.
 VIII-H Case Study on Imbalanced Clusters
@@
 Real graphs often exhibit skewed class distributions, and for this case study, we use the Photo dataset.
-The proportion of $8$ clusters in Photo is $25\%$, $22\%$, $11\%$, $11\%$, $10\%$, $9\%$, $4\%$ and $4\%$.
-We compare IsoSEL with gCool, DGCluster, RGC and LSEnet, and visualize the node embedding in Fig.
+The proportion of $8$ clusters in Photo is $5\%, 22$$2\%$, $\%, 1111\%$, $10\%$, $9\%$, $4\%$ and $4\%$.
+We compare ASIL with gCool, DGCluster, RGC and LSEnet, and visualize the node embedding in Fig.
 4 where different clusters are denoted by different colors.
 The distributions of true labels and predicted labels are given in the top and bottom subfigures, respectively.
-We invoke the normal t-sne for gCool, DGCluster, RGC, while employing the hyperbolic metric for LSEnet and IsoSEL, so that all the visualizations are finally presented in 2D Euclidean space.
+We invoke the normal t-sne for gCool, DGCluster, RGC, while employing the hyperbolic metric for LSEnet and ASIL, so that all the visualizations are finally presented in 2D Euclidean space.
 In Fig.
-5, we obtain three findings as follows.
-(i) Clustering model without $K$ may fail to uncover the correct cluster number, e.g., RGC outputs $7$ node clusters on Photo dataset, but IsoSEL gives the true cluster number $8$.
+4, we obtain three findings as follows.
+(i) Clustering model without $K$ may fail to uncover the correct cluster number, e.g., RGC outputs $7$ node clusters on Photo dataset, but ASIL gives the true cluster number $8$.
 (ii) We observe that deep clustering models tend to generate biased clustering boundary.
 For example, in RGC, the members of violet cluster and the minority gray cluster (at the upper left in Fig.
-5c) are assimilated into the majority cluster colored in purple, comparing the true label in Fig.
-5h.
+4c) are assimilated into the majority cluster colored in green, comparing the true label in Fig.
+4h.
 The reason lies in that the deep models typically perform clustering in the representation space and minority clusters are underrepresented in the self-supervised learning owing to limited samples.
-(iii) Different from other deep models, the minority clusters are easier to be identified and discovered in both LSEnet and IsoSEL.
+(iii) Different from other deep models, the minority clusters are easier to be identified and discovered in both LSEnet and ASIL.
 We highlight that, the differentiable structural information objective effectively discovers node clusters without $K$ and is suitable to identify minority clusters in imbalanced graphs.
-In addition, compared to LSEnet, IsoSEL further mitigates the bias in clustering boundary with the hyperbolic reweighing proposed in Lorentz tree contrastive learning.
+In addition, compared to LSEnet, ASIL further mitigates the bias in clustering boundary with better hyperbolic partitioning tree representations.
 (a) Karate: Results
@@
 VIII-I Visualizing Hyperbolic Partitioning Tree
-We visualize the hyperbolic partitioning trees of IsoSEL.
-In IsoSEL, the partitioning tree is created in the Lorentz model of hyperbolic space.
+We visualize the hyperbolic partitioning trees of ASIL.
+In ASIL, the partitioning tree is created in the Lorentz model of hyperbolic space.
 Note the fact that the 2D Lorentz model is represented in 3D Euclidean space and is hard to be visualized.
@@
 With $\Psi$, we study the clustering results on two toy examples of Karate and Football datasets, and show the visualization in the 2D Poincaré disc in Fig.
-6.
+5.
 The partitioning tree with learnt clusters are plotted in Fig.
@@
 5 (c) and (d).
-Two important findings are summarized as follows: (i) The clustering results of IsoSEL are generally aligned with the ground truth distribution.
-(ii) The learnt embeddings of each clusters are well separated, showing the effectiveness of IsoSEL.
+Two important findings are summarized as follows: (i) The clustering results of ASIL are generally aligned with the ground truth distribution.
+(ii) The learnt embeddings of each clusters are well separated, showing the effectiveness of ASIL.
 IX Related Work
@@
 X Conclusion
-Our work bridges structural information with deep learning and provides a principled solution for graph clustering without without requiring a predefined number of clusters while accounting for the imbalance in real-world graphs.
-Theoretically, we establish differentiable structural information in continuous realm, and prove its effectiveness in clustering without $K$ and detecting minority clusters.
-Accordingly, we present the IsoSEL framework where LSEnet is designed to learn a deep partitioning tree via gradient backpropagation in the Lorentz model of hyperbolic space.
-Furthermore, we propose Lorentz tree contrastive learning with isometric augmentation to refine the deep tree, thereby enhancing cluster assignment.
-Extensive experiments demonstrate that IsoSEL effectively generates node clusters and mitigates the bias in the clustering boundary among imbalanced clusters.
+Our work presents a principled solution for graph clustering from an information-theoretic perspective, considering the practical challenges of unknown cluster number and imbalanced node distribution.
+To this end, we extend the classic structural entropy to the continuous realm, so that its partitioning tree can be learned via gradient backpropagation, uncovering clustering structures.
+Theoretically, we demonstrate its capacity in debiased clustering and connect it to contrastive learning, allowing for effective and efficient clustering via Augmented Structural Information Learning (ASIL).
+The key innovation is the seamless integration of partitioning tree construction and contrastive learning in hyperbolic space.
+Without requiring the number of clusters, ASIL achieves debiased node clustering with linear computational complexity in graph size.
+Empirically, we validate the superiority of ASIL against $20$ strong baselines on multiple benchmark datasets.